### v1

Abstract
The field of Fake Image Detection and Localization (FIDL) is highly fragmented, encompassing four domains: deepfake detection (Deepfake), image manipulation detection and localization (IMDL), artificial intelligence-generated image detection (AIGC), and document image manipulation localization (Doc). Although individual benchmarks exist in some domains, a unified benchmark for all domains in FIDL remains blank. The absence of a unified benchmark results in significant domain silos, where each domain independently constructs its datasets, models, and evaluation protocols without interoperability, preventing cross-domain comparisons and hindering the development of the entire FIDL field. To close the domain silo barrier, we propose ForensicHub, the first unified benchmark & codebase for all-domain fake image detection and localization. Considering drastic variations on dataset, model, and evaluation configurations across all domains, as well as the scarcity of open-sourced baseline models and the lack of individual benchmarks in some domains, ForensicHub: i) proposes a modular and configuration-driven architecture that decomposes forensic pipelines into interchangeable components across datasets, transforms, models, and evaluators, allowing flexible composition across all domains; ii) fully implements 10 baseline models (3 of which are reproduced from scratch), 6 backbones, 2 new benchmarks for AIGC and Doc, and integrates 2 existing benchmarks of DeepfakeBench and IMDLBenCo through an adapter-based design; iii) establishes an image forensic fusion protocol evaluation mechanism that supports unified training and testing of diverse forensic models across tasks; iv) conducts indepth analysis based on the ForensicHub, offering 8 key actionable insights into FIDL model architecture, dataset characteristics, and evaluation standards. Specifically, ForensicHub includes 4 forensic tasks, 23 datasets, 42 baseline models, 6 backbones, 11 GPU-accelerated pixel- and image-level evaluation metrics, and realizes 16 kinds of cross-domain evaluations. ForensicHub represents a significant leap forward in breaking the domain silos in the FIDL field and inspiring future breakthroughs. Code is available at: https://github.com/scu-zjz/ForensicHub.
1 Introduction
"The whole is more than the sum of its parts" - Aristotle
Fake images have become increasingly prevalent, driven by the rapid advancement of various digital image editing techniques in recent years. This highlights the importance of Fake Image Detection and Localization (FIDL), which aims to distinguish partially tampered and fully generated images from real ones. In FIDL, the term Detection refers to classification at the image level, while Localization targets a finer-grained segmentation of manipulated pixels at the pixel level.
However, the research efforts of FIDL have gradually split into four relatively independent research domains over time. 1) Deepfake detection (Deepfake) [37, 26, 25, 8, 38, 4, 66, 42, 28, 33, 54, 71, 49, 65, 67]: detects manipulations such as face swapping, expression editing, or feature replacement. 2) Image manipulation detection/localization (IMDL) [5, 17, 29, 35, 70, 53]: detects and localizes the tampering in natural images. 3) AI-Generated Image Detection (AIGC) [40, 18, 63, 59]: detects the images fully generated by deep generative models such as Stable Diffusion [46]. 4) Document Image Manipulation Localization (Document)  [48, 43, 6, 12, 52]: localizes the forgery of various forms of document images, including receipts, certificates, and identification materials, with a particular focus on detecting modifications to the printed text.
Although these domains have become isolated due to differences in application scenarios, manipulation types, and detection methods, there are still overlaps and similarities among them. As vision tasks, these four domains almost universally adopt SoTA detection or segmentatoin models as pre-trained backbones. Further, since the creators of fake images typically aim to preserve semantically plausible and realistic content, all four domains have placed considerable emphasis on designing low-level visual feature extractors to capture subtle, non-semantic discrepancies for reliable detection. Some research methodologies, such as contrastive learning, are commonly employed across these areas to mine discriminative features. We summarize SoTAs in four domains of the backbone, artifacts strategy, output type, and contribution in Table 1. The differences cause the four FIDL domains to become fragmented, but the similarities call for a unified perspective to understand them cohesively.
Table 1: Summary of representative methods from four forensic domains, detailing model design, backbone, artifact strategy, output format, and core contributions.
Task Model Backbone Artifact Strategy Output Type Contribution
Deepfake Capsule-Net [37](ICASSP19) VGG [50] Dynamic Routing Label Proposes a capsule network with dynamic routing and a VGG19 backbone.
RECCE [4](CVPR22) Xception [7] Reconstruction Label Proposes a graph-based framework leveraging reconstruction differences
SPSL [28](CVPR21) Xception [7] Phase Spectrum Label Proposes phase-spectrum fusion with Xception for face forgery detection.
UCF [66](ICCV23) Xception [7] Multi-task Disentanglement Label Proposes multi-task disentanglement with Xception for deepfake generalization.
SBI [49](CVPR22) EfficientNet [55] Frequency,Blending Boundaries Label Proposes self-blended images to improve deepfake detection generalization.
IMDL MVSS-Net [62](ICCV21) Resnet [19] BayarConv,Sobel Label,Mask Exploit noise and boundary artifacts via multi-view learning for manipulation detection.
CAT-Net [22](IJCV22) HRNet [56] DCT Mask Fuse RGB and DCT streams to learn compression artifacts for splice localization.
PSCC-Net [29](TCSVT22) HRNet [56] Multi-Resolution Conv Label,Mask Progressively refine masks with spatio-channel correlations for high-reso localization.
Trufor [17](CVPR23) Seformer [64] High-Reso,Multi-scale,Edge Label,Mask Fuse RGB and learned noise fingerprints to detect manipulations as anomalies.
IML-ViT [35](Arxiv) ViT [13] BayarConv,SRM Filter Mask Use ViT with high-reso, multi-scale edge-aware design for manipulation localization.
Mesorch [70](AAAI25) Conv. [31],Segfor. [64] DCT Mask Fuse micro and macro cues for mesoscopic image manipulation localization.
AIGC Dire [59](ICCV23) Resnet [19] Diffusion Reconstruction Label Use reconstruction error of diffusion for diffusion-generated images detection.
DualNet [63](APSIPA23) CNN SRM,Low Frequency Label Fuse SRM residual and low-frequency content streams for AIGC detection.
HiFiNet [18](CVPR23) HRNet [56] Multi-branch Feature Extractor Label,Mask Learn hierarchical fine-grained representations of forgery attributes.
Synthbuster [2](OJSP23) None Fourier Transform Label Leverage spectral artifacts in the frequency domain for diffusion detection.
UnivFD [40](CVPR23) CLIP-ViT [45] None Label Use pretrained vision-language model features for unified detection.
Document CAFTB [52](TOMM24) Resnet [19] SRM Mask Proposes CAFTB-Net with dual-branch and cross-attention.
TIFDM [12](TCE24) Resnet [19] None Mask Proposes a robust network with multiscale attention.
DTD [43](CVPR23) Conv. [31], Swin.[30] Frequency Mask Proposes DTD with frequency head and multi-view decoder.
FFDN [6](ECCV24) ConvNext [31] Wavelet, Frequency Mask Proposes FFDN combining visual enhancement and frequency decomposition
Although individual benchmarks exist in some domains, such as DeepfakeBench [67] for Deepfake and IMDLBenCo [36] for IMDL, a unified benchmark for all domains in FIDL remains blank. The absence of such a unified benchmark results in significant domain silos, where each domain independently constructs its datasets, models, and evaluation protocols without interoperability. Domain silos lead to redundant and uneven research across existing FIDL fields, and difficulty in establishing a general and unified FIDL approach, severely hindering the development of the entire FIDL field.
Therefore, establishing a unified benchmark for all domains is critically significant. However, such a benchmark faces the following challenges. Firstly, the drastic variations in datasets, models, and evaluation configurations across all domains require the benchmark to be sufficiently extendable and flexible in its design to support all domains. Secondly, compatibility with existing benchmarks is needed to reduce redundant research, while also addressing the scarcity of open-sourced baseline models and the absence of individual benchmarks in certain domains.
To this end, we propose ForensicHub, which: 1) proposes a modular and configuration-driven architecture that decomposes forensic pipelines into interchangeable components across datasets, transforms, models, and evaluators, allowing flexible composition across all domains; 2) fully implements 10 baseline models (3 of which are reproduced from scratch), 6 backbones, 2 new benchmarks for AIGC and Doc, and integrates 2 existing benchmarks of DeepfakeBench and IMDLBenCo through an adapter-based design. With the above efforts, ForensicHub serves as the first unified benchmark and codebase for all-domain fake image detection and localization. Building on ForensicHub, we establish an image forensic fusion protocol (IFF-Protocol) evaluation mechanism that supports unified training and testing of diverse forensic models across tasks. We conduct deep analysis on 8 issues that are of particular interest to researchers but have not yet been thoroughly investigated, offering new insights into FIDL model architecture, dataset characteristics, and evaluation standards. The introduction of ForensicHub bridges all domains within FIDL, breaks down domain silos, and inspires future breakthroughs.
Figure 1: Overview of our ForensicHub. It is compatible with DeepfakeBench and IMDLBenCo via adapters, and introduces new AIGC and Document benchmarks. ForensicHub allows datasets and models from any domain to be freely combined into custom pipelines.
2 Related Works
Fake image detection and localization span four sub-tasks, summarized in Appendix C. Despite rapid progress, a unified benchmark is lacking—each task uses isolated pipelines, limiting cross-task comparison.
DeepfakeBench [67] is a Deepfake detection benchmark specifically designed to address the lack of uniformity in data processing pipelines, leading to inconsistent data inputs for detection models. IMDLBenCo [36] is a benchmark and codebase for IMDL, aiming to compare IMDL models through a unified training and evaluation protocol. AIGCDetectBenchmark [68] is a repository for experiments on 9 AI-generated image detection methods.
These benchmarks provide models, datasets, and evaluation metrics within their respective tasks, but their underlying designs lack cross-task considerations, making them difficult to integrate across different detection scenarios. For example, DeepfakeBench is tightly coupled with Deepfake-specific data preprocessing steps, such as facial landmarks, while IMDLBenCo requires both datasets and models to output pixel-level masks. AIGCDetectBenchmark does not handle multi-GPU metric computation effectively. Additionally, none of them include a comprehensive set of image-level and pixel-level metrics. These limitations call for a new, unified, and flexible cross-task benchmark.
3 ForensicHub
In this section, we present our ForensicHub, which is a unified benchmark for all-domain fake image detection and localization designed for flexibility and extensibility, as illustrated in Figure 1.
Modular Architecture.
To accommodate different forensic tasks, ForensicHub is designed as a modular architecture consisting of four main components: Datasets, Transforms, Models, and Evaluators. Datasets handle the data loading process and are required to return fields that conform to the ForensicHub specification. Transforms handle the data pre-processing and augmentation for different tasks. Models, through alignment with Datasets and unified output, allow for the inclusion of various state-of-the-art image forensic models. Evaluators cover commonly used image- and pixel-level metrics for different tasks, and are implemented with GPU acceleration to improve evaluation efficiency during training and testing.
Configurable workflow.
ForensicHub provides a codeless approach for users to build training or testing workflows directly through the configuration of YAML files. Based on the modular architecture, users can select different evaluators to train and test any model on any dataset. ForensicHub also provides a code generator for customized purposes, allowing users to integrate with the benchmark with minimal coding effort.
Construction of ForensicHub.
To enable broad interoperability and reduce duplication of effort, ForensicHub adopts an adapter-based design [14] that ensures seamless integration with DeepfakeBench [67] and IMDLBenCo [36], two widely used benchmarks. This mechanism allows users to reuse existing models and datasets without major modification, while also supporting the definition of new models and benchmarks within ForensicHub under the unified protocol. This unified infrastructure simplifies cross-task benchmarking, supports reproducibility, and enables consistent evaluation across domains.
Specifically, ForensicHub supports all 10 models from IMDLBenCo for multi-domain and cross-domain evaluation. From DeepfakeBench, 27 out of 34 image-level detectors are compatible, including 5 general-purpose backbones and 9 domain-specific models that are not applicable to cross-task evaluation. The remaining 13 models support training or inference across different forensic domains. Therefore, 22 models of DeepfakeBench are included. ForensicHub fully implements 5 baseline models for AIGC and 5 baseline models for Document, with details in Sec. 4. In addition, ForensicHub includes 6 commonly used backbones. In total, ForensicHub covers 4 tasks, 23 datasets, 42 models, 6 backbones, and implements 11 commonly used image- and pixel-level metrics.
Datasets used in this paper are: Faceforencs++ [47], Celeb-DF [27], DFD [9], FaceShifter [23] and UADFV [26] for Deepfake, CASIA [11], COVERAGE [60], Columbia [20], IMD2020 [39], NIST16 [16], CocoGlide [17], and Autosplice [21] for IMDL, DiffusionForensics [59], GenImage [69] for AIGC, Doctamper [43], T-SROIE [58], OSTF [44], TPIC-13 [57], RTM [32] for Doc. A brief summary of each dataset is provided in Table 2, with more details in Appendix D.1.
Table 2: Summary of ForensicHub datasets. Pipeline indicates whether manipulations are manual, typically implying higher quality. Split shows if validation and test sets are provided.
Task Dataset Year Content Real Fake Annotation Pipeline Split
Deepfake FaceForensics++ [47] 2019 Face 45,388 127,209 Label Automatic Train,Test
Celeb-DF-v1 [27] 2020 Face 1,203 1,933 Label Automatic Train,Test
Celeb-DF-v2 [27] 2020 Face 5,620 10,800 Label Automatic Train,Test
DeepFakeDetection [15] 2019 Face 10,741 91,800 Label Automatic Train,Test
DFDC [9] 2020 Face 63,265 68,851 Label Automatic Train,Test
DFDCP [10] 2019 Face 5,901 11,321 Label Automatic Train,Test
FaceShifter [24] 2020 Face 4,479 4,479 Label Automatic Train,Test
UADFV [26] 2018 Face 1,548 1,551 Label Automatic Train,Test
IMDL CASIAv1 [11] 2013 General 0 920 Label,Mask Manual Train
CASIAv2 [11] 2013 General 7,491 5,123 Label,Mask Manual Train
COVERAGE [60] 2016 General 100 100 Label,Mask Manual Train
Columbia [20] 2006 General 183 180 Label,Mask Automatic Train
IMD2020 [11] 2020 General 414 2,010 Label,Mask Manual Train
NIST16 [16] 2019 General 873 564 Label,Mask Manual Train
CocoGlide [17] 2023 General 512 512 Label,Mask Automatic Train
Autosplice [21] 2023 General 2,273 3,621 Label,Mask Manual Train
AIGC DiffusionForensics [59] 2023 General 134,000 481,200 Label Automatic Train,Val,Test
GenImage [69] 2023 General 1,331,167 1,350,000 Label Automatic Train,Test
Document Doctamper [43] 2023 Document 0 170,000 Mask Automatic Train,Test
OSTF [44] 2025 Document 16,405 3,685 Mask Manual Train,Test
RealTextManipulation [32] 2025 Document 52,314 17,423 Mask Automatic+Manual Train,Test
T-SROIE [58] 2022 Document 21,268 4,326 Mask Manual Train,Test
Tampered-IC13 [57] 2022 Document 2,810 1,228 Mask Manual Train,Test
Models used in this paper are: Capsule-Net [37], RECCE [4], SPSL [28], UCF [66], and SBI [49] for Deepfake detection, MVSS-Net [5], CAT-Net [22], PSCC-Net [29], Trufor [17], IML-ViT [35], and Mesorch [70] for image manipulation and localization, Dire [59], DualNet [63], HiFiNet [18], Synthbuster [2], and UnivFD [40] for AIGC detection, DTD [43], FFDN [6], CAFTB [52], TIFDM [12] for document detection. These methods are from official repositories and our reimplementations. In addition, ForensicHub also selects 6 commonly used backbones in visual tasks, which are: Resnet [19], Xception [7], EfficientNet [55], Segformer [64], Swin Transformer [30], and ConvNext [31]. Details about models can be found in Appendix D.2.
Metrics used in this paper are: AP, MCC, TNR, TPR, AUC, ACC, F1, and IOU, with pixel- and image-level implementations shown in Fig. 1. Details of each metric can be found in Appendix D.3. In the evaluation, the threshold (if applicable) for all metrics is set to 0.5 to ensure fair comparison.
4 Benchmarks
In addition to being fully compatible with existing benchmarks, DeepfakeBench [67] and IMDLBenCo [36], ForensicHub further extends standardization efforts by introducing unified evaluation protocols for the AIGC and Document domains—two areas that previously lacked widely accepted benchmarks and codebases. We propose two protocols for two domains to evaluate generalization.
4.1 AI Generation Image Detection Benchmark
Datasets.
In the field of AIGC detection, the challenge in dataset construction usually lies not in obtaining a sufficient quantity of samples, since they can be easily generated using existing models, but in ensuring comprehensive coverage of a wide range of generative models. Therefore, we select only two commonly used public datasets: DiffusionForensics [59] and GenImage [69]. The former contains only diffusion-based generated images, while the latter covers a million-scale dataset constructed from eight SoTA generative models. Models are trained on DiffusionForensics and evaluated on different generative models within GenImage to assess generalization, as detection methods typically already achieve good performance on samples from the same generative model [69]. The detailed data splits are summarized in Table D.1.3.
Models.
ForensicHub implements five SoTA methods in AIGC detection: Dire [59], DualNet [63], HiFiNet [18], Synthbuster [2], and UnivFD [40], among which Synthbuster has no official open-source code and is fully reimplemented by us. More details about models and training settings can be found in Appendix E.1.
Results.
Table 3 in green background presents the AUC scores for image-level classification for AIGC benchmark, divided into in-domain results on the test set split of DiffusionForensics [59], and cross-domain results on different generative models and the total set on GenImage [69]. The results show that AIGC SoTAs generally achieve excellent performance on the DiffusionForensics test set, which shares the same source as the training set, and also perform well on datasets composed of diffusion-based generated images like ADM, VQDM, and GLIDE that are similar to the training data. However, the relatively poor generalization to generative models like Midjourney and Wukong highlights areas for improvement and provides guidance for future model development.
4.2 Document Image Manipulation Localization Benchmark
Table 3: AUC scores of image-level detectors. Models are tested in-domain on DiffusionForensics and cross-domain on GenImage sources. AverageC reflects cross-domain performance. In this table, each cell color denotes the model’s associated task domain:   Green indicates AIGC models,   Blue denotes IMDL models. Additionally, in other tables   Yellow indicates Deepfake models,   Orange indicates Document models, and   Gray indicates backbone models.
Method Within-Domain Cross-Domain
DiffusionForensics ADM BigGAN Midjourney VQDM GLIDE SD V1.4 SD V1.5 Wukong AverageC
DualNet [63](APSIPA23) 1.0000 0.9986 0.9172 0.8495 0.9901 0.9835 0.8127 0.8081 0.7925 0.8902
HiFiNet [18](CVPR23) 1.0000 0.9998 0.8407 0.7210 0.9999 0.9991 0.6765 0.6747 0.6675 0.8160
Synthbuster(OJSP23) 0.6662 0.6226 0.8550 0.4565 0.6308 0.7719 0.4606 0.4614 0.3902 0.5762
UnivFD(CVPR23) 0.9947 0.8400 0.9687 0.5427 0.9663 0.9541 0.7078 0.7072 0.6781 0.7920
MVSS-Net(ICCV21) 0.9992 0.9706 0.9520 0.6242 0.8894 0.9380 0.7207 0.7192 0.6089 0.7997
CAT-Net(IJCV22) 0.9985 0.9391 0.9327 0.6423 0.8226 0.7987 0.7171 0.7142 0.6156 0.7707
Trufor(CVPR23) 0.9999 0.9750 0.9316 0.7079 0.9340 0.9144 0.8683 0.8687 0.8018 0.8752
IML-ViT(Arxiv) 1.0000 0.9594 0.9152 0.8086 0.9577 0.9464 0.8924 0.8877 0.8006 0.8957
Mesorch(AAAI25) 0.9998 0.9754 0.9667 0.7011 0.9331 0.9253 0.8168 0.8179 0.7441 0.8582
Dataset.
Existing datasets for document image manipulation localization can be broadly categorized into two types: high-fidelity non-sliced datasets, including T-SROIE [58], OSTF [44], TPIC-13 [57], and RTM [32]; and sliced datasets, represented by Doctamper [43]. The primary difference lies in whether the images are preprocessed using patch-wise slicing.
To ensure consistency for downstream evaluation, we adopt the slicing strategy from Doctamper and apply it to the four non-sliced datasets, resulting in a unified format. Each dataset follows its original train/test split. Notably, Doctamper provides one training set and three distinct test sets—Doctamper-Test, Doctamper-FCD, and Doctamper-SCD—targeting different manipulation scenarios. The detailed data distribution is summarized in Table D.1.4.
Model.
We employ two open-source models, DTD [43] and FFDN [6], and reproduce two closed-source models, CATFB [52] and TIFDM [12]. All details are provided in Appendix D.2.4 and E.2.
Results.
Following the original protocols [43, 6], each detector is trained on its designated training split and evaluated on the corresponding test split. As shown in Table 4, three models consistently achieve top performance: FFDN and DTD, both designed specifically for document forensics, and Cat-Net, an IMDL-based model. Notably, all three methods incorporate JPEG-specific priors, such as DCT coefficients and quantization tables, highlighting the discriminative value of compression artifacts for manipulation localization in document images.
Table 4: Binary-F1 scores of document detectors. AverageD is the mean over three Doctamper test dataset, and AverageAll is averaged across all seven test datasets.
Model Doctamper T-SROIE-train OSTF-train TPIC-13-train RTM-train AverageAll
DocTamperFCD DocTamperSCD DocTamperTest AverageD T-SROIE-test OSTF-test TPIC-13-test RTM-test
CAFTB [52](TOMM24) 0.2917 0.3770 0.3275 0.3321 0.9165 0.6478 0.8394 0.2493 0.5213
DTD [43](CVPR23) 0.6856 0.7392 0.8031 0.7426 0.9205 0.5626 0.8341 0.1720 0.6739
FFDN [6](ECCV24) 0.8773 0.7392 0.8212 0.8126 0.9137 0.5586 0.8186 0.1589 0.6982
TIFDM [12](TCE24) 0.0896 0.2572 0.2585 0.2018 0.8942 0.5410 0.7972 0.0591 0.4138
MVSS-Net [5](ICCV21) 0.2066 0.3710 0.3810 0.3195 0.8756 0.5254 0.7864 0.1314 0.4682
PSCC-Net [29](TCSVT22) 0.3855 0.3931 0.4972 0.4253 0.9305 0.5697 0.7894 0.1971 0.5375
Cat-Net [22](IJCV22) 0.7600 0.6405 0.7644 0.7216 0.8726 0.5371 0.7947 0.1174 0.6410
IML-ViT [35](Arxiv) 0.4688 0.5117 0.4486 0.4764 0.8731 0.5128 0.7202 0.0920 0.5182
Trufor [17](CVPR23) 0.2613 0.3124 0.2517 0.2751 0.9095 0.6485 0.8197 0.2302 0.4905
Mesorch [70](AAAI25) 0.4231 0.4586 0.4651 0.4489 0.9049 0.4804 0.7607 0.1314 0.5177
However, this evaluation setting has a key limitation: all models are trained and tested within the same distribution, limiting the assessment of cross-domain generalization. To address this, we introduce a dedicated Doc Protocol, where models are trained on Doctamper and evaluated on four other document-level test sets. As shown in Table 5, PSCC-Net demonstrates superior generalization, highlighting the benefit of progressive spatial modeling for Doc manipulation localization.
Table 5: Binary F1 evaluation of models trained only on Doctamper and tested on both within-domain and cross-domain datasets. AverageW, AverageC, and AverageAll denote the average performance on Doctamper, external datasets, and all test sets, respectively.
Model Within-Domain Cross-Domin AverageAll
DocTamperFCD DocTamperSCD DocTamperTest AverageW T-SROIE OSTF TPIC-13 RTM AverageC
CAFTB [52](TOMM24) 0.2917 0.3770 0.3275 0.3321 0.2617 0.1194 0.3007 0.0328 0.1787 0.2444
DTD [43](CVPR23) 0.6856 0.7392 0.8031 0.7426 0.5245 0.1241 0.2835 0.0575 0.2474 0.4596
FFDN [6](ECCV24) 0.8773 0.7392 0.8212 0.8126 0.5330 0.2409 0.3572 0.0708 0.3005 0.5199
TIFDM [12](TCE24) 0.0896 0.2572 0.2585 0.2018 0.0582 0.0058 0.0134 0.0176 0.0238 0.1000
MVSS-Net [5](ICCV23) 0.2066 0.3710 0.3810 0.3195 0.1870 0.0373 0.1134 0.0268 0.0911 0.1890
PSCC-Net [29](TCSVT22) 0.3855 0.3931 0.4972 0.4253 0.5168 0.4414 0.5495 0.1255 0.4083 0.4156
Cat-Net [22](IJCV22) 0.7600 0.6405 0.7644 0.7216 0.6085 0.1777 0.3430 0.0630 0.2981 0.4796
IML-ViT [35](Arxiv) 0.4688 0.5117 0.4486 0.4764 0.4269 0.2101 0.2563 0.0764 0.2424 0.3427
Trufor [17](CVPR23) 0.2613 0.3124 0.2517 0.2751 0.2126 0.0464 0.1038 0.0342 0.0993 0.1746
Mesorch [70](AAAI25) 0.4231 0.4586 0.4651 0.4489 0.2937 0.1388 0.2408 0.0405 0.1785 0.2944
5 Image Forensic Fusion Protocol
Protocol.
To explore the performance of different models under a unified forensic protocol, we implement an image forensic fusion protocol (IFF-Protocol) inspired by CAT-Net’s training data construction strategy. The IFF-Protocol defines the training set as a combination of Deepfake, IMDL, AIGC, and Document data, where each training epoch samples an equal number of instances from each domain at random. During training, we select FaceForensics++ [47] from Deepfake, CASIAv2 [11] from IMDL, GenImage [69] from AIGC, and OSTF [44], RealTextManipulation, T-SROIE, and Tampered-IC13 from the Document. We use the smallest dataset, CASIAv2 with 12,641 samples, as the sampling number for each epoch. During testing, we evaluate the models directly on datasets from different domains without fine-tuning.
Implementation Details.
We resize images to 256×256 (except UnivFD, DTD and FFDN, see Appendix F for details) and apply only basic data augmentations, including flipping, brightness and contrast adjustment, compression, and Gaussian blur. All models are trained for 20 epochs using a cosine decay learning rate schedule, decreasing from $41e-$ 4 to $51e-$ 5. For models that output masks (IMDL and Doc), we apply max pooling to the final-layer feature maps to obtain a predicted label and compute the loss using only the label.
Table 6: Comparison of model parameters and FLOPs across representative architectures.
Model Params (M) FLOPs (G) Model Params (M) FLOPs (G) Model Params (M) FLOPs (G)
Resnet [19] $44.55$ $20.59$ Capsule-Net [37] $3.90$ $32.43$ Mesorch [70] $85.75$ $57.95$
Xception [7] $22.86$ $12.05$ RECCE [4] $25.83$ $16.18$ DualNet [63] $7.99$ $66.34$
EfficientNet [55] $19.34$ $4.14$ SPSL [28] $20.81$ $12.06$ HiFiNet [18] $6.89$ $145.00$
ConvNeXt [31] $50.22$ $22.74$ MVSS-Net [5] $147.00$ $83.08$ UnivFD [40] $428.00$ $156.00$
Segformer [64] $44.07$ $16.29$ Trufor [17] $68.70$ $112.00$ DTD [43] $67.07$ $272.00$
Swin [30] $49.61$ $28.65$ IML-ViT [35] $91.78$ $80.37$ FFDN [6] $89.20$ $453.00$
Model Efficiency.
We test the parameters and FLOPs of the backbones and SoTAs of each domain in Table 6. It can be observed that model efficiency is often related to the task’s application scenario. For example, Deepfake models are typically lightweight to support real-time video detection, while IMDL models, which focus on pixel-level classification, often adopt more complex and heavier architectures. These efficiency preferences can influence the experimental results under the IFF-Protocol.
Table 7: Cross-domain AUC evaluation of models trained under IFF-Protocol.
Method Deepfake IMDL AIGC Document Average
FF-c40 CDFv2 DFD Columbia IMD2020 Autosplice DF GenImage T-SROIE OSTF RTM
Resnet [19](CVPR16) 0.681 0.730 0.793 0.482 0.533 0.738 0.619 0.797 0.951 0.681 0.662 0.697
Xception [7](CVPR17) 0.728 0.719 0.870 0.465 0.537 0.756 0.757 0.980 0.966 0.762 0.734 0.752
EfficientNet [55](ICML19) 0.504 0.535 0.517 0.623 0.506 0.483 0.544 0.597 0.884 0.581 0.512 0.571
Segformer [64](NIPS21) 0.691 0.748 0.862 0.409 0.562 0.824 0.805 0.998 0.980 0.866 0.736 0.771
Swin [30](ICCV21) 0.771 0.746 0.901 0.636 0.631 0.864 0.915 0.999 0.990 0.856 0.758 0.824
ConvNext [31](CVPR22) 0.794 0.784 0.911 0.625 0.598 0.825 0.895 1.000 0.994 0.849 0.762 0.822
Capsule-Net [37](ICASSP19) 0.613 0.660 0.699 0.330 0.527 0.745 0.546 0.971 0.946 0.704 0.670 0.674
RECCE [4](CVPR22) 0.634 0.602 0.727 0.506 0.492 0.642 0.684 0.906 0.542 0.688 0.555 0.634
SPSL [28](CVPR21) 0.730 0.726 0.876 0.419 0.545 0.759 0.770 0.987 0.972 0.769 0.738 0.754
MVSS-Net [5](ICCV21) 0.713 0.700 0.857 0.298 0.539 0.795 0.671 0.994 0.978 0.806 0.741 0.736
Trufor [17](CVPR23) 0.642 0.698 0.832 0.306 0.564 0.808 0.726 0.996 0.979 0.805 0.732 0.735
IML-ViT [35](Arxiv) 0.750 0.726 0.851 0.483 0.556 0.819 0.627 0.991 0.972 0.800 0.703 0.753
Mesorch [70](AAAI25) 0.767 0.814 0.867 0.285 0.570 0.773 0.629 0.996 0.982 0.819 0.739 0.749
DualNet [63](APSIPA23) 0.637 0.552 0.540 0.268 0.517 0.748 0.899 0.988 0.935 0.658 0.657 0.673
HiFiNet [18](CVPR23) 0.587 0.611 0.648 0.745 0.534 0.677 0.575 0.756 0.937 0.663 0.615 0.668
UnivFD [40](CVPR23) 0.690 0.671 0.798 0.886 0.786 0.785 0.742 0.813 0.938 0.684 0.569 0.760
DTD [43](CVPR23) 0.498 0.520 0.490 0.679 0.498 0.506 0.457 0.499 0.748 0.595 0.496 0.544
FFDN [6](ECCV24) 0.714 0.699 0.871 0.553 0.624 0.927 0.999 1.000 0.997 0.893 0.782 0.824
Benchmark Result.
Table 7 shows the AUC scores of backbones and domain-specific SoTA methods on datasets from four domains under the IFF-Protocol, in which DFD refers to DeepFakeDetection [15], DF refers to DiffusionForensics [59], and RTM refers to RealTextManipulation [32]. We provide detailed results in Appendix F.
The results show that surprisingly, visual backbones such as ConvNeXt [31] and Swin Transformer [30] outperform almost all domain-specific SoTA methods, indicating that backbones demonstrate greater potential when trained on more unified fake images. Meanwhile, domain-specific SoTAs do not necessarily retain their superiority within their own tasks. For instance, UnivFD [40], a CLIP-based fine-tuned model for AIGC detection, demonstrates strong performance on the IMD2020 [39] from IMDL, revealing valuable insights into the transferability of cross-task methods.
From a task perspective, although IMDL target shifts from pixel-level to image-level classification, it remains challenging due to significant distribution differences across datasets in terms of size and manipulation types. In contrast, AIGC benefits from training on sufficient data from diverse generative models, resulting in higher detection accuracy. This observation reminds us that it is essential not only to include a comprehensive range of manipulation types in the training data but also to focus on enhancing the generalization ability of models.
6 Experiments
Based on ForensicHub, we conduct cross-task experiments, which have been less explored in previous research. The similarities and differences among detection methods across different tasks lead us to the following questions: 1) Are low-level feature extractors effective across all tasks? 2) Do detection methods from one task remain effective when transferred to another task? We answer the above questions through extensive experiments.
6.1 Effectiveness of Low-Level Feature Extractors
Since each domain has proposed specific feature extractors, to explore their effectiveness under a unified domain, we conduct experiments using 6 backbones combined with 4 different extractors in the shallow layer under the aforementioned IFF-Protocol setting. The extractors are BayarConv [3] for noise aritfacts, Sobel [5] for edge arifacts, DCT [1] and FPH [43] for frequency artifacts. Details of each extractor can be found in Appendix G.1.
Results in Table 8 show the AUC differences between versions of each backbone using the four different feature extractors and those without, averaged across all test datasets for each task. All backbones except EfficientNet [55] show performance drops after using feature extractors, indicating that under the IFF-Protocol, where training data includes sufficient manipulation types and image quantity, models do not rely on the additional information provided by feature extractors. However, due to its lightweight nature, EfficientNet still benefits from the use of feature extractors. The results suggest that feature extractors may only be beneficial for detection on small-scale datasets, with limited manipulation types, or when using lightweight models. Details of each domain test datasets AUC scores can be found in Appendix G.2.
Table 8: Mean AUC differences between extractor-enhanced models and their plain counterparts. Positive values (red) indicate gains; negative (blue) indicate drops.
Task Extractor = Sobel [5] Extractor = Bayar[3]
ResNet Xception EfficientNet SwinTransformer SegFormer ConvNext ResNet Xception EfficientNet SwinTransformer SegFormer ConvNext
AIGC -0.0446 -0.0513 0.0359 -0.0983 -0.1047 -0.1374 -0.0582 -0.0258 -0.0105 -0.0709 -0.0705 -0.0672
Deepfake -0.0887 -0.0163 0.0273 -0.1319 -0.1262 -0.1657 -0.0749 -0.0103 0.0088 -0.0963 -0.0973 -0.0878
Document -0.0450 -0.0268 0.0063 -0.0950 -0.0815 -0.1228 -0.0333 -0.0058 0.0025 -0.0760 -0.0653 -0.0885
IMDL 0.0093 -0.0203 0.0058 -0.0838 -0.0642 -0.0687 -0.0118 -0.0227 0.0085 -0.0938 -0.0735 -0.0913
Extractor = FPH [43] Extractor = DCT [1]
ResNet Xception EfficientNet SwinTransformer SegFormer ConvNext ResNet Xception EfficientNet SwinTransformer SegFormer ConvNext
AIGC -0.1314 -0.1060 0.1425 -0.0940 -0.1018 -0.5071 -0.0173 -0.0009 0.0081 -0.0429 -0.0339 -0.0704
Deepfake -0.1380 -0.1003 0.0596 -0.0585 -0.1404 -0.3443 0.0055 -0.0062 0.0108 -0.0078 -0.0153 -0.0993
Document -0.0525 -0.0645 0.0158 -0.0690 -0.0965 -0.2015 -0.0088 0.0000 0.0063 -0.0493 -0.0213 -0.0820
IMDL -0.0128 -0.0240 0.0367 -0.0917 -0.0933 -0.0848 0.0102 0.0067 0.0087 -0.0565 -0.0493 -0.0983
6.2 Transferability of Task-Specific Detectors
6.2.1 Cross-Evaluation Between IMDL and Document Benchmarks
Current Document-level detectors’ input–output formats are fully compatible with those of Image Manipulation Detection and Localization models. Leveraging this consistency, we perform a bidirectional evaluation: IMDL detectors are tested on the Document benchmark, and Document detectors are tested on the IMDL benchmark. This cross-testing enlarges the effective model pool for each benchmark and allows us to probe detector generality beyond their original task scopes.
IMDL → Document.
Table 4 reports the within-domain results obtained under the original Document benchmark split, whereas Table 5 presents the cross-domain scores produced by our newly introduced generalization protocol. Across both settings, IMDL detectors demonstrate strong competitiveness in the document forensics scenario. In the conventional split, the Cat-Net [22, 43, 6] family achieves the best average F1, confirming the merit of its hierarchical “cat-net” paradigm. Under the more challenging cross-domain evaluation, PSCC-Net [29] displays markedly better generalization, suggesting that progressive spatial modeling captures cues for document manipulation localization. We expect future work to further investigate the underlying mechanisms behind PSCC-Net.
Document → IMDL.
Following the MVSS training protocol [36], all Document-oriented models are trained on the CASIAv2 dataset [11] and evaluated on five standard IMDL test sets. As shown in Table 9, the dual-branch architecture of CAFTB [52] achieves the best overall performance among Document models when transferred to IMDL tasks—an outcome that aligns with the design philosophy of the current SoTA model Mesorch [70], which also emphasizes dual-branch learning.
Table 9: Pixel-level binary F1 evaluation on IMDL benchmarks for document-trained detectors.
Model CASIAv1 COVERAGE Columbia IMD2020 NIST16 Average
CAFTB [52](TOMM24) 0.6234 0.2557 0.6731 0.3526 0.3770 0.4564
DTD [43](CVPR23) 0.3535 0.1482 0.6470 0.1749 0.1811 0.3009
FFDN [6](ECCV24) 0.5012 0.2670 0.6085 0.2516 0.2957 0.3848
TIFDM [12](TCE24) 0.3675 0.2087 0.3572 0.1634 0.2333 0.2660
6.2.2 Extending IMDL Detectors to AIGC and Deepfake Benchmarks
IMDL models are designed to produce both pixel-level masks and image-level labels, with most architectures incorporating classification heads alongside segmentation branches. This dual-output design enables direct application to tasks like AIGC and Deepfake detection. For models without label heads, image-level scores are obtained via max-pooling over the predicted masks.
IMDL → AIGC.
We fine-tune representative IMDL detectors on the training split of the AIGC benchmark and report cross-generator performance in Table 3. The training settings and other configurations are consistent with those used in the previously mentioned AIGC benchmark setup. The results show that techniques from IMDL, such as noise print (TruFor) and multiscale analysis (IML-ViT), remain effective for AIGC detection.
IMDL → Deepfake.
We train each IMDL detector on the FF++-c23 training split and evaluate on all remaining deep-fake test sets; the scores are given in Table 10. When compared to all baselines in the deepfakebench [67], Cat-Net attains the best performance in the within-domain setting, while Mesorch achieves the highest average accuracy in the cross-domain evaluation, establishing new state-of-the-art results in both regimes.
Table 10: Image-level AUC evaluation of IMDL-based detectors trained on FF++-c23 and tested across both within-domain and cross-domain deepfake benchmarks.
Model Within Domain Evaluation Cross Domain Evaluation
FF-c23 FF-c40 FF-DF FF-F2F FF-FS FF-NT Avg. CDFv1 CDFv2 DFD DFDC DFDCP Fsh UADFV Avg.
CAT-Net [22](IJCV22) 0.9855 0.8351 0.9946 0.9891 0.9905 0.9689 0.9606 0.7489 0.7511 0.8086 0.7196 0.7043 0.6626 0.9282 0.7605
IML-ViT [35](Arxiv) 0.9614 0.8483 0.9810 0.9728 0.9793 0.9206 0.9439 0.7250 0.7419 0.7770 0.7516 0.7237 0.6149 0.8718 0.7437
Mesorch [70](AAAI25) 0.9815 0.8445 0.9889 0.9890 0.9859 0.9636 0.9589 0.7900 0.7649 0.8468 0.7806 0.7345 0.6638 0.9623 0.7918
MVSS-Net [5](ICCV21) 0.9723 0.8140 0.9779 0.9865 0.9868 0.9458 0.9472 0.6969 0.6839 0.7749 0.6764 0.6785 0.5841 0.8871 0.7117
PSCC-Net [29](TCSVT22) 0.7147 0.6536 0.7751 0.6414 0.6460 0.6065 0.6729 0.4602 0.5898 0.5565 0.5945 0.5845 0.6574 0.8216 0.6092
Trufor [17](CVPR23) 0.7303 0.6848 0.8087 0.7687 0.7084 0.6326 0.7223 0.6551 0.6115 0.5908 0.6017 0.5749 0.5977 0.9526 0.6549
7 Conclusion
This paper proposes ForensicHub, the first unified benchmark and codebase for all-domain fake image detection and localization. It adapts existing benchmarks and extends to other domains. Based on the extensive cross-domain experiments, we summarize 8 key actionable insights for future research:
1) In Doc, PSCC-Net exhibits strong generalization, while Cat-Net effectively adapts to synthetic manipulations, offering valuable Doc model designs. 2) In IMDL, parallel architecture models like CAFTB and Mesorch achieve leading performance, suggesting the effectiveness of multi-branch modeling. 3) Frequency-strategy models like Cat-Net and Mesorch consistently perform well, highlighting the potential of frequency features for FIDL. 4) Less-explored backbones like ConvNeXt and Swin Transformer outperform nearly all domain SoTAs under IFF-Protocol. 5) Shallow concatenation of feature extractors tends to negatively impact performance when the dataset is large and contains a wide variety of manipulation types, while lightweight models such as EfficientNet can benefit from this approach. 6) Current AIGC and Doc evaluations often neglect generalization, leading to overestimated performance. We recommend our proposed AIGC and Doc protocols for future work, which explicitly encourage generalization-aware model design. 7) Existing AIGC and Deepfake datasets are often too simple and lack diversity, limiting meaningful comparisons. Future benchmarks should aim for greater complexity and realism. 8) For all-domain scenarios, we recommend our IFF-Protocol to enable more comprehensive evaluation.
In conclusion, ForensicHub represents an important step toward breaking down domain silos across four fields, offering new insights into FIDL future research across model architecture, dataset characteristics, and evaluation standards.
Appendix A Limitations
Certain experiment limitations still remain. For example, in our study of feature extractors, we only apply shallow fusion of features. More advanced fusion strategies that can better exploit the proposed features remain to be explored. ForensicHub will gradually improve and expand its experiments in future versions.
Appendix B Author Contributions
The author contributions are: Bo Du: framework design, construction of AIGC benchmark, IFF-Protocol, feature extractors, cross-domain evaluation, and manuscript writing. Xuekang Zhu: framework design, construction of benchmark adapters, Document benchmark, cross-domain evaluation, and manuscript writing. Xiaochen Ma: framework design, framework code optimization and manuscript writing. Chenfan Qu: construction of Document benchmark and manuscript writing. Kaiwen Feng: construction of IFF-Protocol, feature extractors and manuscript writing. Zhe Yang: construction of AIGC benchmark. Chi-Man Pun: project advising. Jian Liu: general project advising. Jizhe Zhou: project supervisor and manuscript writing.
Appendix C Task Definitions and Detection Paradigms
To provide contextual understanding of the forensic domains included in our benchmark, we summarize the goals, characteristics, and representative modeling approaches for each task below.
C.1 Deepfake Detection
Deepfake detection aims to identify whether faces in an image have been manipulated, typically formulated as an image-level binary classification task. According to [67], current methods fall into three categories: naive, spatial, and frequency detectors. These approaches primarily target artifacts specific to facial manipulation, such as biological signals, spatial inconsistencies, frequency abnormalities, and auto-learned clues. It’s important to note that artifacts characteristic of Deepfake images often differ from those found in other types of image manipulation.
C.2 Image Manipulation Detection and Localization
IMDL task requires two types of assignments [36]: image-level detection to determine whether manipulation has occurred, and pixel-level localization to identify the manipulated regions. IMDL models are typically composed of a backbone and a low-level feature extractor to capture artifacts left by manipulation, such as edge artifacts [5, 35], frequency artifacts [22], and noise artifacts [17, 5, 70]. IMDL models are generally designed to detect manipulations in natural images rather than targeting specific types of tampering, such as facial forgeries.
C.3 AI-Generated Image Detection
AI-Generated Image Detection focuses on identifying whether an image is generated by generative models, performing binary classification at the image level only. Existing classifiers typically detect AIGC images by leveraging artifacts that differentiate them from real images, such as discrepancies in spatial feature space [40, 18], frequency inconsistency [41, 63, 2], and fingerprints left by specific generative models like diffusion models [59, 51, 34]. As a rapidly evolving technology, AIGC presents challenges to detection methods due to the artifacts left by deep generative models, which differ significantly from those found in traditional manual manipulations.
C.4 Document Image Manipulation Localization
Document Image Manipulation Localization focuses on identifying tampered text on images. Tampered text regions are usually small in size, with subtle appearance anomalies and fewer edge artifacts, due to consistent backgrounds and fonts [52, 48]. Consequently, methods designed for detecting forgeries in natural and face images usually do not perform well when applied directly to this task [12, 58]. To overcome this difficulty, recent studies propose to model the block artifact grids [43] or the texture differences [57], etc. Despite progress, accurately detecting forged text against elaborate tampering processes, advanced text editing models, and diverse image styles remains an open challenge [32, 44].
Appendix D Details of ForensicHub Construction
D.1 Datasets
D.1.1 Deepfake
Most Deepfake datasets are video-based. Following the protocol of DeepfakeBench [67], we extract 32 equally spaced frames from each video to form the image-based datasets used in our experiments.
FaceForensics++.
FaceForensics++ (FF++) [47] is the most widely used benchmark for Deepfake detection. It provides real and fake data generated using four manipulation methods: DeepFakes (FF-DF), Face2Face (FF-F2F), FaceSwap (FF-FS), and NeuralTextures (FF-NT). These four subsets share the same test real images but differ in the fake generation methods. The full dataset contains 27,472 real and 109,800 fake images.
The training set includes 22,993 real and 91,891 fake images. The test set contains 4,479 real images, shared across four manipulation subsets, with fake image counts of 4,473 (FF-DF), 4,480 (FF-F2F), 4,477 (FF-FS), and 4,479 (FF-NT), respectively.
These subsets are designed to assess how detection models perform against different generation types.
Celeb-DF-v1.
Celeb-DF-v1 [27] was released in 2020 with improved realism over early datasets. It includes 7,946 real and 25,362 fake training images, and 1,203 real and 1,933 fake test images.
Celeb-DF-v2.
Celeb-DF-v2 [27] expands on v1 in both quality and size. The training set contains 9,524 real and 179,777 fake images. The test set includes 5,620 real and 10,800 fake images.
DeepFakeDetection.
The DeepFakeDetection (DFD) [15] dataset, released by Google and Jigsaw, provides 10,741 real and 91,800 fake images for both training and testing. It is widely used for large-scale pretraining and evaluation.
DFDCP.
The DFDCP [10] dataset introduces post-compressed versions of fake images to simulate real-world distortions. The training set contains 22,425 real and 103,631 fake images. The test set includes 5,901 real and 11,321 fake images.
DFDC.
The Deepfake Detection Challenge (DFDC) [9] dataset, provided by Facebook, contains only a test set with 63,265 real and 68,851 fake images. The training set is not publicly available.
FaceShifter.
The FaceShifter [23] dataset offers 22,993 real and 22,968 fake training images, and 4,479 real and 4,479 fake images for testing. It is typically used to assess model generalization to unseen generation techniques.
UADFV.
UADFV [26] is one of the earliest Deepfake datasets. Both the training and test sets contain 1,548 real and 1,551 fake images. Due to its small size and early generation style, it is mostly used for cross-dataset evaluation.
D.1.2 IMDL
Due to the difficulty of annotation, IMDL datasets are typically small in scale. Information on datasets such as CASIA [11], Columbia [20], COVERAGE [60], IMD2020 [39], and NIST16 [16] can be found in IMDLBenCo [36]. Notably, considering the recent rise of AI-based partial image inpainting manipulations, we include two inpainting datasets generated using deep generative models: CocoGlide [17] and Autosplice [21]. CocoGlide and AutoSplice include 512 and 3621 images edited by GLIDE diffusion model and DALL-E2, respectively.
D.1.3 AIGC
DiffusionForensics.
DiffusionForensics [59] is a dataset constructed to facilitate the evaluation of detectors targeting diffusion-generated images. It comprises real and synthetic images across three representative domains: LSUN-Bedroom, ImageNet, and CelebA-HQ. The dataset includes outputs from a variety of diffusion models, covering unconditional, class-conditional, and text-to-image generation paradigms. For each image, the dataset provides a triplet: the source image, its reconstruction, and the corresponding DIRE image, enabling more detailed forensic analysis. The design of DiffusionForensics supports both training and testing, with subsets carefully split for each purpose. By encompassing a wide range of generative models and image domains, it serves as a comprehensive benchmark for assessing the generalization and robustness of diffusion image detectors.
GenImage.
GenImage [69] is a large-scale dataset developed to advance the detection of AI-generated images. It contains over one million pairs of synthetic and real images, covering a wide range of image categories. The synthetic images are produced using state-of-the-art generative models, including advanced diffusion models and GANs. GenImage includes the generative models of ADM, BigGAN, Midjourney, VQDM, GLIDE, Stable Diffusion V1.4, Stable Diffusion V1.5, Wukong. Each generative model generates nearly the same numbers of images (approximately 168750), with a total number of 1,350,000 of fake images. GenImage enables the evaluation of detectors under realistic conditions through two tasks: cross-generator classification, which assesses generalization across different generative models, and degraded image classification, which tests robustness to image quality degradation such as compression, blurring, and low resolution. By combining scale, diversity, and challenging evaluation settings, GenImage provides a comprehensive benchmark for developing reliable fake image detectors.
D.1.4 Document
DocTamper.
DocTamper [43]was introduced in 2023 and has become the most widely used dataset for document tampering localization. It contains fully synthetic manipulations on various photographed documents, such as contracts, receipts, invoices, and books. The tampering types include copy-move, splicing, and print-based edits. All images have been preprocessed by cropping to $512\times 512$ resolution, and the corresponding pixel-level masks are cropped accordingly. The training set contains 120,000 fake images, while the test set is split into three subsets: DocTamper-Test (30,000 fake), DocTamper-FCD (2,000 fake), and DocTamper-SCD (18,000 fake). Clean images are not included.
T-SROIE.
Released in 2022, T-SROIE [58] is the first dataset to localize AIGC-style tampering in scanned receipts using a modern IML approach. It contains text tampered by SR-Net [61] and was originally provided as high-resolution uncropped images. To ensure consistency with DocTamper, we apply the same cropping strategy to resize all images to $512\times 512$, and crop the corresponding pixel-level masks in the same manner. After cropping, the training set consists of 12,769 real and 2,747 fake images; the test set contains 8,499 real and 1,579 fake images.
RTM.
RTM [32] was introduced in 2025 and includes both synthetic and manually manipulated document images. The dataset covers a wide range of manipulation types, including copy-move, splicing, print, and erasure, across diverse document types such as scanned forms. The original high-resolution images are not cropped, so we apply the DocTamper-style cropping strategy to obtain $512\times 512$ resolution images, along with their aligned masks. After cropping, the RealTextManipulation-Test set includes 22,334 real and 3,444 fake samples.
OSTF.
OSTF[44], proposed in 2025, contains natural scene texts tampered by eight different AIGC-based text editing models. It focues on evaluating model generalization ability across unseen text tampering models and unseen image styles. Since the original images are high-resolution and unaligned, we perform $512\times 512$ cropping using the DocTamper protocol, and apply the same transformation to the associated masks. The resulting training set includes 1,729 real and 639 fake samples; the test set includes 14,676 real and 3,046 fake samples.
Tampered-IC13.
Tampered-IC13, released in 2022, contains naturally captured scene texts tampered by the AIGC text editing model SR-Net [61]. It also lacks predefined cropping, so we apply the DocTamper-style image and mask cropping to $512\times 512$. After preprocessing, the training set includes 1,729 real and 639 fake images; the test set includes 1,081 real and 589 fake images.
D.2 Models
D.2.1 Deepfake
For Deepfake detection, we design an adapter to directly align with the 27 image-based detectors provided in DeepfakeBench [67]. These detectors cover diverse architectures and training settings. For full details, we refer to the official DeepfakeBench documentation.
D.2.2 IMDL
We adapt all nine detection models from IMDLBenCo [36] via adapters. For detailed information on these models, please refer to the official IMDLBenCo documentation.
D.2.3 AIGC
Dire.
Dire [59] is a novel approach designed to detect diffusion-generated images by leveraging a unique image representation called Diffusion Reconstruction Error (DIRE). Unlike existing detectors, which often struggle to distinguish between real and diffusion-generated images, DIRE measures the reconstruction error between an input image and its counterpart reconstructed by a pre-trained diffusion model. It has been observed that while diffusion-generated images can be effectively reconstructed by a diffusion model, real images cannot, making DIRE a valuable tool for distinguishing between the two. DIRE is robust to various perturbations and generalizes well across different diffusion models, even those not seen during training. Extensive experiments on a comprehensive benchmark dataset demonstrate that DIRE outperforms previous detection methods in identifying AI-generated images, establishing it as a powerful tool for diffusion-based image forensics.
DualNet.
DualNet [63] is a novel detection method developed to address the challenges posed by AI Generated Content (AIGC), particularly text-to-image models like DALL·E2 and DreamStudio. Unlike traditional computer-generated graphics (CG), AIGC images are inherently more deceptive and require less human intervention, making conventional CG detection methods inadequate. To improve detection, DualNet employs a robust dual-stream network consisting of a residual stream and a content stream. The residual stream uses the Spatial Rich Model (SRM) to extract texture information from images, while the content stream captures low-frequency forged traces, providing complementary insights. These two streams are connected through a cross multi-head attention mechanism to enhance information exchange. Extensive experiments on two text-to-image databases and traditional CG benchmarks, such as SPL2018 and DsTok, demonstrate that DualNet consistently outperforms existing detection methods across a range of image resolutions, showing superior robustness and generalization capabilities.
HiFiNet.
HiFiNet [18] is a novel framework designed to address the challenges of image forgery detection and localization (IFDL), particularly when distinguishing between images generated by CNN-based synthesis and image-editing techniques. Due to the significant differences in forgery attributes between these domains, traditional methods struggle to provide a unified solution. HiFiNet tackles this issue by employing a hierarchical fine-grained approach for IFDL representation learning. The method first represents forgery attributes with multiple labels at different levels and performs fine-grained classification using the hierarchical dependencies between them. This encourages the model to learn both comprehensive features and the inherent hierarchical nature of various forgery attributes, improving the detection and localization performance. HiFiNet consists of three key components: a multi-branch feature extractor that classifies forgery attributes at different levels, and localization and classification modules that segment pixel-level forgery regions and detect image-level forgery, respectively. The effectiveness of HiFiNet is demonstrated through experiments on seven different benchmarks, showing significant improvements in both IFDL and forgery attribute classification tasks.
Synthbuster.
Synthbuster [2] is a detection method specifically designed to identify images generated by diffusion models, a type of AI-based generative technique that has gained popularity due to its ability to produce photo-realistic images from simple text prompts. While older detection methods targeting Generative Adversarial Networks (GANs) exist, they are insufficient for detecting images from advanced diffusion models. Synthbuster addresses this gap by focusing on the unique frequency artifacts left behind during the diffusion process. The method uses spectral analysis of the Fourier transform of residual images to highlight these artifacts, enabling the distinction between real and synthetic images. Synthbuster demonstrates strong detection capabilities even in the presence of mild JPEG compression and generalizes effectively to unseen models. This novel approach aims to enhance forensic techniques for detecting AI-generated images and encourages further research into this emerging field.
UnivFD.
UnivFD [40] is a novel approach designed to address the growing need for general-purpose fake image detectors, particularly in the face of rapidly evolving generative models. Traditional methods, which rely on training deep networks for real-vs-fake classification, struggle to detect images generated by newer models when trained only on older generative models like GANs. Analysis reveals that such classifiers become asymmetrically tuned, with the "real" class effectively acting as a catch-all for any image that isn’t fake, leading to poor performance when confronted with images generated by models not seen during training. To overcome this, UnivFD introduces a novel strategy: performing real-vs-fake classification without explicit training, using a feature space that is not designed to distinguish between real and fake images. By leveraging the feature space of large pretrained vision-language models and applying simple methods like nearest neighbor and linear probing, UnivFD achieves surprisingly strong generalization.
D.2.4 Document
PS-Net.
We reproduce PS-Net [48], a tampering localization model that refines both the input and output stages to enhance detection performance. At the input level, a Multi-View Enhancement (MVE) module fuses RGB, noise residual, and texture features to capture richer tampering traces. At the output level, Progressive Supervision (PS) applies multi-scale BCE losses to exploit hierarchical localization cues, while a Detection Assistance (DA) module introduces KL loss to align detection and localization branches. PS-Net demonstrates strong performance on DocTamper, effectively combining fine-grained supervision and global consistency.
CAFTB-Net.
We reproduce CAFTB-Net [52], a dual-branch network designed for document forgery localization in complex and noisy environments. It consists of a Spatial Information Extraction Branch (SIEB) and a Noise Feature Extraction Branch (NFEB), with the latter leveraging a Spatial Rich Model (SRM) filter to extract tampering cues. A Cross-Attention Fusion Module (CAFM) integrates both branches to enhance localization. CAFTB-Net achieves strong performance across benchmarks, particularly in detecting subtle and diverse manipulations.
TIFDM.
We reproduce TIFDM [12], which performs document forgery localization by modeling spatial. It processes RGB and uses attention mechanisms and a multi-scale decoder to improve localization. TIFDM shows robust generalization to mixed tampering types, including splicing, erasure, and generative edits.
DTD.
DTD (Document Tampering Detector) [43] introduces a multi-modality framework for detecting tampered text in document images. It integrates both RGB visual features and frequency cues extracted from JPEG compression artifacts via a dedicated Frequency Perception Head (FPH). A Swin-Transformer encoder combined with a Multi-view Iterative Decoder (MID) enables the model to capture subtle and dispersed tampering signals. Furthermore, DTD incorporates Curriculum Learning for Tampering Detection (CLTD), training the model in an easy-to-hard strategy to enhance robustness against varying compression levels. Extensive evaluations on DocTamper and T-SROIE datasets show that DTD achieves state-of-the-art performance, particularly in scenarios with heavy JPEG compression and complex document layouts.
FFDN.
FFDN (Feature Fusion and Decomposition Network) [6] tackles the challenge of subtle tampering in document images by jointly modeling spatial and frequency domains. It introduces a Visual Enhancement Module (VEM) that fuses visual features with frequency-aware representations using an attention mechanism, and a Wavelet-like Frequency Enhancement (WFE) module that explicitly decomposes features into high- and low-frequency components to capture faint tampering traces. This dual-path architecture enhances both perceptibility and robustness. Evaluated on DocTamper and T-SROIE, FFDN significantly outperforms previous methods, especially in detecting small tampered regions under compression and noise.
D.3 Metrics
AP (Average Precision).
Average Precision (AP) is calculated as the area under the precision-recall curve. It is defined as:
[…]
where $\text{Precision}(r)$ is the precision at recall level $r$. The precision is calculated as:
$\text{Precision}=\frac{TP}{TP+FP}$
and recall is:
$\text{Recall}=\frac{TP}{TP+FN}$
where $TP$ is true positives, $FP$ is false positives, and $FN$ is false negatives.
MCC (Matthews Correlation Coefficient).
The Matthews Correlation Coefficient (MCC) is calculated as:
$\text{MCC}=\frac{TP\cdot TN-FP\cdot FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}$
where $TP$ is true positives, $TN$ is true negatives, $FP$ is false positives, and $FN$ is false negatives.
TNR (True Negative Rate).
The True Negative Rate (TNR) is defined as:
$\text{TNR}=\frac{TN}{TN+FP}$
where $TN$ is true negatives and $FP$ is false positives.
TPR (True Positive Rate).
The True Positive Rate (TPR), also known as recall, is given by:
$\text{TPR}=\frac{TP}{TP+FN}$
where $TP$ is true positives and $FN$ is false negatives.
AUC (Area Under the Curve).
The Area Under the Curve (AUC) is the area under the Receiver Operating Characteristic (ROC) curve. It can be calculated as:
[…]
where $\text{TPR}(FPR)$ is the true positive rate at a given false positive rate ($FPR$).
ACC (Accuracy).
Accuracy is calculated as the ratio of correctly classified instances to the total number of instances:
$\text{ACC}=\frac{TP+TN}{TP+TN+FP+FN}$
where $TP$ is true positives, $TN$ is true negatives, $FP$ is false positives, and $FN$ is false negatives.
F1 (F1 Score).
The F1 score is the harmonic mean of precision and recall, given by:
[…]
where precision and recall are defined as:
$\text{Precision}=\frac{TP}{TP+FP},\quad\text{Recall}=\frac{TP}{TP+FN}$
IOU (Intersection over Union).
Intersection over Union (IoU) is calculated as the ratio of the intersection of predicted and ground truth regions to their union:
$\text{IoU}=\frac{|A\cap B|}{|A\cup B|}$
where $A$ is the predicted region and $B$ is the ground truth region.
Appendix E Details of AIGC and Document Benchmarks
E.1 Training Details for AIGC Benchmark Implementation
We resize images to 224×224 and apply only basic data augmentations, including flipping, brightness and contrast adjustment, compression, and Gaussian blur. All models are trained for 20 epochs using a cosine decay learning rate schedule, decreasing from $41e-$ 4 to $51e-$ 5. We use the ImageNet split of DiffusionForensics as the training set. For Synthbuster, we use a fully connected layer as the classifier.
E.2 Training Details for Document Image Manipulation Localization Benchmark Models
For all models, we adopt a cosine learning rate schedule decaying from $1\mathrm{e}{-4}$ to $5\mathrm{e}{-7}$, using the AdamW optimizer with $\beta_{1}{=}0.9$, $\beta_{2}{=}0.999$, weight decay of 0.05, and gradient accumulation step of 1.
Epoch schedules are adapted to dataset size and complexity: 10 epochs for Doctamper, 75 epochs for RTM, and 150 epochs for all other datasets. We use a batch size of 10 for CATFB, 8 for DTD and FFDN, and 4 for TIFDM.
Table 11: Cross-dataset AUC evaluation on Deepfake benchmarks.
Model FF++c40 FF-DF FF-F2F FF-NT FF-FS CDFv1 CDFv2 DFD DFDC DFDCP FaceShifter UADFV Average
ResNet 0.675 0.938 0.936 0.879 0.935 0.692 0.684 0.834 0.678 0.763 0.613 0.770 0.783
Xception 0.728 0.976 0.967 0.925 0.971 0.708 0.719 0.870 0.702 0.764 0.605 0.885 0.818
EfficientNet 0.504 0.503 0.505 0.508 0.500 0.541 0.535 0.517 0.510 0.454 0.478 0.480 0.503
Swin-Transformer 0.771 0.992 0.988 0.969 0.988 0.704 0.746 0.901 0.754 0.820 0.655 0.721 0.834
SegFormer 0.739 0.993 0.990 0.962 0.993 0.770 0.799 0.885 0.726 0.771 0.719 0.889 0.853
ConvNeXt 0.794 0.996 0.991 0.981 0.994 0.718 0.784 0.911 0.747 0.842 0.756 0.805 0.860
CapsuleNet 0.613 0.855 0.867 0.839 0.800 0.527 0.660 0.699 0.640 0.656 0.624 0.761 0.712
Recce 0.634 0.815 0.843 0.762 0.848 0.572 0.602 0.727 0.625 0.672 0.531 0.816 0.704
Spsl 0.730 0.982 0.966 0.933 0.974 0.690 0.726 0.876 0.706 0.758 0.632 0.878 0.821
MVSS-Net 0.713 0.992 0.981 0.953 0.986 0.655 0.700 0.857 0.707 0.727 0.696 0.737 0.809
Trufor 0.642 0.968 0.967 0.931 0.960 0.683 0.698 0.832 0.670 0.769 0.629 0.746 0.791
IML-ViT 0.750 0.973 0.963 0.906 0.973 0.763 0.726 0.851 0.708 0.663 0.560 0.702 0.795
Mesorch 0.767 0.991 0.983 0.956 0.983 0.803 0.814 0.867 0.760 0.807 0.665 0.889 0.857
DualNet 0.637 0.784 0.685 0.707 0.606 0.471 0.552 0.540 0.547 0.559 0.667 0.530 0.607
HifiNet 0.587 0.727 0.664 0.646 0.598 0.456 0.611 0.648 0.607 0.619 0.544 0.745 0.621
UnivFD 0.690 0.931 0.645 0.549 0.871 0.660 0.671 0.798 0.633 0.674 0.656 0.891 0.722
DTD 0.498 0.509 0.522 0.513 0.507 0.576 0.520 0.490 0.485 0.499 0.484 0.499 0.508
FFDN 0.714 0.997 0.991 0.980 0.997 0.706 0.699 0.871 0.712 0.769 0.741 0.771 0.829
Table 12: Cross-dataset AUC evaluation on Image Manipulation Detection and Localization (IMDL) datasets.
Model COVERAGE Columbia IMD2020 NIST16 Cocoglide Autosplice Average
ResNet 0.492 0.380 0.549 0.540 0.629 0.764 0.559
Xception 0.491 0.465 0.537 0.582 0.666 0.756 0.583
EfficientNet 0.489 0.623 0.506 0.496 0.508 0.483 0.517
Swin-Transformer 0.505 0.636 0.631 0.361 0.816 0.864 0.635
SegFormer 0.501 0.558 0.594 0.502 0.785 0.819 0.627
ConvNeXt 0.489 0.625 0.598 0.529 0.761 0.825 0.638
CapsuleNet 0.504 0.330 0.527 0.486 0.704 0.745 0.549
Recce 0.485 0.506 0.492 0.593 0.616 0.642 0.556
Spsl 0.493 0.419 0.545 0.557 0.675 0.759 0.575
MVSS-Net 0.481 0.298 0.539 0.527 0.683 0.795 0.554
Trufor 0.492 0.306 0.564 0.368 0.710 0.808 0.541
IML-ViT 0.498 0.483 0.556 0.484 0.657 0.819 0.583
Mesorch 0.495 0.285 0.570 0.494 0.653 0.773 0.545
DualNet 0.503 0.268 0.517 0.314 0.619 0.748 0.495
HifiNet 0.500 0.745 0.534 0.503 0.589 0.677 0.591
UnivFD 0.499 0.886 0.786 0.715 0.706 0.785 0.729
DTD 0.500 0.679 0.498 0.468 0.510 0.506 0.527
FFDN 0.522 0.553 0.624 0.574 0.681 0.927 0.647
Table 13: Cross-domain AUC evaluation on AIGC datasets.
Model DiffusionForensics GenImage_all GenImage_ADM GenImage_Midjourney GenImage_wukong GenImage_glide GenImage_BigGAN GenImage_sd14 GenImage_sd15 GenImage_VQDM Average
ResNet 0.614 0.865 0.649 0.818 0.928 0.936 0.971 0.919 0.920 0.761 0.838
Xception 0.757 0.980 0.990 0.949 0.979 0.986 0.997 0.987 0.985 0.963 0.957
EfficientNet 0.544 0.597 0.842 0.485 0.642 0.523 0.542 0.619 0.619 0.489 0.590
Swin-Transformer 0.915 0.999 1.000 0.998 0.999 1.000 1.000 1.000 1.000 1.000 0.991
SegFormer 0.847 0.998 1.000 0.995 0.997 0.999 1.000 0.999 0.999 0.999 0.983
ConvNeXt 0.895 1.000 1.000 0.999 1.000 1.000 1.000 1.000 1.000 1.000 0.989
CapsuleNet 0.546 0.971 0.992 0.921 0.971 0.978 0.992 0.984 0.982 0.944 0.928
Recce 0.684 0.906 0.909 0.856 0.927 0.907 0.948 0.933 0.936 0.819 0.882
Spsl 0.770 0.987 0.994 0.965 0.985 0.990 0.999 0.992 0.991 0.979 0.965
MVSS-Net 0.671 0.994 0.999 0.978 0.991 0.996 1.000 0.996 0.996 0.996 0.962
Trufor 0.726 0.996 0.998 0.990 0.992 0.997 0.998 0.997 0.997 0.997 0.969
IML-ViT 0.627 0.991 0.999 0.960 0.988 0.989 0.999 0.996 0.997 0.997 0.954
Mesorch 0.629 0.996 0.999 0.987 0.995 0.998 0.999 0.998 0.998 0.997 0.960
DualNet 0.899 0.988 0.999 0.926 0.992 0.993 0.999 0.997 0.996 0.998 0.979
HifiNet 0.575 0.756 0.707 0.618 0.867 0.830 0.751 0.820 0.812 0.618 0.735
UnivFD 0.742 0.813 0.767 0.725 0.829 0.852 0.883 0.841 0.839 0.759 0.805
DTD 0.457 0.499 0.542 0.547 0.510 0.531 0.459 0.518 0.505 0.379 0.495
FFDN 0.999 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
Table 14: Cross-dataset AUC evaluation on document manipulation datasets.
Model OSTF RTM T-SROIE Tampered-IC13 Average
ResNet 0.704 0.689 0.953 0.807 0.788
Xception 0.762 0.734 0.966 0.877 0.835
EfficientNet 0.581 0.512 0.884 0.653 0.657
Swin-Transformer 0.856 0.758 0.990 0.966 0.893
SegFormer 0.856 0.705 0.980 0.956 0.874
ConvNeXt 0.849 0.762 0.994 0.957 0.890
CapsuleNet 0.704 0.670 0.946 0.759 0.770
Recce 0.688 0.555 0.542 0.797 0.645
Spsl 0.769 0.738 0.972 0.888 0.842
MVSS-Net 0.806 0.741 0.978 0.913 0.860
Trufor 0.805 0.732 0.979 0.920 0.859
IML-ViT 0.800 0.703 0.972 0.923 0.850
Mesorch 0.819 0.739 0.982 0.954 0.873
DualNet 0.658 0.657 0.935 0.776 0.756
HifiNet 0.663 0.615 0.937 0.762 0.744
UnivFD 0.684 0.569 0.938 0.740 0.733
DTD 0.595 0.496 0.748 0.658 0.624
FFDN 0.893 0.782 0.997 0.975 0.912
Appendix F Details of IFF-Protocol
Implementation Resolution.
We use the commonly used 256×256 resolution for detection tasks, such as Deepfake and AIGC. However, UnivFD uses CLIP-ViT as the backbone, which only supports 224×224 image input. Therefore, the input image is resized to 224×224 for UnivFD. On the other hand, the SoTA for Document is specifically designed for 512×512 resolution, with some models like FFDN even having a fixed input size of 512×512. Therefore, we resize images to 512×512 for Document models.
Results on Domains.
We provide the test results of backbones and domain-specific SoTAs under the IFF-Protocol for each domain, which are Table 11 for Deepfake, Table 12 for IMDL, Table 13 for AIGC, and Table 14 for Document.
Appendix G Details of Experiments
Table 15: Extractor & backbone performance difference in IMDL region
Extractor Model Coverage Columbia IMD2020 NIST16 Cocoglide Autosplice Mean Diff
Sobel [5] ResNet 0.018 0.141 -0.046 0.084 -0.048 -0.093 0.009
Xception -0.005 -0.001 0.015 -0.010 -0.052 -0.069 -0.020
EfficientNet 0.011 -0.045 -0.018 0.060 0.015 0.012 0.006
Swin -0.006 -0.192 -0.118 0.256 -0.196 -0.247 -0.084
SegFormer -0.018 -0.007 -0.071 0.091 -0.191 -0.189 -0.064
ConvNeXt 0.012 -0.019 -0.101 0.023 -0.146 -0.181 -0.069
Bayar [3] ResNet -0.007 0.027 -0.025 0.064 -0.044 -0.086 -0.012
Xception -0.009 -0.049 -0.001 -0.026 -0.007 -0.044 -0.023
EfficientNet 0.019 0.004 0.007 0.006 0.002 0.013 0.009
Swin 0.004 -0.282 -0.128 0.214 -0.156 -0.215 -0.094
SegFormer 0.009 -0.108 -0.069 0.073 -0.173 -0.173 -0.073
ConvNeXt 0.005 -0.234 -0.080 0.028 -0.126 -0.141 -0.091
FPH [43] ResNet -0.002 0.169 -0.044 -0.016 -0.079 -0.105 -0.013
Xception 0.007 -0.016 -0.011 0.013 -0.102 -0.035 -0.024
EfficientNet 0.017 -0.062 -0.002 0.095 0.057 0.115 0.037
Swin -0.010 -0.208 -0.115 0.235 -0.248 -0.204 -0.092
SegFormer -0.007 -0.104 -0.066 0.077 -0.263 -0.197 -0.093
ConvNeXt 0.014 0.203 -0.097 -0.044 -0.272 -0.313 -0.085
DCT [1] ResNet -0.002 0.073 -0.005 0.023 0.002 -0.030 0.010
Xception 0.015 0.025 -0.002 0.001 0.009 -0.008 0.007
EfficientNet 0.007 0.034 -0.007 0.018 -0.005 0.005 0.009
Swin -0.023 -0.162 -0.066 0.206 -0.189 -0.105 -0.057
SegFormer -0.003 -0.156 -0.021 0.067 -0.140 -0.043 -0.049
ConvNeXt -0.002 -0.187 -0.078 0.038 -0.216 -0.145 -0.098
Table 16: Extractor & backbone performance difference in AIGC region
extractor model DiffusionForensics_test GenImage_all GenImage_ADM GenImage_Midjourney GenImage_wukong GenImage_glide GenImage_BigGAN GenImage_sd14 GenImage_sd15 GenImage_VQDM mean_diff
Sobel [5] ResNet -0.037 -0.046 0.312 -0.112 -0.064 -0.073 -0.186 -0.060 -0.061 -0.119 -0.045
Xception -0.133 -0.042 -0.007 -0.090 -0.031 -0.020 -0.007 -0.035 -0.033 -0.115 -0.051
EfficientNet 0.049 0.031 0.015 0.029 -0.023 0.128 0.105 -0.023 -0.036 0.084 0.036
Swin -0.206 -0.086 -0.014 -0.161 -0.073 -0.035 -0.036 -0.081 -0.079 -0.212 -0.098
SegFormer -0.181 -0.095 -0.025 -0.185 -0.085 -0.055 -0.056 -0.083 -0.079 -0.203 -0.105
ConvNeXt -0.212 -0.129 -0.027 -0.278 -0.117 -0.069 -0.004 -0.124 -0.124 -0.290 -0.137
Bayar [3] ResNet -0.027 -0.060 0.007 -0.054 -0.050 -0.029 -0.165 -0.042 -0.039 -0.123 -0.058
Xception -0.054 -0.023 -0.018 -0.034 -0.019 -0.009 -0.006 -0.019 -0.019 -0.057 -0.026
EfficientNet 0.013 -0.013 -0.096 0.008 -0.001 0.001 -0.033 0.003 0.004 0.009 -0.011
Swin -0.240 -0.052 -0.008 -0.117 -0.058 -0.038 -0.004 -0.054 -0.056 -0.082 -0.071
SegFormer -0.215 -0.054 -0.005 -0.113 -0.057 -0.043 -0.010 -0.049 -0.050 -0.109 -0.070
ConvNeXt -0.227 -0.050 -0.015 -0.123 -0.055 -0.036 -0.007 -0.051 -0.053 -0.055 -0.067
FPH [43] ResNet -0.108 -0.132 -0.094 -0.247 -0.060 -0.130 -0.195 -0.082 -0.094 -0.172 -0.131
Xception -0.198 -0.094 -0.033 -0.186 -0.044 -0.084 -0.097 -0.049 -0.048 -0.227 -0.106
EfficientNet 0.031 0.153 0.051 0.118 0.163 0.319 0.315 0.124 0.124 0.027 0.142
Swin -0.264 -0.073 -0.006 -0.153 -0.035 -0.073 -0.073 -0.035 -0.036 -0.192 -0.094
SegFormer -0.217 -0.087 -0.018 -0.201 -0.048 -0.088 -0.081 -0.045 -0.046 -0.187 -0.102
ConvNeXt -0.452 -0.511 -0.547 -0.507 -0.399 -0.539 -0.581 -0.452 -0.456 -0.627 -0.507
DCT [1] ResNet -0.036 -0.015 0.011 -0.016 -0.002 -0.020 -0.036 -0.010 -0.014 -0.035 -0.017
Xception -0.007 0.000 0.000 0.003 $0$0.001 -0.001 0.000 0.001 -0.004 -0.001
EfficientNet 0.006 0.008 -0.017 0.009 0.013 0.004 -0.009 0.021 0.016 0.030 0.008
Swin -0.253 -0.019 -0.002 -0.042 -0.013 -0.019 -0.011 -0.008 -0.011 -0.051 -0.043
SegFormer -0.223 -0.012 -0.004 -0.029 -0.014 -0.011 -0.006 -0.007 -0.008 -0.025 -0.034
ConvNeXt -0.234 -0.051 -0.012 -0.109 -0.031 -0.062 -0.045 -0.018 -0.019 -0.123 -0.070
Table 17: Extractor & backbone performance difference in Deepfake region
Extractor Model FF++c40 FF-DF FF-F2F FF-NT FF-FS CDFv1 CDFv2 DFD DFDC DFDCP FaceShifter UADFV Mean
Sobel [5] ResNet [19] -0.052 -0.111 -0.166 -0.150 -0.212 -0.097 -0.042 -0.144 -0.076 -0.063 -0.046 0.095 -0.089
Xception [7] 0.022 -0.021 -0.034 -0.056 -0.040 0.033 0.034 -0.078 -0.009 -0.023 -0.004 -0.020 -0.016
EfficientNet [55] 0.022 0.044 0.034 0.035 0.017 -0.031 -0.017 -0.006 0.043 0.050 0.069 0.068 0.027
Swin-Transformer [30] -0.130 -0.133 -0.197 -0.232 -0.255 0.002 -0.084 -0.234 -0.102 -0.142 0.001 -0.077 -0.132
SegFormer [64] -0.040 -0.117 -0.174 -0.224 -0.180 -0.034 -0.145 -0.176 -0.089 -0.111 -0.104 -0.120 -0.126
ConvNeXt [31] -0.141 -0.120 -0.169 -0.222 -0.232 -0.066 -0.167 -0.228 -0.115 -0.198 -0.095 -0.235 -0.166
Bayar [3] ResNet [19] -0.006 -0.071 -0.081 -0.097 -0.175 -0.100 -0.042 -0.122 -0.044 -0.088 -0.119 0.046 -0.075
Xception [7] 0.006 -0.007 -0.013 -0.021 -0.011 0.006 -0.007 -0.023 -0.015 -0.025 -0.008 -0.005 -0.010
EfficientNet [55] $0$0.011 0.001 0.000 0.002 0.064 0.030 -0.003 0.024 -0.018 0.012 0.004 0.009
Swin-Transformer [30] -0.050 -0.092 -0.134 -0.185 -0.213 0.056 -0.032 -0.107 -0.097 -0.168 -0.128 -0.006 -0.096
SegFormer [64] 0.023 -0.069 -0.119 -0.154 -0.109 -0.040 -0.123 -0.077 -0.077 -0.084 -0.177 -0.162 -0.097
ConvNeXt [31] -0.019 -0.062 -0.109 -0.151 -0.126 0.028 -0.052 -0.113 -0.085 -0.103 -0.200 -0.061 -0.088
FPH [43] ResNet [19] -0.019 -0.123 -0.305 -0.268 -0.367 -0.105 -0.108 -0.214 -0.094 -0.074 0.013 0.008 -0.138
Xception [7] -0.010 -0.111 -0.106 -0.173 -0.137 0.094 -0.052 -0.211 -0.065 -0.206 -0.141 -0.085 -0.100
EfficientNet [55] 0.054 0.123 0.031 0.052 0.017 -0.057 0.044 0.078 0.045 0.143 0.104 0.081 0.060
Swin-Transformer [30] 0.037 -0.066 -0.117 -0.252 -0.093 0.078 -0.008 -0.161 -0.072 -0.068 -0.104 0.124 -0.059
SegFormer [64] 0.062 -0.108 -0.201 -0.330 -0.164 -0.067 -0.153 -0.271 -0.089 -0.159 -0.111 -0.094 -0.140
ConvNeXt -0.282 -0.482 -0.498 -0.468 -0.473 -0.150 -0.368 -0.380 -0.225 -0.296 -0.237 -0.272 -0.344
DCT [1] ResNet [19] 0.014 -0.006 -0.008 -0.013 -0.011 0.052 0.029 -0.029 -0.001 0.011 0.007 0.021 0.005
Xception [7] -0.006 0.004 -0.003 0.004 -0.001 -0.020 0.006 -0.011 0.005 -0.023 0.003 -0.032 -0.006
EfficientNet [55] -0.004 -0.004 0.005 0.004 0.008 0.035 0.023 -0.006 0.036 0.011 0.020 0.001 0.011
Swin-Transformer [30] 0.040 -0.027 -0.030 -0.105 -0.027 0.092 -0.021 -0.060 0.001 -0.045 -0.071 0.159 -0.008
SegFormer [64] 0.043 -0.019 -0.028 -0.041 -0.026 0.039 -0.037 -0.039 0.007 0.007 -0.116 0.026 -0.015
ConvNeXt [31] 0.041 -0.049 -0.087 -0.173 -0.076 -0.102 -0.127 -0.158 -0.025 -0.135 -0.288 -0.013 -0.099
Table 18: Extractor & backbone performance difference in Document region
extractor model OSTF_test RealTextManipulation_test T-SROIE_test Tampered-IC13_test mean_diff
Sobel [5] ResNet -0.044 -0.046 -0.020 -0.070 -0.045
Xception -0.046 -0.019 -0.012 -0.030 -0.027
EfficientNet -0.025 0.033 0.019 -0.002 0.006
Swin -0.143 -0.051 -0.033 -0.153 -0.095
SegFormer -0.135 -0.008 -0.027 -0.156 -0.082
ConvNeXt -0.160 -0.090 -0.041 -0.200 -0.123
Bayar [3] ResNet -0.023 -0.050 -0.013 -0.047 -0.033
Xception 0.001 -0.015 -0.010 0.001 -0.006
EfficientNet -0.001 0.003 0.007 0.001 0.003
Swin -0.119 -0.040 -0.040 -0.105 -0.076
SegFormer -0.123 0.002 -0.033 -0.107 -0.065
ConvNeXt -0.145 -0.047 -0.036 -0.126 -0.088
FPH [43] ResNet -0.053 -0.071 -0.009 -0.077 -0.052
Xception -0.079 -0.045 -0.012 -0.122 -0.065
EfficientNet 0.049 0.023 -0.053 0.044 0.016
Swin -0.113 -0.034 -0.038 -0.091 -0.069
SegFormer -0.202 0.014 -0.022 -0.176 -0.097
ConvNeXt -0.236 -0.209 -0.072 -0.289 -0.202
DCT [1] ResNet -0.008 -0.005 0.004 -0.026 -0.009
Xception -0.001 0.001 -0.002 0.002 0.000
EfficientNet -0.013 0.028 0.007 0.003 0.006
Swin -0.093 -0.028 -0.020 -0.056 -0.049
SegFormer -0.063 0.018 -0.011 -0.029 -0.021
ConvNeXt -0.132 -0.050 -0.039 -0.107 -0.082
G.1 Details of Feature Extractors
BayarConv.
BayarConv [3] is a constrained convolutional layer, that is able to jointly suppress an image’s content and adaptively learn manipulation detection features. It learns to extract noise artifacts within images.
Sobel.
Sobel layer is proposed to enhance edge-related patterns, whereas the subtle boundary cues are critical for manipulation detection and localization [5]. This is based on the common assumption that manipulations often leave edge artifacts along the tampered boundaries.
DCT.
DCT (Discrete Cosine Transform) [1] is a mathematical technique that transforms spatial domain data into frequency domain components, primarily used to isolate image features based on their frequency to extract frequency features.
FPH.
FPH [43] is designed to find out tampering clues in the frequency domain with DCT coefficients. It receives DCT coefficients and a quantization table as input, and outputs a 256-channel feature map that is downsampled by a factor of 8. This design enables it to effectively capture compression artifacts and frequency-domain inconsistencies for downstream analysis.
G.2 Details for Extractor & Backbone in different tasks
We provide the performance differences of 6 backbones with and without 4 different feature extractors across the 4 domains. The table presents detailed results for each individual test dataset. They are Table 15 for IMDL, Table 16 for AIGC, Tabel 17 for Deepfake, and Table 18 for Document.
Appendix H Computational Resources
The experiments were conducted on three different servers. The first server is equipped with two AMD EPYC 7542 CPUs, 256GB RAM, and 6$\times$×NVIDIA A40 GPUs, which was used for all IFF-related experiments. The remaining experiments were performed on two servers: one with a single AMD EPYC 7542 CPU, 256GB RAM, and 4$\times$×NVIDIA RTX 3090 GPUs, and another with two AMD EPYC 7542 CPUs, 256GB RAM, and 8$\times$×NVIDIA RTX 3090 GPUs.
Appendix I Broader Impacts Discussion
ForensicHub establishes a critical benchmark for all-domain fake image detection and localization, helping to curb the spread of falsified images in society and significantly advancing the development of a more trustworthy digital environment. However, the comprehensive coverage of detection methods in ForensicHub may enable malicious actors to study and develop targeted evasion techniques.

### v2

@@
-
+Bo Du1† Xuekang Zhu1,2† Xiaochen Ma3† Chenfan Qu4, 2† Kaiwen Feng1† Zhe Yang1 Chi-Man Pun5 Jian Liu2‡ Jizhe Zhou1‡
 Abstract
@@
 However, the research efforts of FIDL have gradually split into four relatively independent research domains over time.
-1) Deepfake detection (Deepfake) [37, 26, 25, 8, 38, 4, 66, 42, 28, 33, 54, 71, 49, 65, 67]: detects manipulations such as face swapping, expression editing, or feature replacement.
-2) Image manipulation detection/localization (IMDL) [5, 17, 29, 35, 70, 53]: detects and localizes the tampering in natural images.
-3) AI-Generated Image Detection (AIGC) [40, 18, 63, 59]: detects the images fully generated by deep generative models such as Stable Diffusion [46].
-4) Document Image Manipulation Localization (Document)  [48, 43, 6, 12, 52]: localizes the forgery of various forms of document images, including receipts, certificates, and identification materials, with a particular focus on detecting modifications to the printed text.
+1) Deepfake detection (Deepfake) nguyen2019capsule ; li2018uadfv ; li2020xray ; dang2020ffd ; ni2022core ; cao2022recce ; yan2023ucf ; han2023fcd ; qian2020f3net ; liu2021spsl ; luo2021srm ; sun2022SIA ; zhuang2022uia ; shiohara2022sbi ; yan2024LSDA ; yan2023deepfakebench : detects human-centric manipulations such as face swapping, expression editing, or feature replacement.
+2) Image manipulation detection/localization (IMDL) liu2020d ; MVSS_2021 ; trufor2023 ; liu2022pscc ; NCL_IML_2023 ; ma2023iml ; yu2024diffforensics ; zhu2025mesoscopic ; sparseViT_2025 : detects and localizes the tampering in natural images.
+3) AI-Generated Image Detection (AIGC) univfd ; HiFi-Net2023 ; dualnet ; wang2023dire ; zhang2025towards : detects the images fully generated by deep generative models such as Stable Diffusion stablediffusion .
+4) Document Image Manipulation Localization (Document)  psnet ; qu2023doctamper ; chen2024ffdn ; dong2024tifdm ; song2025caftb ; yu2023learning ; li2026document : localizes the forgery of various forms of document images, including receipts, certificates, and identification materials, with a particular focus on detecting modifications to the printed text.
 Although these domains have become isolated due to differences in application scenarios, manipulation types, and detection methods, there are still overlaps and similarities among them.
-As vision tasks, these four domains almost universally adopt SoTA detection or segmentatoin models as pre-trained backbones.
+As vision tasks, these four domains almost universally adopt SoTA detection or segmentation models as pre-trained backbones.
 Further, since the creators of fake images typically aim to preserve semantically plausible and realistic content, all four domains have placed considerable emphasis on designing low-level visual feature extractors to capture subtle, non-semantic discrepancies for reliable detection.
@@
 Task Model Backbone Artifact Strategy Output Type Contribution
-Deepfake Capsule-Net [37](ICASSP19) VGG [50] Dynamic Routing Label Proposes a capsule network with dynamic routing and a VGG19 backbone.
-RECCE [4](CVPR22) Xception [7] Reconstruction Label Proposes a graph-based framework leveraging reconstruction differences
-SPSL [28](CVPR21) Xception [7] Phase Spectrum Label Proposes phase-spectrum fusion with Xception for face forgery detection.
-UCF [66](ICCV23) Xception [7] Multi-task Disentanglement Label Proposes multi-task disentanglement with Xception for deepfake generalization.
-SBI [49](CVPR22) EfficientNet [55] Frequency,Blending Boundaries Label Proposes self-blended images to improve deepfake detection generalization.
-IMDL MVSS-Net [62](ICCV21) Resnet [19] BayarConv,Sobel Label,Mask Exploit noise and boundary artifacts via multi-view learning for manipulation detection.
-CAT-Net [22](IJCV22) HRNet [56] DCT Mask Fuse RGB and DCT streams to learn compression artifacts for splice localization.
-PSCC-Net [29](TCSVT22) HRNet [56] Multi-Resolution Conv Label,Mask Progressively refine masks with spatio-channel correlations for high-reso localization.
-Trufor [17](CVPR23) Seformer [64] High-Reso,Multi-scale,Edge Label,Mask Fuse RGB and learned noise fingerprints to detect manipulations as anomalies.
-IML-ViT [35](Arxiv) ViT [13] BayarConv,SRM Filter Mask Use ViT with high-reso, multi-scale edge-aware design for manipulation localization.
-Mesorch [70](AAAI25) Conv. [31],Segfor. [64] DCT Mask Fuse micro and macro cues for mesoscopic image manipulation localization.
-AIGC Dire [59](ICCV23) Resnet [19] Diffusion Reconstruction Label Use reconstruction error of diffusion for diffusion-generated images detection.
-DualNet [63](APSIPA23) CNN SRM,Low Frequency Label Fuse SRM residual and low-frequency content streams for AIGC detection.
-HiFiNet [18](CVPR23) HRNet [56] Multi-branch Feature Extractor Label,Mask Learn hierarchical fine-grained representations of forgery attributes.
-Synthbuster [2](OJSP23) None Fourier Transform Label Leverage spectral artifacts in the frequency domain for diffusion detection.
-UnivFD [40](CVPR23) CLIP-ViT [45] None Label Use pretrained vision-language model features for unified detection.
-Document CAFTB [52](TOMM24) Resnet [19] SRM Mask Proposes CAFTB-Net with dual-branch and cross-attention.
-TIFDM [12](TCE24) Resnet [19] None Mask Proposes a robust network with multiscale attention.
-DTD [43](CVPR23) Conv. [31], Swin.[30] Frequency Mask Proposes DTD with frequency head and multi-view decoder.
-FFDN [6](ECCV24) ConvNext [31] Wavelet, Frequency Mask Proposes FFDN combining visual enhancement and frequency decomposition
-Although individual benchmarks exist in some domains, such as DeepfakeBench [67] for Deepfake and IMDLBenCo [36] for IMDL, a unified benchmark for all domains in FIDL remains blank.
+Deepfake Capsule-Net nguyen2019capsule (ICASSP19) VGG VGG_2015 Dynamic Routing Label Proposes a capsule network with dynamic routing and a VGG19 backbone.
+RECCE cao2022recce (CVPR22) Xception chollet2017xception Reconstruction Label Proposes a graph-based framework leveraging reconstruction differences
+SPSL liu2021spsl (CVPR21) Xception chollet2017xception Phase Spectrum Label Proposes phase-spectrum fusion with Xception for face forgery detection.
+UCF yan2023ucf (ICCV23) Xception chollet2017xception Multi-task Disentanglement Label Proposes multi-task disentanglement with Xception for deepfake generalization.
+SBI shiohara2022sbi (CVPR22) EfficientNet tan2019efficientnet Frequency,Blending Boundaries Label Proposes self-blended images to improve deepfake detection generalization.
+IMDL MVSS-Net Mantra_2019 (ICCV21) Resnet Resnet_2016 BayarConv,Sobel Label,Mask Exploit noise and boundary artifacts via multi-view learning for manipulation detection.
+CAT-Net CAT-Net2022 (IJCV22) HRNet HRNet DCT Mask Fuse RGB and DCT streams to learn compression artifacts for splice localization.
+PSCC-Net liu2022pscc (TCSVT22) HRNet HRNet Multi-Resolution Conv Label,Mask Progressively refine masks with spatio-channel correlations for high-reso localization.
+Trufor trufor2023 (CVPR23) Seformer SegFormer_2021 High-Reso,Multi-scale,Edge Label,Mask Fuse RGB and learned noise fingerprints to detect manipulations as anomalies.
+IML-ViT ma2023iml (Arxiv) ViT ViT_2021 BayarConv,SRM Filter Mask Use ViT with high-reso, multi-scale edge-aware design for manipulation localization.
+Mesorch zhu2025mesoscopic (AAAI25) Conv. Convnet_2022 ,Segfor. SegFormer_2021 DCT Mask Fuse micro and macro cues for mesoscopic image manipulation localization.
+AIGC Dire wang2023dire (ICCV23) Resnet Resnet_2016 Diffusion Reconstruction Label Use reconstruction error of diffusion for diffusion-generated images detection.
+DualNet dualnet (APSIPA23) CNN SRM,Low Frequency Label Fuse SRM residual and low-frequency content streams for AIGC detection.
+HiFiNet HiFi-Net2023 (CVPR23) HRNet HRNet Multi-branch Feature Extractor Label,Mask Learn hierarchical fine-grained representations of forgery attributes.
+Synthbuster bammey2023synthbuster (OJSP23) None Fourier Transform Label Leverage spectral artifacts in the frequency domain for diffusion detection.
+UnivFD univfd (CVPR23) CLIP-ViT clip None Label Use pretrained vision-language model features for unified detection.
+Document CAFTB song2025caftb (TOMM24) Resnet Resnet_2016 SRM Mask Proposes CAFTB-Net with dual-branch and cross-attention.
+TIFDM dong2024tifdm (TCE24) Resnet Resnet_2016 None Mask Proposes a robust network with multiscale attention.
+DTD qu2023doctamper (CVPR23) Conv. Convnet_2022 , Swin.Swin_2021 Frequency Mask Proposes DTD with frequency head and multi-view decoder.
+FFDN chen2024ffdn (ECCV24) ConvNext Convnet_2022 Wavelet, Frequency Mask Proposes FFDN combining visual enhancement and frequency decomposition
+Although individual benchmarks exist in some domains, such as DeepfakeBench yan2023deepfakebench for Deepfake and IMDLBenCo ma2024imdl for IMDL, a unified benchmark for all domains in FIDL remains blank.
 The absence of such a unified benchmark results in significant domain silos, where each domain independently constructs its datasets, models, and evaluation protocols without interoperability.
 Domain silos lead to redundant and uneven research across existing FIDL fields, and difficulty in establishing a general and unified FIDL approach, severely hindering the development of the entire FIDL field.
+Besides, in real-world scenarios, it is often impossible to predetermine the type of manipulation (deepfake, imdl, aigc and document) present in an image, making unified detection particularly important for users.
 Therefore, establishing a unified benchmark for all domains is critically significant.
@@
 2 Related Works
-Fake image detection and localization span four sub-tasks, summarized in Appendix C.
+Fake image detection and localization encompass four sub-tasks: 1) Deepfake Detection, 2) Image Manipulation Detection and Localization, 3) AI-Generated Image Detection, and 4) Document Image Manipulation Localization.
+The characteristics of each task are summarized in Appendix C.
 Despite rapid progress, a unified benchmark is lacking—each task uses isolated pipelines, limiting cross-task comparison.
-DeepfakeBench [67] is a Deepfake detection benchmark specifically designed to address the lack of uniformity in data processing pipelines, leading to inconsistent data inputs for detection models.
-IMDLBenCo [36] is a benchmark and codebase for IMDL, aiming to compare IMDL models through a unified training and evaluation protocol.
-AIGCDetectBenchmark [68] is a repository for experiments on 9 AI-generated image detection methods.
+Despite the rapid development of these tasks, there is a lack of a unified benchmark, with some task having its isolated benchmark, creating barriers between them.
+DeepfakeBench yan2023deepfakebench is a Deepfake detection benchmark specifically designed to address the lack of uniformity in data processing pipelines, leading to inconsistent data inputs for detection models.
+IMDLBenCo ma2024imdl is a benchmark and codebase for IMDL, aiming to compare IMDL models through a unified training and evaluation protocol.
+AIGCDetectBenchmark EkkoAIGCDetectBenchmark is a repository for experiments on 9 AI-generated image detection methods.
 These benchmarks provide models, datasets, and evaluation metrics within their respective tasks, but their underlying designs lack cross-task considerations, making them difficult to integrate across different detection scenarios.
@@
 To accommodate different forensic tasks, ForensicHub is designed as a modular architecture consisting of four main components: Datasets, Transforms, Models, and Evaluators.
-Datasets handle the data loading process and are required to return fields that conform to the ForensicHub specification.
-Transforms handle the data pre-processing and augmentation for different tasks.
-Models, through alignment with Datasets and unified output, allow for the inclusion of various state-of-the-art image forensic models.
-Evaluators cover commonly used image- and pixel-level metrics for different tasks, and are implemented with GPU acceleration to improve evaluation efficiency during training and testing.
+1) Datasets handle the data loading process and are required to return fields that conform to the ForensicHub specification.
+2) Transforms handle the data pre-processing and augmentation for different tasks.
+3) Models, through alignment with Datasets and unified output, allow for the inclusion of various state-of-the-art image forensic models.
+4) Evaluators cover commonly used image- and pixel-level metrics for different tasks, and are implemented with GPU acceleration to improve evaluation efficiency during training and testing.
 Configurable workflow.
@@
 Construction of ForensicHub.
-To enable broad interoperability and reduce duplication of effort, ForensicHub adopts an adapter-based design [14] that ensures seamless integration with DeepfakeBench [67] and IMDLBenCo [36], two widely used benchmarks.
+To enable broad interoperability and reduce duplication of effort, ForensicHub adopts an adapter-based design gamma1995design that ensures seamless integration with DeepfakeBench yan2023deepfakebench and IMDLBenCo ma2024imdl , two widely used benchmarks.
 This mechanism allows users to reuse existing models and datasets without major modification, while also supporting the definition of new models and benchmarks within ForensicHub under the unified protocol.
@@
 In total, ForensicHub covers 4 tasks, 23 datasets, 42 models, 6 backbones, and implements 11 commonly used image- and pixel-level metrics.
-Datasets used in this paper are: Faceforencs++ [47], Celeb-DF [27], DFD [9], FaceShifter [23] and UADFV [26] for Deepfake, CASIA [11], COVERAGE [60], Columbia [20], IMD2020 [39], NIST16 [16], CocoGlide [17], and Autosplice [21] for IMDL, DiffusionForensics [59], GenImage [69] for AIGC, Doctamper [43], T-SROIE [58], OSTF [44], TPIC-13 [57], RTM [32] for Doc.
+Datasets used in this paper are: FaceForensics++ rossler2019faceforensics++ , Celeb-DF li2020celeb , DFD dolhansky2020dfd , FaceShifter li2019faceshifter and UADFV li2018uadfv for Deepfake; CASIA CASIA_2013 , COVERAGE Coverage_2016 , Columbia Columbia_2006 , IMD2020 IMD20_2020 , NIST16 NIST16_2019 , CocoGlide trufor2023 , and Autosplice jia2023autosplice for IMDL; DiffusionForensics wang2023dire , GenImage zhu2023genimage for AIGC; Doctamper qu2023doctamper , T-SROIE wang2022tsroie , OSTF qu2025ostf , TPIC-13 wang2022tpic , RTM luo2025rtm for Doc.
 A brief summary of each dataset is provided in Table 2, with more details in Appendix D.1.
@@
 Task Dataset Year Content Real Fake Annotation Pipeline Split
-Deepfake FaceForensics++ [47] 2019 Face 45,388 127,209 Label Automatic Train,Test
-Celeb-DF-v1 [27] 2020 Face 1,203 1,933 Label Automatic Train,Test
-Celeb-DF-v2 [27] 2020 Face 5,620 10,800 Label Automatic Train,Test
-DeepFakeDetection [15] 2019 Face 10,741 91,800 Label Automatic Train,Test
-DFDC [9] 2020 Face 63,265 68,851 Label Automatic Train,Test
-DFDCP [10] 2019 Face 5,901 11,321 Label Automatic Train,Test
-FaceShifter [24] 2020 Face 4,479 4,479 Label Automatic Train,Test
-UADFV [26] 2018 Face 1,548 1,551 Label Automatic Train,Test
-IMDL CASIAv1 [11] 2013 General 0 920 Label,Mask Manual Train
-CASIAv2 [11] 2013 General 7,491 5,123 Label,Mask Manual Train
-COVERAGE [60] 2016 General 100 100 Label,Mask Manual Train
-Columbia [20] 2006 General 183 180 Label,Mask Automatic Train
-IMD2020 [11] 2020 General 414 2,010 Label,Mask Manual Train
-NIST16 [16] 2019 General 873 564 Label,Mask Manual Train
-CocoGlide [17] 2023 General 512 512 Label,Mask Automatic Train
-Autosplice [21] 2023 General 2,273 3,621 Label,Mask Manual Train
-AIGC DiffusionForensics [59] 2023 General 134,000 481,200 Label Automatic Train,Val,Test
-GenImage [69] 2023 General 1,331,167 1,350,000 Label Automatic Train,Test
-Document Doctamper [43] 2023 Document 0 170,000 Mask Automatic Train,Test
-OSTF [44] 2025 Document 16,405 3,685 Mask Manual Train,Test
-RealTextManipulation [32] 2025 Document 52,314 17,423 Mask Automatic+Manual Train,Test
-T-SROIE [58] 2022 Document 21,268 4,326 Mask Manual Train,Test
-Tampered-IC13 [57] 2022 Document 2,810 1,228 Mask Manual Train,Test
-Models used in this paper are: Capsule-Net [37], RECCE [4], SPSL [28], UCF [66], and SBI [49] for Deepfake detection, MVSS-Net [5], CAT-Net [22], PSCC-Net [29], Trufor [17], IML-ViT [35], and Mesorch [70] for image manipulation and localization, Dire [59], DualNet [63], HiFiNet [18], Synthbuster [2], and UnivFD [40] for AIGC detection, DTD [43], FFDN [6], CAFTB [52], TIFDM [12] for document detection.
+Deepfake FaceForensics++ rossler2019faceforensics++ 2019 Face 45,388 127,209 Label Automatic Train,Test
+Celeb-DF-v1 li2020celeb 2020 Face 1,203 1,933 Label Automatic Train,Test
+Celeb-DF-v2 li2020celeb 2020 Face 5,620 10,800 Label Automatic Train,Test
+DeepFakeDetection dfd2019 2019 Face 10,741 91,800 Label Automatic Train,Test
+DFDC dolhansky2020dfd 2020 Face 63,265 68,851 Label Automatic Train,Test
+DFDCP dolhansky2019deepfake 2019 Face 5,901 11,321 Label Automatic Train,Test
+FaceShifter li2020advancing 2020 Face 4,479 4,479 Label Automatic Train,Test
+UADFV li2018uadfv 2018 Face 1,548 1,551 Label Automatic Train,Test
+IMDL CASIAv1 CASIA_2013 2013 General 0 920 Label,Mask Manual Train
+CASIAv2 CASIA_2013 2013 General 7,491 5,123 Label,Mask Manual Train
+COVERAGE Coverage_2016 2016 General 100 100 Label,Mask Manual Train
+Columbia Columbia_2006 2006 General 183 180 Label,Mask Automatic Train
+IMD2020 CASIA_2013 2020 General 414 2,010 Label,Mask Manual Train
+NIST16 NIST16_2019 2019 General 873 564 Label,Mask Manual Train
+CocoGlide trufor2023 2023 General 512 512 Label,Mask Automatic Train
+Autosplice jia2023autosplice 2023 General 2,273 3,621 Label,Mask Manual Train
+AIGC DiffusionForensics wang2023dire 2023 General 134,000 481,200 Label Automatic Train,Val,Test
+GenImage zhu2023genimage 2023 General 1,331,167 1,350,000 Label Automatic Train,Test
+Document Doctamper qu2023doctamper 2023 Document 0 170,000 Mask Automatic Train,Test
+OSTF qu2025ostf 2025 Document 16,405 3,685 Mask Manual Train,Test
+RealTextManipulation luo2025rtm 2025 Document 52,314 17,423 Mask Automatic+Manual Train,Test
+T-SROIE wang2022tsroie 2022 Document 21,268 4,326 Mask Manual Train,Test
+Tampered-IC13 wang2022tpic 2022 Document 2,810 1,228 Mask Manual Train,Test
+Models used in this paper are: Capsule-Net nguyen2019capsule , RECCE cao2022recce , SPSL liu2021spsl , UCF yan2023ucf , and SBI shiohara2022sbi for Deepfake detection; MVSS-Net MVSS_2021 , CAT-Net CAT-Net2022 , PSCC-Net liu2022pscc , Trufor trufor2023 , IML-ViT ma2023iml , and Mesorch zhu2025mesoscopic for image manipulation and localization; Dire wang2023dire , DualNet dualnet , HiFiNet HiFi-Net2023 , Synthbuster bammey2023synthbuster , and UnivFD univfd for AIGC detection; DTD qu2023doctamper , FFDN chen2024ffdn , CAFTB song2025caftb , TIFDM dong2024tifdm for document detection.
 These methods are from official repositories and our reimplementations.
-In addition, ForensicHub also selects 6 commonly used backbones in visual tasks, which are: Resnet [19], Xception [7], EfficientNet [55], Segformer [64], Swin Transformer [30], and ConvNext [31].
+In addition, ForensicHub also selects 6 commonly used backbones in visual tasks, which are: Resnet Resnet_2016 , Xception chollet2017xception , EfficientNet tan2019efficientnet , Segformer SegFormer_2021 , Swin Transformer Swin_2021 , and ConvNext Convnet_2022 .
 Details about models can be found in Appendix D.2.
@@
 4 Benchmarks
-In addition to being fully compatible with existing benchmarks, DeepfakeBench [67] and IMDLBenCo [36], ForensicHub further extends standardization efforts by introducing unified evaluation protocols for the AIGC and Document domains—two areas that previously lacked widely accepted benchmarks and codebases.
+In addition to being fully compatible with existing benchmarks, DeepfakeBench yan2023deepfakebench and IMDLBenCo ma2024imdl , ForensicHub further extends standardization efforts by introducing unified evaluation protocols for the AIGC and Document domains—two areas that previously lacked widely accepted benchmarks and codebases.
 We propose two protocols for two domains to evaluate generalization.
@@
 In the field of AIGC detection, the challenge in dataset construction usually lies not in obtaining a sufficient quantity of samples, since they can be easily generated using existing models, but in ensuring comprehensive coverage of a wide range of generative models.
-Therefore, we select only two commonly used public datasets: DiffusionForensics [59] and GenImage [69].
+Therefore, we select only two commonly used public datasets: DiffusionForensics wang2023dire and GenImage zhu2023genimage .
 The former contains only diffusion-based generated images, while the latter covers a million-scale dataset constructed from eight SoTA generative models.
-Models are trained on DiffusionForensics and evaluated on different generative models within GenImage to assess generalization, as detection methods typically already achieve good performance on samples from the same generative model [69].
+Models are trained on DiffusionForensics and evaluated on different generative models within GenImage to assess generalization, as detection methods typically already achieve good performance on samples from the same generative model zhu2023genimage .
 The detailed data splits are summarized in Table D.1.3.
 Models.
-ForensicHub implements five SoTA methods in AIGC detection: Dire [59], DualNet [63], HiFiNet [18], Synthbuster [2], and UnivFD [40], among which Synthbuster has no official open-source code and is fully reimplemented by us.
+ForensicHub implements five SoTA methods in AIGC detection: Dire wang2023dire , DualNet dualnet , HiFiNet HiFi-Net2023 , Synthbuster bammey2023synthbuster , and UnivFD univfd , among which Synthbuster has no official open-source code and is fully reimplemented by us.
 More details about models and training settings can be found in Appendix E.1.
 Results.
-Table 3 in green background presents the AUC scores for image-level classification for AIGC benchmark, divided into in-domain results on the test set split of DiffusionForensics [59], and cross-domain results on different generative models and the total set on GenImage [69].
+Table 3 in green background presents the AUC scores for image-level classification for AIGC benchmark, divided into in-domain results on the test set split of DiffusionForensics wang2023dire , and cross-domain results on different generative models and the total set on GenImage zhu2023genimage .
 The results show that AIGC SoTAs generally achieve excellent performance on the DiffusionForensics test set, which shares the same source as the training set, and also perform well on datasets composed of diffusion-based generated images like ADM, VQDM, and GLIDE that are similar to the training data.
@@
 DiffusionForensics ADM BigGAN Midjourney VQDM GLIDE SD V1.4 SD V1.5 Wukong AverageC
-DualNet [63](APSIPA23) 1.0000 0.9986 0.9172 0.8495 0.9901 0.9835 0.8127 0.8081 0.7925 0.8902
-HiFiNet [18](CVPR23) 1.0000 0.9998 0.8407 0.7210 0.9999 0.9991 0.6765 0.6747 0.6675 0.8160
-Synthbuster(OJSP23) 0.6662 0.6226 0.8550 0.4565 0.6308 0.7719 0.4606 0.4614 0.3902 0.5762
-UnivFD(CVPR23) 0.9947 0.8400 0.9687 0.5427 0.9663 0.9541 0.7078 0.7072 0.6781 0.7920
-MVSS-Net(ICCV21) 0.9992 0.9706 0.9520 0.6242 0.8894 0.9380 0.7207 0.7192 0.6089 0.7997
-CAT-Net(IJCV22) 0.9985 0.9391 0.9327 0.6423 0.8226 0.7987 0.7171 0.7142 0.6156 0.7707
-Trufor(CVPR23) 0.9999 0.9750 0.9316 0.7079 0.9340 0.9144 0.8683 0.8687 0.8018 0.8752
-IML-ViT(Arxiv) 1.0000 0.9594 0.9152 0.8086 0.9577 0.9464 0.8924 0.8877 0.8006 0.8957
-Mesorch(AAAI25) 0.9998 0.9754 0.9667 0.7011 0.9331 0.9253 0.8168 0.8179 0.7441 0.8582
+DualNet dualnet (APSIPA23) 1.0000 0.9986 0.9172 0.8495 0.9901 0.9835 0.8127 0.8081 0.7925 0.8902
+HiFiNet HiFi-Net2023 (CVPR23) 1.0000 0.9998 0.8407 0.7210 0.9999 0.9991 0.6765 0.6747 0.6675 0.8160
+Synthbuster bammey2023synthbuster (OJSP23) 0.6662 0.6226 0.8550 0.4565 0.6308 0.7719 0.4606 0.4614 0.3902 0.5762
+UnivFD univfd (CVPR23) 0.9947 0.8400 0.9687 0.5427 0.9663 0.9541 0.7078 0.7072 0.6781 0.7920
+MVSS-Net MVSS_2021 (ICCV21) 0.9992 0.9706 0.9520 0.6242 0.8894 0.9380 0.7207 0.7192 0.6089 0.7997
+CAT-Net CAT-Net2022 (IJCV22) 0.9985 0.9391 0.9327 0.6423 0.8226 0.7987 0.7171 0.7142 0.6156 0.7707
+Trufor trufor2023 (CVPR23) 0.9999 0.9750 0.9316 0.7079 0.9340 0.9144 0.8683 0.8687 0.8018 0.8752
+IML-ViT ma2023iml (Arxiv) 1.0000 0.9594 0.9152 0.8086 0.9577 0.9464 0.8924 0.8877 0.8006 0.8957
+Mesorch zhu2025mesoscopic (AAAI25) 0.9998 0.9754 0.9667 0.7011 0.9331 0.9253 0.8168 0.8179 0.7441 0.8582
 Dataset.
-Existing datasets for document image manipulation localization can be broadly categorized into two types: high-fidelity non-sliced datasets, including T-SROIE [58], OSTF [44], TPIC-13 [57], and RTM [32]; and sliced datasets, represented by Doctamper [43].
+Existing datasets for document image manipulation localization can be broadly categorized into two types: high-fidelity non-sliced datasets, including T-SROIE wang2022tsroie , OSTF qu2025ostf , TPIC-13 wang2022tpic , and RTM luo2025rtm ; and sliced datasets, represented by Doctamper qu2023doctamper .
 The primary difference lies in whether the images are preprocessed using patch-wise slicing.
@@
 Model.
-We employ two open-source models, DTD [43] and FFDN [6], and reproduce two closed-source models, CATFB [52] and TIFDM [12]. All details are provided in Appendix D.2.4 and E.2.
+We employ two open-source models, DTD qu2023doctamper and FFDN chen2024ffdn , and reproduce two closed-source models, CATFB song2025caftb and TIFDM dong2024tifdm .
+All details are provided in Appendix D.2.4 and E.2.
 Results.
-Following the original protocols [43, 6], each detector is trained on its designated training split and evaluated on the corresponding test split.
+Following the original protocols qu2023doctamper ; chen2024ffdn , each detector is trained on its designated training split and evaluated on the corresponding test split.
 As shown in Table 4, three models consistently achieve top performance: FFDN and DTD, both designed specifically for document forensics, and Cat-Net, an IMDL-based model.
 Notably, all three methods incorporate JPEG-specific priors, such as DCT coefficients and quantization tables, highlighting the discriminative value of compression artifacts for manipulation localization in document images.
-Table 4: Binary-F1 scores of document detectors. AverageD is the mean over three Doctamper test dataset, and AverageAll is averaged across all seven test datasets.
+Table 4: Binary‑F1 scores of document detectors. AverageD is the mean over three Doctamper test dataset, and AverageAll is averaged across all seven test datasets.
 Model Doctamper T-SROIE-train OSTF-train TPIC-13-train RTM-train AverageAll
 DocTamperFCD DocTamperSCD DocTamperTest AverageD T-SROIE-test OSTF-test TPIC-13-test RTM-test
-CAFTB [52](TOMM24) 0.2917 0.3770 0.3275 0.3321 0.9165 0.6478 0.8394 0.2493 0.5213
-DTD [43](CVPR23) 0.6856 0.7392 0.8031 0.7426 0.9205 0.5626 0.8341 0.1720 0.6739
-FFDN [6](ECCV24) 0.8773 0.7392 0.8212 0.8126 0.9137 0.5586 0.8186 0.1589 0.6982
-TIFDM [12](TCE24) 0.0896 0.2572 0.2585 0.2018 0.8942 0.5410 0.7972 0.0591 0.4138
-MVSS-Net [5](ICCV21) 0.2066 0.3710 0.3810 0.3195 0.8756 0.5254 0.7864 0.1314 0.4682
-PSCC-Net [29](TCSVT22) 0.3855 0.3931 0.4972 0.4253 0.9305 0.5697 0.7894 0.1971 0.5375
-Cat-Net [22](IJCV22) 0.7600 0.6405 0.7644 0.7216 0.8726 0.5371 0.7947 0.1174 0.6410
-IML-ViT [35](Arxiv) 0.4688 0.5117 0.4486 0.4764 0.8731 0.5128 0.7202 0.0920 0.5182
-Trufor [17](CVPR23) 0.2613 0.3124 0.2517 0.2751 0.9095 0.6485 0.8197 0.2302 0.4905
-Mesorch [70](AAAI25) 0.4231 0.4586 0.4651 0.4489 0.9049 0.4804 0.7607 0.1314 0.5177
+CAFTB song2025caftb (TOMM24) 0.2917 0.3770 0.3275 0.3321 0.9165 0.6478 0.8394 0.2493 0.5213
+DTD qu2023doctamper (CVPR23) 0.6856 0.7392 0.8031 0.7426 0.9205 0.5626 0.8341 0.1720 0.6739
+FFDN chen2024ffdn (ECCV24) 0.8773 0.7392 0.8212 0.8126 0.9137 0.5586 0.8186 0.1589 0.6982
+TIFDM dong2024tifdm (TCE24) 0.0896 0.2572 0.2585 0.2018 0.8942 0.5410 0.7972 0.0591 0.4138
+MVSS-Net MVSS_2021 (ICCV21) 0.2066 0.3710 0.3810 0.3195 0.8756 0.5254 0.7864 0.1314 0.4682
+PSCC-Net liu2022pscc (TCSVT22) 0.3855 0.3931 0.4972 0.4253 0.9305 0.5697 0.7894 0.1971 0.5375
+Cat-Net CAT-Net2022 (IJCV22) 0.7600 0.6405 0.7644 0.7216 0.8726 0.5371 0.7947 0.1174 0.6410
+IML-ViT ma2023iml (Arxiv) 0.4688 0.5117 0.4486 0.4764 0.8731 0.5128 0.7202 0.0920 0.5182
+Trufor trufor2023 (CVPR23) 0.2613 0.3124 0.2517 0.2751 0.9095 0.6485 0.8197 0.2302 0.4905
+Mesorch zhu2025mesoscopic (AAAI25) 0.4231 0.4586 0.4651 0.4489 0.9049 0.4804 0.7607 0.1314 0.5177
 However, this evaluation setting has a key limitation: all models are trained and tested within the same distribution, limiting the assessment of cross-domain generalization.
@@
 DocTamperFCD DocTamperSCD DocTamperTest AverageW T-SROIE OSTF TPIC-13 RTM AverageC
-CAFTB [52](TOMM24) 0.2917 0.3770 0.3275 0.3321 0.2617 0.1194 0.3007 0.0328 0.1787 0.2444
-DTD [43](CVPR23) 0.6856 0.7392 0.8031 0.7426 0.5245 0.1241 0.2835 0.0575 0.2474 0.4596
-FFDN [6](ECCV24) 0.8773 0.7392 0.8212 0.8126 0.5330 0.2409 0.3572 0.0708 0.3005 0.5199
-TIFDM [12](TCE24) 0.0896 0.2572 0.2585 0.2018 0.0582 0.0058 0.0134 0.0176 0.0238 0.1000
-MVSS-Net [5](ICCV23) 0.2066 0.3710 0.3810 0.3195 0.1870 0.0373 0.1134 0.0268 0.0911 0.1890
-PSCC-Net [29](TCSVT22) 0.3855 0.3931 0.4972 0.4253 0.5168 0.4414 0.5495 0.1255 0.4083 0.4156
-Cat-Net [22](IJCV22) 0.7600 0.6405 0.7644 0.7216 0.6085 0.1777 0.3430 0.0630 0.2981 0.4796
-IML-ViT [35](Arxiv) 0.4688 0.5117 0.4486 0.4764 0.4269 0.2101 0.2563 0.0764 0.2424 0.3427
-Trufor [17](CVPR23) 0.2613 0.3124 0.2517 0.2751 0.2126 0.0464 0.1038 0.0342 0.0993 0.1746
-Mesorch [70](AAAI25) 0.4231 0.4586 0.4651 0.4489 0.2937 0.1388 0.2408 0.0405 0.1785 0.2944
+CAFTB song2025caftb (TOMM24) 0.2917 0.3770 0.3275 0.3321 0.2617 0.1194 0.3007 0.0328 0.1787 0.2444
+DTD qu2023doctamper (CVPR23) 0.6856 0.7392 0.8031 0.7426 0.5245 0.1241 0.2835 0.0575 0.2474 0.4596
+FFDN chen2024ffdn (ECCV24) 0.8773 0.7392 0.8212 0.8126 0.5330 0.2409 0.3572 0.0708 0.3005 0.5199
+TIFDM dong2024tifdm (TCE24) 0.0896 0.2572 0.2585 0.2018 0.0582 0.0058 0.0134 0.0176 0.0238 0.1000
+MVSS-Net MVSS_2021 (ICCV23) 0.2066 0.3710 0.3810 0.3195 0.1870 0.0373 0.1134 0.0268 0.0911 0.1890
+PSCC-Net liu2022pscc (TCSVT22) 0.3855 0.3931 0.4972 0.4253 0.5168 0.4414 0.5495 0.1255 0.4083 0.4156
+Cat-Net CAT-Net2022 (IJCV22) 0.7600 0.6405 0.7644 0.7216 0.6085 0.1777 0.3430 0.0630 0.2981 0.4796
+IML-ViT ma2023iml (Arxiv) 0.4688 0.5117 0.4486 0.4764 0.4269 0.2101 0.2563 0.0764 0.2424 0.3427
+Trufor trufor2023 (CVPR23) 0.2613 0.3124 0.2517 0.2751 0.2126 0.0464 0.1038 0.0342 0.0993 0.1746
+Mesorch zhu2025mesoscopic (AAAI25) 0.4231 0.4586 0.4651 0.4489 0.2937 0.1388 0.2408 0.0405 0.1785 0.2944
 5 Image Forensic Fusion Protocol
@@
 The IFF-Protocol defines the training set as a combination of Deepfake, IMDL, AIGC, and Document data, where each training epoch samples an equal number of instances from each domain at random.
-During training, we select FaceForensics++ [47] from Deepfake, CASIAv2 [11] from IMDL, GenImage [69] from AIGC, and OSTF [44], RealTextManipulation, T-SROIE, and Tampered-IC13 from the Document.
+During training, we select FaceForensics++ rossler2019faceforensics++ from Deepfake, CASIAv2 CASIA_2013 from IMDL, GenImage zhu2023genimage from AIGC, and OSTF qu2025ostf , RealTextManipulation luo2025rtm , T-SROIE wang2022tsroie , and Tampered-IC13 wang2022tpic from the Document.
 We use the smallest dataset, CASIAv2 with 12,641 samples, as the sampling number for each epoch.
@@
 We resize images to 256×256 (except UnivFD, DTD and FFDN, see Appendix F for details) and apply only basic data augmentations, including flipping, brightness and contrast adjustment, compression, and Gaussian blur.
-All models are trained for 20 epochs using a cosine decay learning rate schedule, decreasing from $41e-$ 4 to $51e-$ 5.
+All models are trained for 20 epochs using a cosine decay learning rate schedule, decreasing from 1e41e-4 to 1e51e-5.
 For models that output masks (IMDL and Doc), we apply max pooling to the final-layer feature maps to obtain a predicted label and compute the loss using only the label.
@@
 Model Params (M) FLOPs (G) Model Params (M) FLOPs (G) Model Params (M) FLOPs (G)
-Resnet [19] $44.55$ $20.59$ Capsule-Net [37] $3.90$ $32.43$ Mesorch [70] $85.75$ $57.95$
-Xception [7] $22.86$ $12.05$ RECCE [4] $25.83$ $16.18$ DualNet [63] $7.99$ $66.34$
-EfficientNet [55] $19.34$ $4.14$ SPSL [28] $20.81$ $12.06$ HiFiNet [18] $6.89$ $145.00$
-ConvNeXt [31] $50.22$ $22.74$ MVSS-Net [5] $147.00$ $83.08$ UnivFD [40] $428.00$ $156.00$
-Segformer [64] $44.07$ $16.29$ Trufor [17] $68.70$ $112.00$ DTD [43] $67.07$ $272.00$
-Swin [30] $49.61$ $28.65$ IML-ViT [35] $91.78$ $80.37$ FFDN [6] $89.20$ $453.00$
+Resnet Resnet_2016 $44.55$ $20.59$ Capsule-Net nguyen2019capsule $3.90$ $32.43$ Mesorch zhu2025mesoscopic $85.75$ $57.95$
+Xception chollet2017xception $22.86$ $12.05$ RECCE cao2022recce $25.83$ $16.18$ DualNet dualnet $7.99$ $66.34$
+EfficientNet tan2019efficientnet $19.34$ $4.14$ SPSL liu2021spsl $20.81$ $12.06$ HiFiNet HiFi-Net2023 $6.89$ $145.00$
+ConvNeXt Convnet_2022 $50.22$ $22.74$ MVSS-Net MVSS_2021 $147.00$ $83.08$ UnivFD univfd $428.00$ $156.00$
+Segformer SegFormer_2021 $44.07$ $16.29$ Trufor trufor2023 $68.70$ $112.00$ DTD qu2023doctamper $67.07$ $272.00$
+Swin Swin_2021 $49.61$ $28.65$ IML-ViT ma2023iml $91.78$ $80.37$ FFDN chen2024ffdn $89.20$ $453.00$
 Model Efficiency.
@@
 Method Deepfake IMDL AIGC Document Average
-FF-c40 CDFv2 DFD Columbia IMD2020 Autosplice DF GenImage T-SROIE OSTF RTM
-Resnet [19](CVPR16) 0.681 0.730 0.793 0.482 0.533 0.738 0.619 0.797 0.951 0.681 0.662 0.697
-Xception [7](CVPR17) 0.728 0.719 0.870 0.465 0.537 0.756 0.757 0.980 0.966 0.762 0.734 0.752
-EfficientNet [55](ICML19) 0.504 0.535 0.517 0.623 0.506 0.483 0.544 0.597 0.884 0.581 0.512 0.571
-Segformer [64](NIPS21) 0.691 0.748 0.862 0.409 0.562 0.824 0.805 0.998 0.980 0.866 0.736 0.771
-Swin [30](ICCV21) 0.771 0.746 0.901 0.636 0.631 0.864 0.915 0.999 0.990 0.856 0.758 0.824
-ConvNext [31](CVPR22) 0.794 0.784 0.911 0.625 0.598 0.825 0.895 1.000 0.994 0.849 0.762 0.822
-Capsule-Net [37](ICASSP19) 0.613 0.660 0.699 0.330 0.527 0.745 0.546 0.971 0.946 0.704 0.670 0.674
-RECCE [4](CVPR22) 0.634 0.602 0.727 0.506 0.492 0.642 0.684 0.906 0.542 0.688 0.555 0.634
-SPSL [28](CVPR21) 0.730 0.726 0.876 0.419 0.545 0.759 0.770 0.987 0.972 0.769 0.738 0.754
-MVSS-Net [5](ICCV21) 0.713 0.700 0.857 0.298 0.539 0.795 0.671 0.994 0.978 0.806 0.741 0.736
-Trufor [17](CVPR23) 0.642 0.698 0.832 0.306 0.564 0.808 0.726 0.996 0.979 0.805 0.732 0.735
-IML-ViT [35](Arxiv) 0.750 0.726 0.851 0.483 0.556 0.819 0.627 0.991 0.972 0.800 0.703 0.753
-Mesorch [70](AAAI25) 0.767 0.814 0.867 0.285 0.570 0.773 0.629 0.996 0.982 0.819 0.739 0.749
-DualNet [63](APSIPA23) 0.637 0.552 0.540 0.268 0.517 0.748 0.899 0.988 0.935 0.658 0.657 0.673
-HiFiNet [18](CVPR23) 0.587 0.611 0.648 0.745 0.534 0.677 0.575 0.756 0.937 0.663 0.615 0.668
-UnivFD [40](CVPR23) 0.690 0.671 0.798 0.886 0.786 0.785 0.742 0.813 0.938 0.684 0.569 0.760
-DTD [43](CVPR23) 0.498 0.520 0.490 0.679 0.498 0.506 0.457 0.499 0.748 0.595 0.496 0.544
-FFDN [6](ECCV24) 0.714 0.699 0.871 0.553 0.624 0.927 0.999 1.000 0.997 0.893 0.782 0.824
+$F$-c40 CDFv2 DFD Columbia IMD2020 Autosplice DF GenImage T-SROIE OSTF RTM
+Resnet Resnet_2016 (CVPR16) 0.681 0.730 0.793 0.482 0.533 0.738 0.619 0.797 0.951 0.681 0.662 0.697
+Xception chollet2017xception (CVPR17) 0.728 0.719 0.870 0.465 0.537 0.756 0.757 0.980 0.966 0.762 0.734 0.752
+EfficientNet tan2019efficientnet (ICML19) 0.504 0.535 0.517 0.623 0.506 0.483 0.544 0.597 0.884 0.581 0.512 0.571
+Segformer SegFormer_2021 (NIPS21) 0.691 0.748 0.862 0.409 0.562 0.824 0.805 0.998 0.980 0.866 0.736 0.771
+Swin Swin_2021 (ICCV21) 0.771 0.746 0.901 0.636 0.631 0.864 0.915 0.999 0.990 0.856 0.758 0.824
+ConvNext Convnet_2022 (CVPR22) 0.794 0.784 0.911 0.625 0.598 0.825 0.895 1.000 0.994 0.849 0.762 0.822
+Capsule-Net nguyen2019capsule (ICASSP19) 0.613 0.660 0.699 0.330 0.527 0.745 0.546 0.971 0.946 0.704 0.670 0.674
+RECCE cao2022recce (CVPR22) 0.634 0.602 0.727 0.506 0.492 0.642 0.684 0.906 0.542 0.688 0.555 0.634
+SPSL liu2021spsl (CVPR21) 0.730 0.726 0.876 0.419 0.545 0.759 0.770 0.987 0.972 0.769 0.738 0.754
+Sia sia (ECCV22) 0.629 0.584 0.671 0.653 0.483 0.626 0.593 0.748 0.610 0.677 0.574 0.622
+Effort effort (ICML25) 0.805 0.846 0.930 0.979 0.861 0.943 0.930 0.992 0.960 0.834 0.732 0.892
+MVSS-Net MVSS_2021 (ICCV21) 0.713 0.700 0.857 0.298 0.539 0.795 0.671 0.994 0.978 0.806 0.741 0.736
+Trufor trufor2023 (CVPR23) 0.642 0.698 0.832 0.306 0.564 0.808 0.726 0.996 0.979 0.805 0.732 0.735
+IML-ViT ma2023iml (Arxiv) 0.750 0.726 0.851 0.483 0.556 0.819 0.627 0.991 0.972 0.800 0.703 0.753
+Mesorch zhu2025mesoscopic (AAAI25) 0.767 0.814 0.867 0.285 0.570 0.773 0.629 0.996 0.982 0.819 0.739 0.749
+DualNet dualnet (APSIPA23) 0.637 0.552 0.540 0.268 0.517 0.748 0.899 0.988 0.935 0.658 0.657 0.673
+HiFiNet HiFi-Net2023 (CVPR23) 0.587 0.611 0.648 0.745 0.534 0.677 0.575 0.756 0.937 0.663 0.615 0.668
+UnivFD univfd (CVPR23) 0.690 0.671 0.798 0.886 0.786 0.785 0.742 0.813 0.938 0.684 0.569 0.760
+FatFormer fatformer (CVPR24) 0.842 0.770 0.866 0.199 0.585 0.784 0.941 0.999 0.983 0.806 0.751 0.758
+CO-SPY co-spy (CVPR25) 0.819 0.780 0.875 0.460 0.716 0.779 0.940 0.989 0.969 0.836 0.748 0.829
+DTD qu2023doctamper (CVPR23) 0.498 0.520 0.490 0.679 0.498 0.506 0.457 0.499 0.748 0.595 0.496 0.544
+FFDN chen2024ffdn (ECCV24) 0.714 0.699 0.871 0.553 0.624 0.927 0.999 1.000 0.997 0.893 0.782 0.824
 Benchmark Result.
-Table 7 shows the AUC scores of backbones and domain-specific SoTA methods on datasets from four domains under the IFF-Protocol, in which DFD refers to DeepFakeDetection [15], DF refers to DiffusionForensics [59], and RTM refers to RealTextManipulation [32].
+Table 7 shows the AUC scores of backbones and domain-specific SoTA methods on datasets from four domains under the IFF-Protocol, in which DFD refers to DeepFakeDetection dfd2019 , DF refers to DiffusionForensics wang2023dire , and RTM refers to RealTextManipulation luo2025rtm .
 We provide detailed results in Appendix F.
-The results show that surprisingly, visual backbones such as ConvNeXt [31] and Swin Transformer [30] outperform almost all domain-specific SoTA methods, indicating that backbones demonstrate greater potential when trained on more unified fake images.
+The results show that surprisingly, visual backbones such as ConvNeXt Convnet_2022 and Swin Transformer Swin_2021 outperform almost all domain-specific SoTA methods, indicating that backbones demonstrate greater potential when trained on more unified fake images.
 Meanwhile, domain-specific SoTAs do not necessarily retain their superiority within their own tasks.
-For instance, UnivFD [40], a CLIP-based fine-tuned model for AIGC detection, demonstrates strong performance on the IMD2020 [39] from IMDL, revealing valuable insights into the transferability of cross-task methods.
+For instance, UnivFD univfd , a CLIP-based fine-tuned model for AIGC detection, demonstrates strong performance on the IMD2020 IMD20_2020 from IMDL, revealing valuable insights into the transferability of cross-task methods.
 From a task perspective, although IMDL target shifts from pixel-level to image-level classification, it remains challenging due to significant distribution differences across datasets in terms of size and manipulation types.
@@
 Since each domain has proposed specific feature extractors, to explore their effectiveness under a unified domain, we conduct experiments using 6 backbones combined with 4 different extractors in the shallow layer under the aforementioned IFF-Protocol setting.
-The extractors are BayarConv [3] for noise aritfacts, Sobel [5] for edge arifacts, DCT [1] and FPH [43] for frequency artifacts.
+The extractors are BayarConv Bayar_2018 for noise aritfacts, Sobel MVSS_2021 for edge arifacts, DCT ahmed2006discrete and FPH qu2023doctamper for frequency artifacts.
 Details of each extractor can be found in Appendix G.1.
 Results in Table 8 show the AUC differences between versions of each backbone using the four different feature extractors and those without, averaged across all test datasets for each task.
-All backbones except EfficientNet [55] show performance drops after using feature extractors, indicating that under the IFF-Protocol, where training data includes sufficient manipulation types and image quantity, models do not rely on the additional information provided by feature extractors.
+All backbones except EfficientNet tan2019efficientnet show performance drops after using feature extractors, indicating that under the IFF-Protocol, where training data includes sufficient manipulation types and image quantity, models do not rely on the additional information provided by feature extractors.
 However, due to its lightweight nature, EfficientNet still benefits from the use of feature extractors.
@@
 Table 8: Mean AUC differences between extractor-enhanced models and their plain counterparts. Positive values (red) indicate gains; negative (blue) indicate drops.
-Task Extractor = Sobel [5] Extractor = Bayar[3]
+Task Extractor = Sobel MVSS_2021 Extractor = BayarBayar_2018
 ResNet Xception EfficientNet SwinTransformer SegFormer ConvNext ResNet Xception EfficientNet SwinTransformer SegFormer ConvNext
@@
 IMDL 0.0093 -0.0203 0.0058 -0.0838 -0.0642 -0.0687 -0.0118 -0.0227 0.0085 -0.0938 -0.0735 -0.0913
-Extractor = FPH [43] Extractor = DCT [1]
+Extractor = FPH qu2023doctamper Extractor = DCT ahmed2006discrete
 ResNet Xception EfficientNet SwinTransformer SegFormer ConvNext ResNet Xception EfficientNet SwinTransformer SegFormer ConvNext
@@
 6.2 Transferability of Task-Specific Detectors
-6.2.1 Cross-Evaluation Between IMDL and Document Benchmarks
+6.2.1 Cross‑Evaluation Between IMDL and Document Benchmarks
 Current Document-level detectors’ input–output formats are fully compatible with those of Image Manipulation Detection and Localization models.
 Leveraging this consistency, we perform a bidirectional evaluation: IMDL detectors are tested on the Document benchmark, and Document detectors are tested on the IMDL benchmark.
-This cross-testing enlarges the effective model pool for each benchmark and allows us to probe detector generality beyond their original task scopes.
+This cross‑testing enlarges the effective model pool for each benchmark and allows us to probe detector generality beyond their original task scopes.
 IMDL → Document.
-Table 4 reports the within-domain results obtained under the original Document benchmark split, whereas Table 5 presents the cross-domain scores produced by our newly introduced generalization protocol.
+Table 4 reports the within‑domain results obtained under the original Document benchmark split, whereas Table 5 presents the cross‑domain scores produced by our newly introduced generalization protocol.
 Across both settings, IMDL detectors demonstrate strong competitiveness in the document forensics scenario.
-In the conventional split, the Cat-Net [22, 43, 6] family achieves the best average F1, confirming the merit of its hierarchical “cat-net” paradigm.
-Under the more challenging cross-domain evaluation, PSCC-Net [29] displays markedly better generalization, suggesting that progressive spatial modeling captures cues for document manipulation localization.
+In the conventional split, the Cat-Net CAT-Net2022 ; qu2023doctamper ; chen2024ffdn family achieves the best average F1, confirming the merit of its hierarchical “cat‑net” paradigm.
+Under the more challenging cross‑domain evaluation, PSCC-Net liu2022pscc displays markedly better generalization, suggesting that progressive spatial modeling captures cues for document manipulation localization.
 We expect future work to further investigate the underlying mechanisms behind PSCC-Net.
 Document → IMDL.
-Following the MVSS training protocol [36], all Document-oriented models are trained on the CASIAv2 dataset [11] and evaluated on five standard IMDL test sets.
-As shown in Table 9, the dual-branch architecture of CAFTB [52] achieves the best overall performance among Document models when transferred to IMDL tasks—an outcome that aligns with the design philosophy of the current SoTA model Mesorch [70], which also emphasizes dual-branch learning.
+Following the MVSS training protocol ma2024imdl , all Document-oriented models are trained on the CASIAv2 dataset CASIA_2013 and evaluated on five standard IMDL test sets.
+As shown in Table 9, the dual-branch architecture of CAFTB song2025caftb achieves the best overall performance among Document models when transferred to IMDL tasks—an outcome that aligns with the design philosophy of the current SoTA model Mesorch zhu2025mesoscopic , which also emphasizes dual-branch learning.
 Table 9: Pixel-level binary F1 evaluation on IMDL benchmarks for document-trained detectors.
 Model CASIAv1 COVERAGE Columbia IMD2020 NIST16 Average
-CAFTB [52](TOMM24) 0.6234 0.2557 0.6731 0.3526 0.3770 0.4564
-DTD [43](CVPR23) 0.3535 0.1482 0.6470 0.1749 0.1811 0.3009
-FFDN [6](ECCV24) 0.5012 0.2670 0.6085 0.2516 0.2957 0.3848
-TIFDM [12](TCE24) 0.3675 0.2087 0.3572 0.1634 0.2333 0.2660
+CAFTB song2025caftb (TOMM24) 0.6234 0.2557 0.6731 0.3526 0.3770 0.4564
+DTD qu2023doctamper (CVPR23) 0.3535 0.1482 0.6470 0.1749 0.1811 0.3009
+FFDN chen2024ffdn (ECCV24) 0.5012 0.2670 0.6085 0.2516 0.2957 0.3848
+TIFDM dong2024tifdm (TCE24) 0.3675 0.2087 0.3572 0.1634 0.2333 0.2660
 6.2.2 Extending IMDL Detectors to AIGC and Deepfake Benchmarks
@@
 IMDL → AIGC.
-We fine-tune representative IMDL detectors on the training split of the AIGC benchmark and report cross-generator performance in Table 3.
+We fine‑tune representative IMDL detectors on the training split of the AIGC benchmark and report cross‑generator performance in Table 3.
 The training settings and other configurations are consistent with those used in the previously mentioned AIGC benchmark setup.
@@
 IMDL → Deepfake.
-We train each IMDL detector on the FF++-c23 training split and evaluate on all remaining deep-fake test sets; the scores are given in Table 10.
-When compared to all baselines in the deepfakebench [67], Cat-Net attains the best performance in the within-domain setting, while Mesorch achieves the highest average accuracy in the cross-domain evaluation, establishing new state-of-the-art results in both regimes.
+We train each IMDL detector on the FF++‑c23 training split and evaluate on all remaining deep‑fake test sets; the scores are given in Table 10.
+When compared to all baselines in the deepfakebench yan2023deepfakebench , Cat-Net attains the best performance in the within‑domain setting, while Mesorch achieves the highest average accuracy in the cross‑domain evaluation, establishing new state‑of‑the‑art results in both regimes.
 Table 10: Image-level AUC evaluation of IMDL-based detectors trained on FF++-c23 and tested across both within-domain and cross-domain deepfake benchmarks.
 Model Within Domain Evaluation Cross Domain Evaluation
-FF-c23 FF-c40 FF-DF FF-F2F FF-FS FF-NT Avg. CDFv1 CDFv2 DFD DFDC DFDCP Fsh UADFV Avg.
-CAT-Net [22](IJCV22) 0.9855 0.8351 0.9946 0.9891 0.9905 0.9689 0.9606 0.7489 0.7511 0.8086 0.7196 0.7043 0.6626 0.9282 0.7605
-IML-ViT [35](Arxiv) 0.9614 0.8483 0.9810 0.9728 0.9793 0.9206 0.9439 0.7250 0.7419 0.7770 0.7516 0.7237 0.6149 0.8718 0.7437
-Mesorch [70](AAAI25) 0.9815 0.8445 0.9889 0.9890 0.9859 0.9636 0.9589 0.7900 0.7649 0.8468 0.7806 0.7345 0.6638 0.9623 0.7918
-MVSS-Net [5](ICCV21) 0.9723 0.8140 0.9779 0.9865 0.9868 0.9458 0.9472 0.6969 0.6839 0.7749 0.6764 0.6785 0.5841 0.8871 0.7117
-PSCC-Net [29](TCSVT22) 0.7147 0.6536 0.7751 0.6414 0.6460 0.6065 0.6729 0.4602 0.5898 0.5565 0.5945 0.5845 0.6574 0.8216 0.6092
-Trufor [17](CVPR23) 0.7303 0.6848 0.8087 0.7687 0.7084 0.6326 0.7223 0.6551 0.6115 0.5908 0.6017 0.5749 0.5977 0.9526 0.6549
+$F$-c23 $F$-c40 $F$-DF $F$-F2F $F$-FS $F$-NT Avg. CDFv1 CDFv2 DFD DFDC DFDCP Fsh UADFV Avg.
+CAT-Net CAT-Net2022 (IJCV22) 0.9855 0.8351 0.9946 0.9891 0.9905 0.9689 0.9606 0.7489 0.7511 0.8086 0.7196 0.7043 0.6626 0.9282 0.7605
+IML-ViT ma2023iml (Arxiv) 0.9614 0.8483 0.9810 0.9728 0.9793 0.9206 0.9439 0.7250 0.7419 0.7770 0.7516 0.7237 0.6149 0.8718 0.7437
+Mesorch zhu2025mesoscopic (AAAI25) 0.9815 0.8445 0.9889 0.9890 0.9859 0.9636 0.9589 0.7900 0.7649 0.8468 0.7806 0.7345 0.6638 0.9623 0.7918
+MVSS-Net MVSS_2021 (ICCV21) 0.9723 0.8140 0.9779 0.9865 0.9868 0.9458 0.9472 0.6969 0.6839 0.7749 0.6764 0.6785 0.5841 0.8871 0.7117
+PSCC-Net liu2022pscc (TCSVT22) 0.7147 0.6536 0.7751 0.6414 0.6460 0.6065 0.6729 0.4602 0.5898 0.5565 0.5945 0.5845 0.6574 0.8216 0.6092
+Trufor trufor2023 (CVPR23) 0.7303 0.6848 0.8087 0.7687 0.7084 0.6326 0.7223 0.6551 0.6115 0.5908 0.6017 0.5749 0.5977 0.9526 0.6549
+6.3 Grad-CAM Visualization
+We use Grad-CAM to visualize the heatmaps of models from the four domains (Capsule-Net (Deepfake), MVSS-Net (IMDL), UnivFD (AIGC), DTD (Doc)) on datasets from each domain, aiming to explore their attention regions, as shown in Figure 2.
+We use Grad-CAM to visualize the heatmaps of models from the four domains.
+Models from different domains show both common and distinct attention patterns.
+For Deepfake, CapsuleNet focuses on specific facial features, while MVSS-Net attends to larger areas.
+For Doc, CapsuleNet, MVSS-Net, and UnivFD capture overall tampered regions, whereas DTD targets subtle traces like edges and curves of characters.
+Figure 2: Grad-CAM visualization (zoomed in for better visualization).
 7 Conclusion
@@
 In conclusion, ForensicHub represents an important step toward breaking down domain silos across four fields, offering new insights into FIDL future research across model architecture, dataset characteristics, and evaluation standards.
+8 Acknowledgment
+This research was supported by the Sichuan Province Major Special Project (2024ZDZX0001-3), Sichuan Province Natural Science Foundation (Grant No.2024YFHZ0355), and the Science and Technology Development Fund, Macau SAR, under Grant 0193/2023/RIA3 and 0079/2025/AFJ.
+The authors would like to give special thanks to Dr.
+Wentao Feng for the workplace, computation power, and physical infrastructure support.
 Appendix A Limitations
@@
 Average Precision (AP) is calculated as the area under the precision-recall curve. It is defined as:
-[…]
+$\text{AP}=\int_{0}^{1}\text{Precision}(r)\,dr$
 where $\text{Precision}(r)$ is the precision at recall level $r$. The precision is calculated as:
@@
 The Area Under the Curve (AUC) is the area under the Receiver Operating Characteristic (ROC) curve. It can be calculated as:
-[…]
+$\text{AUC}=\int_{0}^{1}\text{TPR}(FPR)\,dFPR$
 where $\text{TPR}(FPR)$ is the true positive rate at a given false positive rate ($FPR$).
@@
 The F1 score is the harmonic mean of precision and recall, given by:
-[…]
+$\text{F1}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$
 where precision and recall are defined as:
@@
 We resize images to 224×224 and apply only basic data augmentations, including flipping, brightness and contrast adjustment, compression, and Gaussian blur.
-All models are trained for 20 epochs using a cosine decay learning rate schedule, decreasing from $41e-$ 4 to $51e-$ 5.
+All models are trained for 20 epochs using a cosine decay learning rate schedule, decreasing from 1e41e-4 to 1e51e-5.
 We use the ImageNet split of DiffusionForensics as the training set.
@@
 Table 11: Cross-dataset AUC evaluation on Deepfake benchmarks.
-Model FF++c40 FF-DF FF-F2F FF-NT FF-FS CDFv1 CDFv2 DFD DFDC DFDCP FaceShifter UADFV Average
+Model FF++c40 $F$-DF $F$-F2F $F$-NT $F$-FS CDFv1 CDFv2 DFD DFDC DFDCP FaceShifter UADFV Average
 ResNet 0.675 0.938 0.936 0.879 0.935 0.692 0.684 0.834 0.678 0.763 0.613 0.770 0.783
@@
 We provide the test results of backbones and domain-specific SoTAs under the IFF-Protocol for each domain, which are Table 11 for Deepfake, Table 12 for IMDL, Table 13 for AIGC, and Table 14 for Document.
+Experiments on Recent Datasets.
+We added experiments on the DF40 [74] and Chameleon [71] dataset, along with evaluations of two recent Deepfake SoTAs: Sia [60] (ECCV22) and Effort [73] (ICML25), and two recent AIGC SoTAs: FatFormer [33] (CVPR24) and CO-SPY [7] (CVPR25).
+Results are shown in Table 15.
+Table 15: AUC performance on recent DF40 and Chameleon datasets.
+Model DF40_CollabDiff DF40_deepfacelab DF40_heygen Chameleon
+ConvNeXt 0.935 0.795 0.744 0.626
+Capsule-Net 0.984 0.638 0.714 0.676
+MVSS-Net 0.983 0.699 0.456 0.725
+UnivFD 0.629 0.752 0.946 0.798
+DTD 0.494 0.477 0.517 0.631
+Sia 0.808 0.764 0.641 0.569
+Effort 0.995 0.894 0.949 0.604
+FatFormer 0.983 0.716 0.611 0.707
+CO-SPY 0.890 0.792 0.656 0.767
+Common features and conflicting patterns across domains.
+We selected two IMDL models: MVSS-Net and IML-ViT, and used IFF-Protocol weights (where models are trained across-domain) as pretrained weights.
+These models were then trained on the IMDL task to investigate whether the artifacts learned across domains could benefit finetuning within a single domain.
+Results are shown in the Table 16.
+Table 16: Common features and conflicting patterns across domains.
+Method Coverage Columbia NIST16 CASIAv1 IMD2020 Avg.
+MVSS-Net (Image-Net) 0.259 0.386 0.246 0.534 0.279 0.341
+MVSS-Net (IFF-Protocol) 0.268 0.395 0.259 0.562 0.292 0.355
+IML-ViT (Image-Net) 0.435 0.780 0.331 0.721 0.327 0.519
+IML-ViT (IFF-Protocol) 0.427 0.767 0.279 0.715 0.351 0.508
 Appendix G Details of Experiments
@@
 DCT [1] ResNet -0.036 -0.015 0.011 -0.016 -0.002 -0.020 -0.036 -0.010 -0.014 -0.035 -0.017
-Xception -0.007 0.000 0.000 0.003 $0$0.001 -0.001 0.000 0.001 -0.004 -0.001
+Xception -0.007 0.000 0.000 0.003 0.000 -0.001 -0.001 0.000 0.001 -0.004 -0.001
 EfficientNet 0.006 0.008 -0.017 0.009 0.013 0.004 -0.009 0.021 0.016 0.030 0.008
@@
 Table 19: Extractor & backbone performance difference in Deepfake region
-Extractor Model FF++c40 FF-DF FF-F2F FF-NT FF-FS CDFv1 CDFv2 DFD DFDC DFDCP FaceShifter UADFV Mean
+Extractor Model FF++c40 $F$-DF $F$-F2F $F$-NT $F$-FS CDFv1 CDFv2 DFD DFDC DFDCP FaceShifter UADFV Mean
 Sobel [5] ResNet [21] -0.052 -0.111 -0.166 -0.150 -0.212 -0.097 -0.042 -0.144 -0.076 -0.063 -0.046 0.095 -0.089
@@
 Xception [8] 0.006 -0.007 -0.013 -0.021 -0.011 0.006 -0.007 -0.023 -0.015 -0.025 -0.008 -0.005 -0.010
-EfficientNet [55] $0$0.011 0.001 0.000 0.002 0.064 0.030 -0.003 0.024 -0.018 0.012 0.004 0.009
+EfficientNet [61] 0.000 -0.011 0.001 0.000 0.002 0.064 0.030 -0.003 0.024 -0.018 0.012 0.004 0.009
 Swin-Transformer [35] -0.050 -0.092 -0.134 -0.185 -0.213 0.056 -0.032 -0.107 -0.097 -0.168 -0.128 -0.006 -0.096
@@
 The table presents detailed results for each individual test dataset.
-They are Table 15 for IMDL, Table 16 for AIGC, Tabel 17 for Deepfake, and Table 18 for Document.
+They are Table 17 for IMDL, Table 18 for AIGC, Tabel 19 for Deepfake, and Table 20 for Document.
 Appendix H Computational Resources
-The experiments were conducted on three different servers.
-The first server is equipped with two AMD EPYC 7542 CPUs, 256GB RAM, and 6$\times$×NVIDIA A40 GPUs, which was used for all IFF-related experiments.
-The remaining experiments were performed on two servers: one with a single AMD EPYC 7542 CPU, 256GB RAM, and 4$\times$×NVIDIA RTX 3090 GPUs, and another with two AMD EPYC 7542 CPUs, 256GB RAM, and 8$\times$×NVIDIA RTX 3090 GPUs.
+[…]
 Appendix I Broader Impacts Discussion
@@
 However, the comprehensive coverage of detection methods in ForensicHub may enable malicious actors to study and develop targeted evasion techniques.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: The main claims reflect the paper’s contributions and scope, which we build the first unified benchmark for all-domain fake image detection and localization.
+  5. 2.
+    Limitations
+  6. Question: Does the paper discuss the limitations of the work performed by the authors?
+  7. Answer: [Yes]
+  8. Justification: We discuss the limitations in the Appendix.
+  9. 3.
+    Theory assumptions and proofs
+  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  11. Answer: [N/A]
+  12. Justification: This paper does not include theory assumptions and proofs.
+  13. 4.
+    Experimental result reproducibility
+  14.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  15. Answer: [Yes]
+  16. Justification: We provide detailed experimental settings in this paper and open-source our code on GitHub.
+  17. 5.
+    Open access to data and code
+  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  19. Answer: [Yes]
+  20. Justification: We provide detailed experimental settings in this paper and open-source our code on GitHub.
+  21. 6.
+    Experimental setting/details
+  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  23. Answer: [Yes]
+  24. Justification: We provide detailed experimental settings in this paper and open-source our code on GitHub.
+  25. 7.
+    Experiment statistical significance
+  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  27. Answer: [No]
+  28. Justification: All experiments were conducted only once, but given the sufficiently large scale, the results can still serve as a valuable reference.
+  29. 8.
+    Experiments compute resources
+  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  31. Answer: [Yes]
+  32. Justification: We provide details of the experimental hardware setup in the Appendix.
+  33. 9.
+    Code of ethics
+  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  35. Answer: [Yes]
+  36. Justification: All experiments in this paper are conducted using existing public datasets and comply with ethical standards.
+  37. 10.
+    Broader impacts
+  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  39. Answer: [Yes]
+  40. Justification: We discuss the proposed paper’s positive and negative societal impacts in the Appendix.
+  41. 11.
+    Safeguards
+  42.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  43. Answer: [N/A]
+  44. Justification: This paper posed no such risks.
+  45. 12.
+    Licenses for existing assets
+  46.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  47. Answer: [Yes]
+  48. Justification: All resources used in this paper are open-source and publicly available.
+  49. 13.
+    New assets
+  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  51. Answer: [Yes]
+  52. Justification: We provide the detailed document along with the open-source code.
+  53. 14.
+    Crowdsourcing and research with human subjects
+  54.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  55. Answer: [N/A]
+  56. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  57. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  58.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  59. Answer: [N/A]
+  60. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  61. 16.
+    Declaration of LLM usage
+  62.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  63. Answer: [N/A]
+  64. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.

### v3

@@
-Bo Du1† Xuekang Zhu1,2† Xiaochen Ma3† Chenfan Qu4, 2† Kaiwen Feng1† Zhe Yang1 Chi-Man Pun5 Jian Liu2‡ Jizhe Zhou1‡
+Bo Du1† Xuekang Zhu1,2† Xiaochen Ma3† Chenfan Qu4, 2† Kaiwen Feng1† Zhe Yang1 Chi-Man Pun5 Jian Liu2‡ Ji-Zhe Zhou1‡
 Abstract
@@
 However, the research efforts of FIDL have gradually split into four relatively independent research domains over time.
-1) Deepfake detection (Deepfake) nguyen2019capsule ; li2018uadfv ; li2020xray ; dang2020ffd ; ni2022core ; cao2022recce ; yan2023ucf ; han2023fcd ; qian2020f3net ; liu2021spsl ; luo2021srm ; sun2022SIA ; zhuang2022uia ; shiohara2022sbi ; yan2024LSDA ; yan2023deepfakebench : detects human-centric manipulations such as face swapping, expression editing, or feature replacement.
-2) Image manipulation detection/localization (IMDL) liu2020d ; MVSS_2021 ; trufor2023 ; liu2022pscc ; NCL_IML_2023 ; ma2023iml ; yu2024diffforensics ; zhu2025mesoscopic ; sparseViT_2025 : detects and localizes the tampering in natural images.
-3) AI-Generated Image Detection (AIGC) univfd ; HiFi-Net2023 ; dualnet ; wang2023dire ; zhang2025towards : detects the images fully generated by deep generative models such as Stable Diffusion stablediffusion .
-4) Document Image Manipulation Localization (Document)  psnet ; qu2023doctamper ; chen2024ffdn ; dong2024tifdm ; song2025caftb ; yu2023learning ; li2026document : localizes the forgery of various forms of document images, including receipts, certificates, and identification materials, with a particular focus on detecting modifications to the printed text.
+1) Deepfake detection (Deepfake) Nguyen et al.
+(2019); Li and Lyu (2018); Li et al.
+(2020b); Dang et al.
+(2020); Ni et al.
+(2022); Cao et al.
+(2022); Yan et al.
+(2023a); Han et al.
+(2023); Qian et al.
+(2020); Liu et al.
+(2021a); Luo et al.
+(2021); Sun et al.
+(2022b); Zhuang et al.
+(2022); Shiohara and Yamasaki (2022); Yan et al.
+(2024b, 2023b): detects human-centric manipulations such as face swapping, expression editing, or feature replacement.
+2) Image manipulation detection/localization (IMDL) Liu et al.
+(2020); Chen et al.
+(2021); Guillaro et al.
+(2023); Liu et al.
+(2022a); Zhou et al.
+(2023); Ma et al.
+(2023b); Yu et al.
+(2024); Zhu et al.
+(2025); Su et al.
+(2025): detects and localizes the tampering in natural images.
+3) AI-Generated Image Detection (AIGC) Ojha et al.
+(2023); Guo et al.
+(2023); Xi et al.
+(2023); Wang et al.
+(2023); Zhang et al.
+(2025): detects the images fully generated by deep generative models such as Stable Diffusion Rombach et al.
+(2022).
+4) Document Image Manipulation Localization (Document)  Shao et al.
+(2023); Qu et al.
+(2023); Chen et al.
+(2024); Dong et al.
+(2024); Song et al.
+(2025); Yu et al.
+(2023); Li et al.
+(2026): localizes the forgery of various forms of document images, including receipts, certificates, and identification materials, with a particular focus on detecting modifications to the printed text.
 Although these domains have become isolated due to differences in application scenarios, manipulation types, and detection methods, there are still overlaps and similarities among them.
@@
 Task Model Backbone Artifact Strategy Output Type Contribution
-Deepfake Capsule-Net nguyen2019capsule (ICASSP19) VGG VGG_2015 Dynamic Routing Label Proposes a capsule network with dynamic routing and a VGG19 backbone.
-RECCE cao2022recce (CVPR22) Xception chollet2017xception Reconstruction Label Proposes a graph-based framework leveraging reconstruction differences
-SPSL liu2021spsl (CVPR21) Xception chollet2017xception Phase Spectrum Label Proposes phase-spectrum fusion with Xception for face forgery detection.
-UCF yan2023ucf (ICCV23) Xception chollet2017xception Multi-task Disentanglement Label Proposes multi-task disentanglement with Xception for deepfake generalization.
-SBI shiohara2022sbi (CVPR22) EfficientNet tan2019efficientnet Frequency,Blending Boundaries Label Proposes self-blended images to improve deepfake detection generalization.
-IMDL MVSS-Net Mantra_2019 (ICCV21) Resnet Resnet_2016 BayarConv,Sobel Label,Mask Exploit noise and boundary artifacts via multi-view learning for manipulation detection.
-CAT-Net CAT-Net2022 (IJCV22) HRNet HRNet DCT Mask Fuse RGB and DCT streams to learn compression artifacts for splice localization.
-PSCC-Net liu2022pscc (TCSVT22) HRNet HRNet Multi-Resolution Conv Label,Mask Progressively refine masks with spatio-channel correlations for high-reso localization.
-Trufor trufor2023 (CVPR23) Seformer SegFormer_2021 High-Reso,Multi-scale,Edge Label,Mask Fuse RGB and learned noise fingerprints to detect manipulations as anomalies.
-IML-ViT ma2023iml (Arxiv) ViT ViT_2021 BayarConv,SRM Filter Mask Use ViT with high-reso, multi-scale edge-aware design for manipulation localization.
-Mesorch zhu2025mesoscopic (AAAI25) Conv. Convnet_2022 ,Segfor. SegFormer_2021 DCT Mask Fuse micro and macro cues for mesoscopic image manipulation localization.
-AIGC Dire wang2023dire (ICCV23) Resnet Resnet_2016 Diffusion Reconstruction Label Use reconstruction error of diffusion for diffusion-generated images detection.
-DualNet dualnet (APSIPA23) CNN SRM,Low Frequency Label Fuse SRM residual and low-frequency content streams for AIGC detection.
-HiFiNet HiFi-Net2023 (CVPR23) HRNet HRNet Multi-branch Feature Extractor Label,Mask Learn hierarchical fine-grained representations of forgery attributes.
-Synthbuster bammey2023synthbuster (OJSP23) None Fourier Transform Label Leverage spectral artifacts in the frequency domain for diffusion detection.
-UnivFD univfd (CVPR23) CLIP-ViT clip None Label Use pretrained vision-language model features for unified detection.
-Document CAFTB song2025caftb (TOMM24) Resnet Resnet_2016 SRM Mask Proposes CAFTB-Net with dual-branch and cross-attention.
-TIFDM dong2024tifdm (TCE24) Resnet Resnet_2016 None Mask Proposes a robust network with multiscale attention.
-DTD qu2023doctamper (CVPR23) Conv. Convnet_2022 , Swin.Swin_2021 Frequency Mask Proposes DTD with frequency head and multi-view decoder.
-FFDN chen2024ffdn (ECCV24) ConvNext Convnet_2022 Wavelet, Frequency Mask Proposes FFDN combining visual enhancement and frequency decomposition
-Although individual benchmarks exist in some domains, such as DeepfakeBench yan2023deepfakebench for Deepfake and IMDLBenCo ma2024imdl for IMDL, a unified benchmark for all domains in FIDL remains blank.
+Deepfake Capsule-Net Nguyen et al. (2019)(ICASSP19) VGG Simonyan and Zisserman (2015) Dynamic Routing Label Proposes a capsule network with dynamic routing and a VGG19 backbone.
+RECCE Cao et al. (2022)(CVPR22) Xception Chollet (2017) Reconstruction Label Proposes a graph-based framework leveraging reconstruction differences
+SPSL Liu et al. (2021a)(CVPR21) Xception Chollet (2017) Phase Spectrum Label Proposes phase-spectrum fusion with Xception for face forgery detection.
+UCF Yan et al. (2023a)(ICCV23) Xception Chollet (2017) Multi-task Disentanglement Label Proposes multi-task disentanglement with Xception for deepfake generalization.
+SBI Shiohara and Yamasaki (2022)(CVPR22) EfficientNet Tan and Le (2019) Frequency,Blending Boundaries Label Proposes self-blended images to improve deepfake detection generalization.
+IMDL MVSS-Net Wu and others (2019)(ICCV21) Resnet He et al. (2016) BayarConv,Sobel Label,Mask Exploit noise and boundary artifacts via multi-view learning for manipulation detection.
+CAT-Net Kwon et al. (2022)(IJCV22) HRNet Wang et al. (2020) DCT Mask Fuse RGB and DCT streams to learn compression artifacts for splice localization.
+PSCC-Net Liu et al. (2022a)(TCSVT22) HRNet Wang et al. (2020) Multi-Resolution Conv Label,Mask Progressively refine masks with spatio-channel correlations for high-reso localization.
+Trufor Guillaro et al. (2023)(CVPR23) Seformer Xie et al. (2021) High-Reso,Multi-scale,Edge Label,Mask Fuse RGB and learned noise fingerprints to detect manipulations as anomalies.
+IML-ViT Ma et al. (2023b)(Arxiv) ViT Dosovitskiy et al. (2021) None Mask Use ViT with high-reso, multi-scale edge-aware design for manipulation localization.
+Mesorch Zhu et al. (2025)(AAAI25) Conv. Liu et al. (2022b),Segfor. Xie et al. (2021) DCT Mask Fuse micro and macro cues for mesoscopic image manipulation localization.
+AIGC Dire Wang et al. (2023)(ICCV23) Resnet He et al. (2016) Diffusion Reconstruction Label Use reconstruction error of diffusion for diffusion-generated images detection.
+DualNet Xi et al. (2023)(APSIPA23) CNN SRM,Low Frequency Label Fuse SRM residual and low-frequency content streams for AIGC detection.
+HiFiNet Guo et al. (2023)(CVPR23) HRNet Wang et al. (2020) Multi-branch Feature Extractor Label,Mask Learn hierarchical fine-grained representations of forgery attributes.
+Synthbuster Bammey (2023)(OJSP23) None Fourier Transform Label Leverage spectral artifacts in the frequency domain for diffusion detection.
+UnivFD Ojha et al. (2023)(CVPR23) CLIP-ViT Radford et al. (2021) None Label Use pretrained vision-language model features for unified detection.
+Document CAFTB Song et al. (2025)(TOMM24) Resnet He et al. (2016) SRM Mask Proposes CAFTB-Net with dual-branch and cross-attention.
+TIFDM Dong et al. (2024)(TCE24) Resnet He et al. (2016) None Mask Proposes a robust network with multiscale attention.
+DTD Qu et al. (2023)(CVPR23) Conv. Liu et al. (2022b), Swin.Liu et al. (2021b) Frequency Mask Proposes DTD with frequency head and multi-view decoder.
+FFDN Chen et al. (2024)(ECCV24) ConvNext Liu et al. (2022b) Wavelet, Frequency Mask Proposes FFDN combining visual enhancement and frequency decomposition
+Although individual benchmarks exist in some domains, such as DeepfakeBench Yan et al.
+(2023b) for Deepfake and IMDLBenCo Ma et al.
+(2024) for IMDL, a unified benchmark for all domains in FIDL remains blank.
 The absence of such a unified benchmark results in significant domain silos, where each domain independently constructs its datasets, models, and evaluation protocols without interoperability.
@@
 Despite the rapid development of these tasks, there is a lack of a unified benchmark, with some task having its isolated benchmark, creating barriers between them.
-DeepfakeBench yan2023deepfakebench is a Deepfake detection benchmark specifically designed to address the lack of uniformity in data processing pipelines, leading to inconsistent data inputs for detection models.
-IMDLBenCo ma2024imdl is a benchmark and codebase for IMDL, aiming to compare IMDL models through a unified training and evaluation protocol.
-AIGCDetectBenchmark EkkoAIGCDetectBenchmark is a repository for experiments on 9 AI-generated image detection methods.
+DeepfakeBench Yan et al.
+(2023b) is a Deepfake detection benchmark specifically designed to address the lack of uniformity in data processing pipelines, leading to inconsistent data inputs for detection models.
+IMDLBenCo Ma et al.
+(2024) is a benchmark and codebase for IMDL, aiming to compare IMDL models through a unified training and evaluation protocol.
+AIGCDetectBenchmark Zhong et al.
+(2023) is a repository for experiments on 9 AI-generated image detection methods.
 These benchmarks provide models, datasets, and evaluation metrics within their respective tasks, but their underlying designs lack cross-task considerations, making them difficult to integrate across different detection scenarios.
@@
 Construction of ForensicHub.
-To enable broad interoperability and reduce duplication of effort, ForensicHub adopts an adapter-based design gamma1995design that ensures seamless integration with DeepfakeBench yan2023deepfakebench and IMDLBenCo ma2024imdl , two widely used benchmarks.
+To enable broad interoperability and reduce duplication of effort, ForensicHub adopts an adapter-based design Gamma (1995) that ensures seamless integration with DeepfakeBench Yan et al.
+(2023b) and IMDLBenCo Ma et al.
+(2024), two widely used benchmarks.
 This mechanism allows users to reuse existing models and datasets without major modification, while also supporting the definition of new models and benchmarks within ForensicHub under the unified protocol.
@@
 In total, ForensicHub covers 4 tasks, 23 datasets, 42 models, 6 backbones, and implements 11 commonly used image- and pixel-level metrics.
-Datasets used in this paper are: FaceForensics++ rossler2019faceforensics++ , Celeb-DF li2020celeb , DFD dolhansky2020dfd , FaceShifter li2019faceshifter and UADFV li2018uadfv for Deepfake; CASIA CASIA_2013 , COVERAGE Coverage_2016 , Columbia Columbia_2006 , IMD2020 IMD20_2020 , NIST16 NIST16_2019 , CocoGlide trufor2023 , and Autosplice jia2023autosplice for IMDL; DiffusionForensics wang2023dire , GenImage zhu2023genimage for AIGC; Doctamper qu2023doctamper , T-SROIE wang2022tsroie , OSTF qu2025ostf , TPIC-13 wang2022tpic , RTM luo2025rtm for Doc.
+Datasets used in this paper are: FaceForensics++ Rossler et al.
+(2019), Celeb-DF Li et al.
+(2020c), DFD Dolhansky et al.
+(2020), FaceShifter Li et al.
+(2019) and UADFV Li and Lyu (2018) for Deepfake; CASIA Dong et al.
+(2013), COVERAGE Wen et al.
+(2016), Columbia Hsu and Chang (2006), IMD2020 Novozamsky et al.
+(2020), NIST16 Guan et al.
+(2019), CocoGlide Guillaro et al.
+(2023), and Autosplice Jia et al.
+(2023) for IMDL; DiffusionForensics Wang et al.
+(2023), GenImage Zhu et al.
+(2023) for AIGC; Doctamper Qu et al.
+(2023), T-SROIE Wang et al.
+(2022b), OSTF Qu et al.
+(2025), TPIC-13 Wang et al.
+(2022a), RTM Luo et al.
+(2025) for Doc.
 A brief summary of each dataset is provided in Table 2, with more details in Appendix D.1.
@@
 Task Dataset Year Content Real Fake Annotation Pipeline Split
-Deepfake FaceForensics++ rossler2019faceforensics++ 2019 Face 45,388 127,209 Label Automatic Train,Test
-Celeb-DF-v1 li2020celeb 2020 Face 1,203 1,933 Label Automatic Train,Test
-Celeb-DF-v2 li2020celeb 2020 Face 5,620 10,800 Label Automatic Train,Test
-DeepFakeDetection dfd2019 2019 Face 10,741 91,800 Label Automatic Train,Test
-DFDC dolhansky2020dfd 2020 Face 63,265 68,851 Label Automatic Train,Test
-DFDCP dolhansky2019deepfake 2019 Face 5,901 11,321 Label Automatic Train,Test
-FaceShifter li2020advancing 2020 Face 4,479 4,479 Label Automatic Train,Test
-UADFV li2018uadfv 2018 Face 1,548 1,551 Label Automatic Train,Test
-IMDL CASIAv1 CASIA_2013 2013 General 0 920 Label,Mask Manual Train
-CASIAv2 CASIA_2013 2013 General 7,491 5,123 Label,Mask Manual Train
-COVERAGE Coverage_2016 2016 General 100 100 Label,Mask Manual Train
-Columbia Columbia_2006 2006 General 183 180 Label,Mask Automatic Train
-IMD2020 CASIA_2013 2020 General 414 2,010 Label,Mask Manual Train
-NIST16 NIST16_2019 2019 General 873 564 Label,Mask Manual Train
-CocoGlide trufor2023 2023 General 512 512 Label,Mask Automatic Train
-Autosplice jia2023autosplice 2023 General 2,273 3,621 Label,Mask Manual Train
-AIGC DiffusionForensics wang2023dire 2023 General 134,000 481,200 Label Automatic Train,Val,Test
-GenImage zhu2023genimage 2023 General 1,331,167 1,350,000 Label Automatic Train,Test
-Document Doctamper qu2023doctamper 2023 Document 0 170,000 Mask Automatic Train,Test
-OSTF qu2025ostf 2025 Document 16,405 3,685 Mask Manual Train,Test
-RealTextManipulation luo2025rtm 2025 Document 52,314 17,423 Mask Automatic+Manual Train,Test
-T-SROIE wang2022tsroie 2022 Document 21,268 4,326 Mask Manual Train,Test
-Tampered-IC13 wang2022tpic 2022 Document 2,810 1,228 Mask Manual Train,Test
-Models used in this paper are: Capsule-Net nguyen2019capsule , RECCE cao2022recce , SPSL liu2021spsl , UCF yan2023ucf , and SBI shiohara2022sbi for Deepfake detection; MVSS-Net MVSS_2021 , CAT-Net CAT-Net2022 , PSCC-Net liu2022pscc , Trufor trufor2023 , IML-ViT ma2023iml , and Mesorch zhu2025mesoscopic for image manipulation and localization; Dire wang2023dire , DualNet dualnet , HiFiNet HiFi-Net2023 , Synthbuster bammey2023synthbuster , and UnivFD univfd for AIGC detection; DTD qu2023doctamper , FFDN chen2024ffdn , CAFTB song2025caftb , TIFDM dong2024tifdm for document detection.
+Deepfake FaceForensics++ Rossler et al. (2019) 2019 Face 45,388 127,209 Label Automatic Train,Test
+Celeb-DF-v1 Li et al. (2020c) 2020 Face 1,203 1,933 Label Automatic Train,Test
+Celeb-DF-v2 Li et al. (2020c) 2020 Face 5,620 10,800 Label Automatic Train,Test
+DeepFakeDetection Google AI Blog (2019) 2019 Face 10,741 91,800 Label Automatic Train,Test
+DFDC Dolhansky et al. (2020) 2020 Face 63,265 68,851 Label Automatic Train,Test
+DFDCP Dolhansky et al. (2019) 2019 Face 5,901 11,321 Label Automatic Train,Test
+FaceShifter Li et al. (2020a) 2020 Face 4,479 4,479 Label Automatic Train,Test
+UADFV Li and Lyu (2018) 2018 Face 1,548 1,551 Label Automatic Train,Test
+IMDL CASIAv1 Dong et al. (2013) 2013 General 0 920 Label,Mask Manual Train
+CASIAv2 Dong et al. (2013) 2013 General 7,491 5,123 Label,Mask Manual Train
+COVERAGE Wen et al. (2016) 2016 General 100 100 Label,Mask Manual Train
+Columbia Hsu and Chang (2006) 2006 General 183 180 Label,Mask Automatic Train
+IMD2020 Dong et al. (2013) 2020 General 414 2,010 Label,Mask Manual Train
+NIST16 Guan et al. (2019) 2019 General 873 564 Label,Mask Manual Train
+CocoGlide Guillaro et al. (2023) 2023 General 512 512 Label,Mask Automatic Train
+Autosplice Jia et al. (2023) 2023 General 2,273 3,621 Label,Mask Manual Train
+AIGC DiffusionForensics Wang et al. (2023) 2023 General 134,000 481,200 Label Automatic Train,Val,Test
+GenImage Zhu et al. (2023) 2023 General 1,331,167 1,350,000 Label Automatic Train,Test
+Document Doctamper Qu et al. (2023) 2023 Document 0 170,000 Mask Automatic Train,Test
+OSTF Qu et al. (2025) 2025 Document 16,405 3,685 Mask Manual Train,Test
+RealTextManipulation Luo et al. (2025) 2025 Document 52,314 17,423 Mask Automatic+Manual Train,Test
+T-SROIE Wang et al. (2022b) 2022 Document 21,268 4,326 Mask Manual Train,Test
+Tampered-IC13 Wang et al. (2022a) 2022 Document 2,810 1,228 Mask Manual Train,Test
+Models used in this paper are: Capsule-Net Nguyen et al.
+(2019), RECCE Cao et al.
+(2022), SPSL Liu et al.
+(2021a), UCF Yan et al.
+(2023a), and SBI Shiohara and Yamasaki (2022) for Deepfake detection; MVSS-Net Chen et al.
+(2021), CAT-Net Kwon et al.
+(2022), PSCC-Net Liu et al.
+(2022a), Trufor Guillaro et al.
+(2023), IML-ViT Ma et al.
+(2023b), and Mesorch Zhu et al.
+(2025) for image manipulation and localization; Dire Wang et al.
+(2023), DualNet Xi et al.
+(2023), HiFiNet Guo et al.
+(2023), Synthbuster Bammey (2023), and UnivFD Ojha et al.
+(2023) for AIGC detection; DTD Qu et al.
+(2023), FFDN Chen et al.
+(2024), CAFTB Song et al.
+(2025), TIFDM Dong et al.
+(2024) for document detection.
 These methods are from official repositories and our reimplementations.
-In addition, ForensicHub also selects 6 commonly used backbones in visual tasks, which are: Resnet Resnet_2016 , Xception chollet2017xception , EfficientNet tan2019efficientnet , Segformer SegFormer_2021 , Swin Transformer Swin_2021 , and ConvNext Convnet_2022 .
+In addition, ForensicHub also selects 6 commonly used backbones in visual tasks, which are: Resnet He et al.
+(2016), Xception Chollet (2017), EfficientNet Tan and Le (2019), Segformer Xie et al.
+(2021), Swin Transformer Liu et al.
+(2021b), and ConvNext Liu et al.
+(2022b).
 Details about models can be found in Appendix D.2.
@@
 4 Benchmarks
-In addition to being fully compatible with existing benchmarks, DeepfakeBench yan2023deepfakebench and IMDLBenCo ma2024imdl , ForensicHub further extends standardization efforts by introducing unified evaluation protocols for the AIGC and Document domains—two areas that previously lacked widely accepted benchmarks and codebases.
+In addition to being fully compatible with existing benchmarks, DeepfakeBench Yan et al.
+(2023b) and IMDLBenCo Ma et al.
+(2024), ForensicHub further extends standardization efforts by introducing unified evaluation protocols for the AIGC and Document domains—two areas that previously lacked widely accepted benchmarks and codebases.
 We propose two protocols for two domains to evaluate generalization.
@@
 In the field of AIGC detection, the challenge in dataset construction usually lies not in obtaining a sufficient quantity of samples, since they can be easily generated using existing models, but in ensuring comprehensive coverage of a wide range of generative models.
-Therefore, we select only two commonly used public datasets: DiffusionForensics wang2023dire and GenImage zhu2023genimage .
+Therefore, we select only two commonly used public datasets: DiffusionForensics Wang et al.
+(2023) and GenImage Zhu et al.
+(2023).
 The former contains only diffusion-based generated images, while the latter covers a million-scale dataset constructed from eight SoTA generative models.
-Models are trained on DiffusionForensics and evaluated on different generative models within GenImage to assess generalization, as detection methods typically already achieve good performance on samples from the same generative model zhu2023genimage .
+Models are trained on DiffusionForensics and evaluated on different generative models within GenImage to assess generalization, as detection methods typically already achieve good performance on samples from the same generative model Zhu et al.
+(2023).
 The detailed data splits are summarized in Table D.1.3.
 Models.
-ForensicHub implements five SoTA methods in AIGC detection: Dire wang2023dire , DualNet dualnet , HiFiNet HiFi-Net2023 , Synthbuster bammey2023synthbuster , and UnivFD univfd , among which Synthbuster has no official open-source code and is fully reimplemented by us.
+ForensicHub implements five SoTA methods in AIGC detection: Dire Wang et al.
+(2023), DualNet Xi et al.
+(2023), HiFiNet Guo et al.
+(2023), Synthbuster Bammey (2023), and UnivFD Ojha et al.
+(2023), among which Synthbuster has no official open-source code and is fully reimplemented by us.
 More details about models and training settings can be found in Appendix E.1.
 Results.
-Table 3 in green background presents the AUC scores for image-level classification for AIGC benchmark, divided into in-domain results on the test set split of DiffusionForensics wang2023dire , and cross-domain results on different generative models and the total set on GenImage zhu2023genimage .
+Table 3 in green background presents the AUC scores for image-level classification for AIGC benchmark, divided into in-domain results on the test set split of DiffusionForensics Wang et al.
+(2023), and cross-domain results on different generative models and the total set on GenImage Zhu et al.
+(2023).
 The results show that AIGC SoTAs generally achieve excellent performance on the DiffusionForensics test set, which shares the same source as the training set, and also perform well on datasets composed of diffusion-based generated images like ADM, VQDM, and GLIDE that are similar to the training data.
@@
 DiffusionForensics ADM BigGAN Midjourney VQDM GLIDE SD V1.4 SD V1.5 Wukong AverageC
-DualNet dualnet (APSIPA23) 1.0000 0.9986 0.9172 0.8495 0.9901 0.9835 0.8127 0.8081 0.7925 0.8902
-HiFiNet HiFi-Net2023 (CVPR23) 1.0000 0.9998 0.8407 0.7210 0.9999 0.9991 0.6765 0.6747 0.6675 0.8160
-Synthbuster bammey2023synthbuster (OJSP23) 0.6662 0.6226 0.8550 0.4565 0.6308 0.7719 0.4606 0.4614 0.3902 0.5762
-UnivFD univfd (CVPR23) 0.9947 0.8400 0.9687 0.5427 0.9663 0.9541 0.7078 0.7072 0.6781 0.7920
-MVSS-Net MVSS_2021 (ICCV21) 0.9992 0.9706 0.9520 0.6242 0.8894 0.9380 0.7207 0.7192 0.6089 0.7997
-CAT-Net CAT-Net2022 (IJCV22) 0.9985 0.9391 0.9327 0.6423 0.8226 0.7987 0.7171 0.7142 0.6156 0.7707
-Trufor trufor2023 (CVPR23) 0.9999 0.9750 0.9316 0.7079 0.9340 0.9144 0.8683 0.8687 0.8018 0.8752
-IML-ViT ma2023iml (Arxiv) 1.0000 0.9594 0.9152 0.8086 0.9577 0.9464 0.8924 0.8877 0.8006 0.8957
-Mesorch zhu2025mesoscopic (AAAI25) 0.9998 0.9754 0.9667 0.7011 0.9331 0.9253 0.8168 0.8179 0.7441 0.8582
+DualNet Xi et al. (2023)(APSIPA23) 1.0000 0.9986 0.9172 0.8495 0.9901 0.9835 0.8127 0.8081 0.7925 0.8902
+HiFiNet Guo et al. (2023)(CVPR23) 1.0000 0.9998 0.8407 0.7210 0.9999 0.9991 0.6765 0.6747 0.6675 0.8160
+Synthbuster Bammey (2023)(OJSP23) 0.6662 0.6226 0.8550 0.4565 0.6308 0.7719 0.4606 0.4614 0.3902 0.5762
+UnivFD Ojha et al. (2023)(CVPR23) 0.9947 0.8400 0.9687 0.5427 0.9663 0.9541 0.7078 0.7072 0.6781 0.7920
+MVSS-Net Chen et al. (2021)(ICCV21) 0.9992 0.9706 0.9520 0.6242 0.8894 0.9380 0.7207 0.7192 0.6089 0.7997
+CAT-Net Kwon et al. (2022)(IJCV22) 0.9985 0.9391 0.9327 0.6423 0.8226 0.7987 0.7171 0.7142 0.6156 0.7707
+Trufor Guillaro et al. (2023)(CVPR23) 0.9999 0.9750 0.9316 0.7079 0.9340 0.9144 0.8683 0.8687 0.8018 0.8752
+IML-ViT Ma et al. (2023b)(Arxiv) 1.0000 0.9594 0.9152 0.8086 0.9577 0.9464 0.8924 0.8877 0.8006 0.8957
+Mesorch Zhu et al. (2025)(AAAI25) 0.9998 0.9754 0.9667 0.7011 0.9331 0.9253 0.8168 0.8179 0.7441 0.8582
 Dataset.
-Existing datasets for document image manipulation localization can be broadly categorized into two types: high-fidelity non-sliced datasets, including T-SROIE wang2022tsroie , OSTF qu2025ostf , TPIC-13 wang2022tpic , and RTM luo2025rtm ; and sliced datasets, represented by Doctamper qu2023doctamper .
+Existing datasets for document image manipulation localization can be broadly categorized into two types: high-fidelity non-sliced datasets, including T-SROIE Wang et al.
+(2022b), OSTF Qu et al.
+(2025), TPIC-13 Wang et al.
+(2022a), and RTM Luo et al.
+(2025); and sliced datasets, represented by Doctamper Qu et al.
+(2023).
 The primary difference lies in whether the images are preprocessed using patch-wise slicing.
@@
 Model.
-We employ two open-source models, DTD qu2023doctamper and FFDN chen2024ffdn , and reproduce two closed-source models, CATFB song2025caftb and TIFDM dong2024tifdm .
+We employ two open-source models, DTD Qu et al.
+(2023) and FFDN Chen et al.
+(2024), and reproduce two closed-source models, CATFB Song et al.
+(2025) and TIFDM Dong et al.
+(2024).
 All details are provided in Appendix D.2.4 and E.2.
 Results.
-Following the original protocols qu2023doctamper ; chen2024ffdn , each detector is trained on its designated training split and evaluated on the corresponding test split.
+Following the original protocols Qu et al.
+(2023); Chen et al.
+(2024), each detector is trained on its designated training split and evaluated on the corresponding test split.
 As shown in Table 4, three models consistently achieve top performance: FFDN and DTD, both designed specifically for document forensics, and Cat-Net, an IMDL-based model.
@@
 DocTamperFCD DocTamperSCD DocTamperTest AverageD T-SROIE-test OSTF-test TPIC-13-test RTM-test
-CAFTB song2025caftb (TOMM24) 0.2917 0.3770 0.3275 0.3321 0.9165 0.6478 0.8394 0.2493 0.5213
-DTD qu2023doctamper (CVPR23) 0.6856 0.7392 0.8031 0.7426 0.9205 0.5626 0.8341 0.1720 0.6739
-FFDN chen2024ffdn (ECCV24) 0.8773 0.7392 0.8212 0.8126 0.9137 0.5586 0.8186 0.1589 0.6982
-TIFDM dong2024tifdm (TCE24) 0.0896 0.2572 0.2585 0.2018 0.8942 0.5410 0.7972 0.0591 0.4138
-MVSS-Net MVSS_2021 (ICCV21) 0.2066 0.3710 0.3810 0.3195 0.8756 0.5254 0.7864 0.1314 0.4682
-PSCC-Net liu2022pscc (TCSVT22) 0.3855 0.3931 0.4972 0.4253 0.9305 0.5697 0.7894 0.1971 0.5375
-Cat-Net CAT-Net2022 (IJCV22) 0.7600 0.6405 0.7644 0.7216 0.8726 0.5371 0.7947 0.1174 0.6410
-IML-ViT ma2023iml (Arxiv) 0.4688 0.5117 0.4486 0.4764 0.8731 0.5128 0.7202 0.0920 0.5182
-Trufor trufor2023 (CVPR23) 0.2613 0.3124 0.2517 0.2751 0.9095 0.6485 0.8197 0.2302 0.4905
-Mesorch zhu2025mesoscopic (AAAI25) 0.4231 0.4586 0.4651 0.4489 0.9049 0.4804 0.7607 0.1314 0.5177
+CAFTB Song et al. (2025)(TOMM24) 0.2917 0.3770 0.3275 0.3321 0.9165 0.6478 0.8394 0.2493 0.5213
+DTD Qu et al. (2023)(CVPR23) 0.6856 0.7392 0.8031 0.7426 0.9205 0.5626 0.8341 0.1720 0.6739
+FFDN Chen et al. (2024)(ECCV24) 0.8773 0.7392 0.8212 0.8126 0.9137 0.5586 0.8186 0.1589 0.6982
+TIFDM Dong et al. (2024)(TCE24) 0.0896 0.2572 0.2585 0.2018 0.8942 0.5410 0.7972 0.0591 0.4138
+MVSS-Net Chen et al. (2021)(ICCV21) 0.2066 0.3710 0.3810 0.3195 0.8756 0.5254 0.7864 0.1314 0.4682
+PSCC-Net Liu et al. (2022a)(TCSVT22) 0.3855 0.3931 0.4972 0.4253 0.9305 0.5697 0.7894 0.1971 0.5375
+Cat-Net Kwon et al. (2022)(IJCV22) 0.7600 0.6405 0.7644 0.7216 0.8726 0.5371 0.7947 0.1174 0.6410
+IML-ViT Ma et al. (2023b)(Arxiv) 0.4688 0.5117 0.4486 0.4764 0.8731 0.5128 0.7202 0.0920 0.5182
+Trufor Guillaro et al. (2023)(CVPR23) 0.2613 0.3124 0.2517 0.2751 0.9095 0.6485 0.8197 0.2302 0.4905
+Mesorch Zhu et al. (2025)(AAAI25) 0.4231 0.4586 0.4651 0.4489 0.9049 0.4804 0.7607 0.1314 0.5177
 However, this evaluation setting has a key limitation: all models are trained and tested within the same distribution, limiting the assessment of cross-domain generalization.
@@
 DocTamperFCD DocTamperSCD DocTamperTest AverageW T-SROIE OSTF TPIC-13 RTM AverageC
-CAFTB song2025caftb (TOMM24) 0.2917 0.3770 0.3275 0.3321 0.2617 0.1194 0.3007 0.0328 0.1787 0.2444
-DTD qu2023doctamper (CVPR23) 0.6856 0.7392 0.8031 0.7426 0.5245 0.1241 0.2835 0.0575 0.2474 0.4596
-FFDN chen2024ffdn (ECCV24) 0.8773 0.7392 0.8212 0.8126 0.5330 0.2409 0.3572 0.0708 0.3005 0.5199
-TIFDM dong2024tifdm (TCE24) 0.0896 0.2572 0.2585 0.2018 0.0582 0.0058 0.0134 0.0176 0.0238 0.1000
-MVSS-Net MVSS_2021 (ICCV23) 0.2066 0.3710 0.3810 0.3195 0.1870 0.0373 0.1134 0.0268 0.0911 0.1890
-PSCC-Net liu2022pscc (TCSVT22) 0.3855 0.3931 0.4972 0.4253 0.5168 0.4414 0.5495 0.1255 0.4083 0.4156
-Cat-Net CAT-Net2022 (IJCV22) 0.7600 0.6405 0.7644 0.7216 0.6085 0.1777 0.3430 0.0630 0.2981 0.4796
-IML-ViT ma2023iml (Arxiv) 0.4688 0.5117 0.4486 0.4764 0.4269 0.2101 0.2563 0.0764 0.2424 0.3427
-Trufor trufor2023 (CVPR23) 0.2613 0.3124 0.2517 0.2751 0.2126 0.0464 0.1038 0.0342 0.0993 0.1746
-Mesorch zhu2025mesoscopic (AAAI25) 0.4231 0.4586 0.4651 0.4489 0.2937 0.1388 0.2408 0.0405 0.1785 0.2944
+CAFTB Song et al. (2025)(TOMM24) 0.2917 0.3770 0.3275 0.3321 0.2617 0.1194 0.3007 0.0328 0.1787 0.2444
+DTD Qu et al. (2023)(CVPR23) 0.6856 0.7392 0.8031 0.7426 0.5245 0.1241 0.2835 0.0575 0.2474 0.4596
+FFDN Chen et al. (2024)(ECCV24) 0.8773 0.7392 0.8212 0.8126 0.5330 0.2409 0.3572 0.0708 0.3005 0.5199
+TIFDM Dong et al. (2024)(TCE24) 0.0896 0.2572 0.2585 0.2018 0.0582 0.0058 0.0134 0.0176 0.0238 0.1000
+MVSS-Net Chen et al. (2021)(ICCV23) 0.2066 0.3710 0.3810 0.3195 0.1870 0.0373 0.1134 0.0268 0.0911 0.1890
+PSCC-Net Liu et al. (2022a)(TCSVT22) 0.3855 0.3931 0.4972 0.4253 0.5168 0.4414 0.5495 0.1255 0.4083 0.4156
+Cat-Net Kwon et al. (2022)(IJCV22) 0.7600 0.6405 0.7644 0.7216 0.6085 0.1777 0.3430 0.0630 0.2981 0.4796
+IML-ViT Ma et al. (2023b)(Arxiv) 0.4688 0.5117 0.4486 0.4764 0.4269 0.2101 0.2563 0.0764 0.2424 0.3427
+Trufor Guillaro et al. (2023)(CVPR23) 0.2613 0.3124 0.2517 0.2751 0.2126 0.0464 0.1038 0.0342 0.0993 0.1746
+Mesorch Zhu et al. (2025)(AAAI25) 0.4231 0.4586 0.4651 0.4489 0.2937 0.1388 0.2408 0.0405 0.1785 0.2944
 5 Image Forensic Fusion Protocol
@@
 The IFF-Protocol defines the training set as a combination of Deepfake, IMDL, AIGC, and Document data, where each training epoch samples an equal number of instances from each domain at random.
-During training, we select FaceForensics++ rossler2019faceforensics++ from Deepfake, CASIAv2 CASIA_2013 from IMDL, GenImage zhu2023genimage from AIGC, and OSTF qu2025ostf , RealTextManipulation luo2025rtm , T-SROIE wang2022tsroie , and Tampered-IC13 wang2022tpic from the Document.
+During training, we select FaceForensics++ Rossler et al.
+(2019) from Deepfake, CASIAv2 Dong et al.
+(2013) from IMDL, GenImage Zhu et al.
+(2023) from AIGC, and OSTF Qu et al.
+(2025), RealTextManipulation Luo et al.
+(2025), T-SROIE Wang et al.
+(2022b), and Tampered-IC13 Wang et al.
+(2022a) from the Document.
 We use the smallest dataset, CASIAv2 with 12,641 samples, as the sampling number for each epoch.
@@
 Model Params (M) FLOPs (G) Model Params (M) FLOPs (G) Model Params (M) FLOPs (G)
-Resnet Resnet_2016 $44.55$ $20.59$ Capsule-Net nguyen2019capsule $3.90$ $32.43$ Mesorch zhu2025mesoscopic $85.75$ $57.95$
-Xception chollet2017xception $22.86$ $12.05$ RECCE cao2022recce $25.83$ $16.18$ DualNet dualnet $7.99$ $66.34$
-EfficientNet tan2019efficientnet $19.34$ $4.14$ SPSL liu2021spsl $20.81$ $12.06$ HiFiNet HiFi-Net2023 $6.89$ $145.00$
-ConvNeXt Convnet_2022 $50.22$ $22.74$ MVSS-Net MVSS_2021 $147.00$ $83.08$ UnivFD univfd $428.00$ $156.00$
-Segformer SegFormer_2021 $44.07$ $16.29$ Trufor trufor2023 $68.70$ $112.00$ DTD qu2023doctamper $67.07$ $272.00$
-Swin Swin_2021 $49.61$ $28.65$ IML-ViT ma2023iml $91.78$ $80.37$ FFDN chen2024ffdn $89.20$ $453.00$
+Resnet He et al. (2016) $44.55$ $20.59$ Capsule-Net Nguyen et al. (2019) $3.90$ $32.43$ Mesorch Zhu et al. (2025) $85.75$ $57.95$
+Xception Chollet (2017) $22.86$ $12.05$ RECCE Cao et al. (2022) $25.83$ $16.18$ DualNet Xi et al. (2023) $7.99$ $66.34$
+EfficientNet Tan and Le (2019) $19.34$ $4.14$ SPSL Liu et al. (2021a) $20.81$ $12.06$ HiFiNet Guo et al. (2023) $6.89$ $145.00$
+ConvNeXt Liu et al. (2022b) $50.22$ $22.74$ MVSS-Net Chen et al. (2021) $147.00$ $83.08$ UnivFD Ojha et al. (2023) $428.00$ $156.00$
+Segformer Xie et al. (2021) $44.07$ $16.29$ Trufor Guillaro et al. (2023) $68.70$ $112.00$ DTD Qu et al. (2023) $67.07$ $272.00$
+Swin Liu et al. (2021b) $49.61$ $28.65$ IML-ViT Ma et al. (2023b) $91.78$ $80.37$ FFDN Chen et al. (2024) $89.20$ $453.00$
 Model Efficiency.
@@
 $F$-c40 CDFv2 DFD Columbia IMD2020 Autosplice DF GenImage T-SROIE OSTF RTM
-Resnet Resnet_2016 (CVPR16) 0.681 0.730 0.793 0.482 0.533 0.738 0.619 0.797 0.951 0.681 0.662 0.697
-Xception chollet2017xception (CVPR17) 0.728 0.719 0.870 0.465 0.537 0.756 0.757 0.980 0.966 0.762 0.734 0.752
-EfficientNet tan2019efficientnet (ICML19) 0.504 0.535 0.517 0.623 0.506 0.483 0.544 0.597 0.884 0.581 0.512 0.571
-Segformer SegFormer_2021 (NIPS21) 0.691 0.748 0.862 0.409 0.562 0.824 0.805 0.998 0.980 0.866 0.736 0.771
-Swin Swin_2021 (ICCV21) 0.771 0.746 0.901 0.636 0.631 0.864 0.915 0.999 0.990 0.856 0.758 0.824
-ConvNext Convnet_2022 (CVPR22) 0.794 0.784 0.911 0.625 0.598 0.825 0.895 1.000 0.994 0.849 0.762 0.822
-Capsule-Net nguyen2019capsule (ICASSP19) 0.613 0.660 0.699 0.330 0.527 0.745 0.546 0.971 0.946 0.704 0.670 0.674
-RECCE cao2022recce (CVPR22) 0.634 0.602 0.727 0.506 0.492 0.642 0.684 0.906 0.542 0.688 0.555 0.634
-SPSL liu2021spsl (CVPR21) 0.730 0.726 0.876 0.419 0.545 0.759 0.770 0.987 0.972 0.769 0.738 0.754
-Sia sia (ECCV22) 0.629 0.584 0.671 0.653 0.483 0.626 0.593 0.748 0.610 0.677 0.574 0.622
-Effort effort (ICML25) 0.805 0.846 0.930 0.979 0.861 0.943 0.930 0.992 0.960 0.834 0.732 0.892
-MVSS-Net MVSS_2021 (ICCV21) 0.713 0.700 0.857 0.298 0.539 0.795 0.671 0.994 0.978 0.806 0.741 0.736
-Trufor trufor2023 (CVPR23) 0.642 0.698 0.832 0.306 0.564 0.808 0.726 0.996 0.979 0.805 0.732 0.735
-IML-ViT ma2023iml (Arxiv) 0.750 0.726 0.851 0.483 0.556 0.819 0.627 0.991 0.972 0.800 0.703 0.753
-Mesorch zhu2025mesoscopic (AAAI25) 0.767 0.814 0.867 0.285 0.570 0.773 0.629 0.996 0.982 0.819 0.739 0.749
-DualNet dualnet (APSIPA23) 0.637 0.552 0.540 0.268 0.517 0.748 0.899 0.988 0.935 0.658 0.657 0.673
-HiFiNet HiFi-Net2023 (CVPR23) 0.587 0.611 0.648 0.745 0.534 0.677 0.575 0.756 0.937 0.663 0.615 0.668
-UnivFD univfd (CVPR23) 0.690 0.671 0.798 0.886 0.786 0.785 0.742 0.813 0.938 0.684 0.569 0.760
-FatFormer fatformer (CVPR24) 0.842 0.770 0.866 0.199 0.585 0.784 0.941 0.999 0.983 0.806 0.751 0.758
-CO-SPY co-spy (CVPR25) 0.819 0.780 0.875 0.460 0.716 0.779 0.940 0.989 0.969 0.836 0.748 0.829
-DTD qu2023doctamper (CVPR23) 0.498 0.520 0.490 0.679 0.498 0.506 0.457 0.499 0.748 0.595 0.496 0.544
-FFDN chen2024ffdn (ECCV24) 0.714 0.699 0.871 0.553 0.624 0.927 0.999 1.000 0.997 0.893 0.782 0.824
+Resnet He et al. (2016)(CVPR16) 0.681 0.730 0.793 0.482 0.533 0.738 0.619 0.797 0.951 0.681 0.662 0.697
+Xception Chollet (2017)(CVPR17) 0.728 0.719 0.870 0.465 0.537 0.756 0.757 0.980 0.966 0.762 0.734 0.752
+EfficientNet Tan and Le (2019)(ICML19) 0.504 0.535 0.517 0.623 0.506 0.483 0.544 0.597 0.884 0.581 0.512 0.571
+Segformer Xie et al. (2021)(NIPS21) 0.691 0.748 0.862 0.409 0.562 0.824 0.805 0.998 0.980 0.866 0.736 0.771
+Swin Liu et al. (2021b)(ICCV21) 0.771 0.746 0.901 0.636 0.631 0.864 0.915 0.999 0.990 0.856 0.758 0.824
+ConvNext Liu et al. (2022b)(CVPR22) 0.794 0.784 0.911 0.625 0.598 0.825 0.895 1.000 0.994 0.849 0.762 0.822
+Capsule-Net Nguyen et al. (2019)(ICASSP19) 0.613 0.660 0.699 0.330 0.527 0.745 0.546 0.971 0.946 0.704 0.670 0.674
+RECCE Cao et al. (2022)(CVPR22) 0.634 0.602 0.727 0.506 0.492 0.642 0.684 0.906 0.542 0.688 0.555 0.634
+SPSL Liu et al. (2021a)(CVPR21) 0.730 0.726 0.876 0.419 0.545 0.759 0.770 0.987 0.972 0.769 0.738 0.754
+Sia Sun et al. (2022a)(ECCV22) 0.629 0.584 0.671 0.653 0.483 0.626 0.593 0.748 0.610 0.677 0.574 0.622
+Effort Yan et al. (2024c)(ICML25) 0.805 0.846 0.930 0.979 0.861 0.943 0.930 0.992 0.960 0.834 0.732 0.892
+MVSS-Net Chen et al. (2021)(ICCV21) 0.713 0.700 0.857 0.298 0.539 0.795 0.671 0.994 0.978 0.806 0.741 0.736
+Trufor Guillaro et al. (2023)(CVPR23) 0.642 0.698 0.832 0.306 0.564 0.808 0.726 0.996 0.979 0.805 0.732 0.735
+IML-ViT Ma et al. (2023b)(Arxiv) 0.750 0.726 0.851 0.483 0.556 0.819 0.627 0.991 0.972 0.800 0.703 0.753
+Mesorch Zhu et al. (2025)(AAAI25) 0.767 0.814 0.867 0.285 0.570 0.773 0.629 0.996 0.982 0.819 0.739 0.749
+DualNet Xi et al. (2023)(APSIPA23) 0.637 0.552 0.540 0.268 0.517 0.748 0.899 0.988 0.935 0.658 0.657 0.673
+HiFiNet Guo et al. (2023)(CVPR23) 0.587 0.611 0.648 0.745 0.534 0.677 0.575 0.756 0.937 0.663 0.615 0.668
+UnivFD Ojha et al. (2023)(CVPR23) 0.690 0.671 0.798 0.886 0.786 0.785 0.742 0.813 0.938 0.684 0.569 0.760
+FatFormer Liu et al. (2024)(CVPR24) 0.842 0.770 0.866 0.199 0.585 0.784 0.941 0.999 0.983 0.806 0.751 0.758
+CO-SPY Cheng et al. (2025)(CVPR25) 0.819 0.780 0.875 0.460 0.716 0.779 0.940 0.989 0.969 0.836 0.748 0.829
+DTD Qu et al. (2023)(CVPR23) 0.498 0.520 0.490 0.679 0.498 0.506 0.457 0.499 0.748 0.595 0.496 0.544
+FFDN Chen et al. (2024)(ECCV24) 0.714 0.699 0.871 0.553 0.624 0.927 0.999 1.000 0.997 0.893 0.782 0.824
 Benchmark Result.
-Table 7 shows the AUC scores of backbones and domain-specific SoTA methods on datasets from four domains under the IFF-Protocol, in which DFD refers to DeepFakeDetection dfd2019 , DF refers to DiffusionForensics wang2023dire , and RTM refers to RealTextManipulation luo2025rtm .
+Table 7 shows the AUC scores of backbones and domain-specific SoTA methods on datasets from four domains under the IFF-Protocol, in which DFD refers to DeepFakeDetection Google AI Blog (2019), DF refers to DiffusionForensics Wang et al.
+(2023), and RTM refers to RealTextManipulation Luo et al.
+(2025).
 We provide detailed results in Appendix F.
-The results show that surprisingly, visual backbones such as ConvNeXt Convnet_2022 and Swin Transformer Swin_2021 outperform almost all domain-specific SoTA methods, indicating that backbones demonstrate greater potential when trained on more unified fake images.
+The results show that surprisingly, visual backbones such as ConvNeXt Liu et al.
+(2022b) and Swin Transformer Liu et al.
+(2021b) outperform almost all domain-specific SoTA methods, indicating that backbones demonstrate greater potential when trained on more unified fake images.
 Meanwhile, domain-specific SoTAs do not necessarily retain their superiority within their own tasks.
-For instance, UnivFD univfd , a CLIP-based fine-tuned model for AIGC detection, demonstrates strong performance on the IMD2020 IMD20_2020 from IMDL, revealing valuable insights into the transferability of cross-task methods.
+For instance, UnivFD Ojha et al.
+(2023), a CLIP-based fine-tuned model for AIGC detection, demonstrates strong performance on the IMD2020 Novozamsky et al.
+(2020) from IMDL, revealing valuable insights into the transferability of cross-task methods.
 From a task perspective, although IMDL target shifts from pixel-level to image-level classification, it remains challenging due to significant distribution differences across datasets in terms of size and manipulation types.
@@
 Since each domain has proposed specific feature extractors, to explore their effectiveness under a unified domain, we conduct experiments using 6 backbones combined with 4 different extractors in the shallow layer under the aforementioned IFF-Protocol setting.
-The extractors are BayarConv Bayar_2018 for noise aritfacts, Sobel MVSS_2021 for edge arifacts, DCT ahmed2006discrete and FPH qu2023doctamper for frequency artifacts.
+The extractors are BayarConv Bayar and Stamm (2018) for noise aritfacts, Sobel Chen et al.
+(2021) for edge arifacts, DCT Ahmed et al.
+(2006) and FPH Qu et al.
+(2023) for frequency artifacts.
 Details of each extractor can be found in Appendix G.1.
 Results in Table 8 show the AUC differences between versions of each backbone using the four different feature extractors and those without, averaged across all test datasets for each task.
-All backbones except EfficientNet tan2019efficientnet show performance drops after using feature extractors, indicating that under the IFF-Protocol, where training data includes sufficient manipulation types and image quantity, models do not rely on the additional information provided by feature extractors.
+All backbones except EfficientNet Tan and Le (2019) show performance drops after using feature extractors, indicating that under the IFF-Protocol, where training data includes sufficient manipulation types and image quantity, models do not rely on the additional information provided by feature extractors.
 However, due to its lightweight nature, EfficientNet still benefits from the use of feature extractors.
@@
 Table 8: Mean AUC differences between extractor-enhanced models and their plain counterparts. Positive values (red) indicate gains; negative (blue) indicate drops.
-Task Extractor = Sobel MVSS_2021 Extractor = BayarBayar_2018
+Task Extractor = Sobel Chen et al. (2021) Extractor = BayarBayar and Stamm (2018)
 ResNet Xception EfficientNet SwinTransformer SegFormer ConvNext ResNet Xception EfficientNet SwinTransformer SegFormer ConvNext
@@
 IMDL 0.0093 -0.0203 0.0058 -0.0838 -0.0642 -0.0687 -0.0118 -0.0227 0.0085 -0.0938 -0.0735 -0.0913
-Extractor = FPH qu2023doctamper Extractor = DCT ahmed2006discrete
+Extractor = FPH Qu et al. (2023) Extractor = DCT Ahmed et al. (2006)
 ResNet Xception EfficientNet SwinTransformer SegFormer ConvNext ResNet Xception EfficientNet SwinTransformer SegFormer ConvNext
@@
 Across both settings, IMDL detectors demonstrate strong competitiveness in the document forensics scenario.
-In the conventional split, the Cat-Net CAT-Net2022 ; qu2023doctamper ; chen2024ffdn family achieves the best average F1, confirming the merit of its hierarchical “cat‑net” paradigm.
-Under the more challenging cross‑domain evaluation, PSCC-Net liu2022pscc displays markedly better generalization, suggesting that progressive spatial modeling captures cues for document manipulation localization.
+In the conventional split, the Cat-Net Kwon et al.
+(2022); Qu et al.
+(2023); Chen et al.
+(2024) family achieves the best average F1, confirming the merit of its hierarchical “cat‑net” paradigm.
+Under the more challenging cross‑domain evaluation, PSCC-Net Liu et al.
+(2022a) displays markedly better generalization, suggesting that progressive spatial modeling captures cues for document manipulation localization.
 We expect future work to further investigate the underlying mechanisms behind PSCC-Net.
 Document → IMDL.
-Following the MVSS training protocol ma2024imdl , all Document-oriented models are trained on the CASIAv2 dataset CASIA_2013 and evaluated on five standard IMDL test sets.
-As shown in Table 9, the dual-branch architecture of CAFTB song2025caftb achieves the best overall performance among Document models when transferred to IMDL tasks—an outcome that aligns with the design philosophy of the current SoTA model Mesorch zhu2025mesoscopic , which also emphasizes dual-branch learning.
+Following the MVSS training protocol Ma et al.
+(2024), all Document-oriented models are trained on the CASIAv2 dataset Dong et al.
+(2013) and evaluated on five standard IMDL test sets.
+As shown in Table 9, the dual-branch architecture of CAFTB Song et al.
+(2025) achieves the best overall performance among Document models when transferred to IMDL tasks—an outcome that aligns with the design philosophy of the current SoTA model Mesorch Zhu et al.
+(2025), which also emphasizes dual-branch learning.
 Table 9: Pixel-level binary F1 evaluation on IMDL benchmarks for document-trained detectors.
 Model CASIAv1 COVERAGE Columbia IMD2020 NIST16 Average
-CAFTB song2025caftb (TOMM24) 0.6234 0.2557 0.6731 0.3526 0.3770 0.4564
-DTD qu2023doctamper (CVPR23) 0.3535 0.1482 0.6470 0.1749 0.1811 0.3009
-FFDN chen2024ffdn (ECCV24) 0.5012 0.2670 0.6085 0.2516 0.2957 0.3848
-TIFDM dong2024tifdm (TCE24) 0.3675 0.2087 0.3572 0.1634 0.2333 0.2660
+CAFTB Song et al. (2025)(TOMM24) 0.6234 0.2557 0.6731 0.3526 0.3770 0.4564
+DTD Qu et al. (2023)(CVPR23) 0.3535 0.1482 0.6470 0.1749 0.1811 0.3009
+FFDN Chen et al. (2024)(ECCV24) 0.5012 0.2670 0.6085 0.2516 0.2957 0.3848
+TIFDM Dong et al. (2024)(TCE24) 0.3675 0.2087 0.3572 0.1634 0.2333 0.2660
 6.2.2 Extending IMDL Detectors to AIGC and Deepfake Benchmarks
@@
 We train each IMDL detector on the FF++‑c23 training split and evaluate on all remaining deep‑fake test sets; the scores are given in Table 10.
-When compared to all baselines in the deepfakebench yan2023deepfakebench , Cat-Net attains the best performance in the within‑domain setting, while Mesorch achieves the highest average accuracy in the cross‑domain evaluation, establishing new state‑of‑the‑art results in both regimes.
+When compared to all baselines in the deepfakebench Yan et al.
+(2023b), Cat-Net attains the best performance in the within‑domain setting, while Mesorch achieves the highest average accuracy in the cross‑domain evaluation, establishing new state‑of‑the‑art results in both regimes.
 Table 10: Image-level AUC evaluation of IMDL-based detectors trained on FF++-c23 and tested across both within-domain and cross-domain deepfake benchmarks.
@@
 $F$-c23 $F$-c40 $F$-DF $F$-F2F $F$-FS $F$-NT Avg. CDFv1 CDFv2 DFD DFDC DFDCP Fsh UADFV Avg.
-CAT-Net CAT-Net2022 (IJCV22) 0.9855 0.8351 0.9946 0.9891 0.9905 0.9689 0.9606 0.7489 0.7511 0.8086 0.7196 0.7043 0.6626 0.9282 0.7605
-IML-ViT ma2023iml (Arxiv) 0.9614 0.8483 0.9810 0.9728 0.9793 0.9206 0.9439 0.7250 0.7419 0.7770 0.7516 0.7237 0.6149 0.8718 0.7437
-Mesorch zhu2025mesoscopic (AAAI25) 0.9815 0.8445 0.9889 0.9890 0.9859 0.9636 0.9589 0.7900 0.7649 0.8468 0.7806 0.7345 0.6638 0.9623 0.7918
-MVSS-Net MVSS_2021 (ICCV21) 0.9723 0.8140 0.9779 0.9865 0.9868 0.9458 0.9472 0.6969 0.6839 0.7749 0.6764 0.6785 0.5841 0.8871 0.7117
-PSCC-Net liu2022pscc (TCSVT22) 0.7147 0.6536 0.7751 0.6414 0.6460 0.6065 0.6729 0.4602 0.5898 0.5565 0.5945 0.5845 0.6574 0.8216 0.6092
-Trufor trufor2023 (CVPR23) 0.7303 0.6848 0.8087 0.7687 0.7084 0.6326 0.7223 0.6551 0.6115 0.5908 0.6017 0.5749 0.5977 0.9526 0.6549
+CAT-Net Kwon et al. (2022)(IJCV22) 0.9855 0.8351 0.9946 0.9891 0.9905 0.9689 0.9606 0.7489 0.7511 0.8086 0.7196 0.7043 0.6626 0.9282 0.7605
+IML-ViT Ma et al. (2023b)(Arxiv) 0.9614 0.8483 0.9810 0.9728 0.9793 0.9206 0.9439 0.7250 0.7419 0.7770 0.7516 0.7237 0.6149 0.8718 0.7437
+Mesorch Zhu et al. (2025)(AAAI25) 0.9815 0.8445 0.9889 0.9890 0.9859 0.9636 0.9589 0.7900 0.7649 0.8468 0.7806 0.7345 0.6638 0.9623 0.7918
+MVSS-Net Chen et al. (2021)(ICCV21) 0.9723 0.8140 0.9779 0.9865 0.9868 0.9458 0.9472 0.6969 0.6839 0.7749 0.6764 0.6785 0.5841 0.8871 0.7117
+PSCC-Net Liu et al. (2022a)(TCSVT22) 0.7147 0.6536 0.7751 0.6414 0.6460 0.6065 0.6729 0.4602 0.5898 0.5565 0.5945 0.5845 0.6574 0.8216 0.6092
+Trufor Guillaro et al. (2023)(CVPR23) 0.7303 0.6848 0.8087 0.7687 0.7084 0.6326 0.7223 0.6551 0.6115 0.5908 0.6017 0.5749 0.5977 0.9526 0.6549
 6.3 Grad-CAM Visualization
@@
 Wentao Feng for the workplace, computation power, and physical infrastructure support.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: The main claims reflect the paper’s contributions and scope, which we build the first unified benchmark for all-domain fake image detection and localization.
+  5. 2.
+    Limitations
+  6. Question: Does the paper discuss the limitations of the work performed by the authors?
+  7. Answer: [Yes]
+  8. Justification: We discuss the limitations in the Appendix.
+  9. 3.
+    Theory assumptions and proofs
+  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  11. Answer: [N/A]
+  12. Justification: This paper does not include theory assumptions and proofs.
+  13. 4.
+    Experimental result reproducibility
+  14.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  15. Answer: [Yes]
+  16. Justification: We provide detailed experimental settings in this paper and open-source our code on GitHub.
+  17. 5.
+    Open access to data and code
+  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  19. Answer: [Yes]
+  20. Justification: We provide detailed experimental settings in this paper and open-source our code on GitHub.
+  21. 6.
+    Experimental setting/details
+  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  23. Answer: [Yes]
+  24. Justification: We provide detailed experimental settings in this paper and open-source our code on GitHub.
+  25. 7.
+    Experiment statistical significance
+  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  27. Answer: [No]
+  28. Justification: All experiments were conducted only once, but given the sufficiently large scale, the results can still serve as a valuable reference.
+  29. 8.
+    Experiments compute resources
+  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  31. Answer: [Yes]
+  32. Justification: We provide details of the experimental hardware setup in the Appendix.
+  33. 9.
+    Code of ethics
+  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  35. Answer: [Yes]
+  36. Justification: All experiments in this paper are conducted using existing public datasets and comply with ethical standards.
+  37. 10.
+    Broader impacts
+  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  39. Answer: [Yes]
+  40. Justification: We discuss the proposed paper’s positive and negative societal impacts in the Appendix.
+  41. 11.
+    Safeguards
+  42.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  43. Answer: [N/A]
+  44. Justification: This paper posed no such risks.
+  45. 12.
+    Licenses for existing assets
+  46.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  47. Answer: [Yes]
+  48. Justification: All resources used in this paper are open-source and publicly available.
+  49. 13.
+    New assets
+  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  51. Answer: [Yes]
+  52. Justification: We provide the detailed document along with the open-source code.
+  53. 14.
+    Crowdsourcing and research with human subjects
+  54.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  55. Answer: [N/A]
+  56. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  57. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  58.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  59. Answer: [N/A]
+  60. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  61. 16.
+    Declaration of LLM usage
+  62.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  63. Answer: [N/A]
+  64. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
 Appendix A Limitations
@@
 However, the comprehensive coverage of detection methods in ForensicHub may enable malicious actors to study and develop targeted evasion techniques.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: The main claims reflect the paper’s contributions and scope, which we build the first unified benchmark for all-domain fake image detection and localization.
-  5. 2.
-    Limitations
-  6. Question: Does the paper discuss the limitations of the work performed by the authors?
-  7. Answer: [Yes]
-  8. Justification: We discuss the limitations in the Appendix.
-  9. 3.
-    Theory assumptions and proofs
-  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  11. Answer: [N/A]
-  12. Justification: This paper does not include theory assumptions and proofs.
-  13. 4.
-    Experimental result reproducibility
-  14.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  15. Answer: [Yes]
-  16. Justification: We provide detailed experimental settings in this paper and open-source our code on GitHub.
-  17. 5.
-    Open access to data and code
-  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  19. Answer: [Yes]
-  20. Justification: We provide detailed experimental settings in this paper and open-source our code on GitHub.
-  21. 6.
-    Experimental setting/details
-  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  23. Answer: [Yes]
-  24. Justification: We provide detailed experimental settings in this paper and open-source our code on GitHub.
-  25. 7.
-    Experiment statistical significance
-  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  27. Answer: [No]
-  28. Justification: All experiments were conducted only once, but given the sufficiently large scale, the results can still serve as a valuable reference.
-  29. 8.
-    Experiments compute resources
-  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  31. Answer: [Yes]
-  32. Justification: We provide details of the experimental hardware setup in the Appendix.
-  33. 9.
-    Code of ethics
-  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  35. Answer: [Yes]
-  36. Justification: All experiments in this paper are conducted using existing public datasets and comply with ethical standards.
-  37. 10.
-    Broader impacts
-  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  39. Answer: [Yes]
-  40. Justification: We discuss the proposed paper’s positive and negative societal impacts in the Appendix.
-  41. 11.
-    Safeguards
-  42.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  43. Answer: [N/A]
-  44. Justification: This paper posed no such risks.
-  45. 12.
-    Licenses for existing assets
-  46.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  47. Answer: [Yes]
-  48. Justification: All resources used in this paper are open-source and publicly available.
-  49. 13.
-    New assets
-  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  51. Answer: [Yes]
-  52. Justification: We provide the detailed document along with the open-source code.
-  53. 14.
-    Crowdsourcing and research with human subjects
-  54.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  55. Answer: [N/A]
-  56. Justification: The paper does not involve crowdsourcing nor research with human subjects.
-  57. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  58.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  59. Answer: [N/A]
-  60. Justification: The paper does not involve crowdsourcing nor research with human subjects.
-  61. 16.
-    Declaration of LLM usage
-  62.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  63. Answer: [N/A]
-  64. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.